### v1

# On oscillator death in the Winfree model I: the one-dimensional case

Abstract.
We show that for the standard sinusoidal Winfree model, a coupling strength exceeding twice the maximal magnitude of the intrinsic frequencies guarantees the convergence of the system for Lebesgue almost every initial data. This is proven by first showing, via an order parameter bootstrapping argument, that the pathwise critical coupling strength is upper bounded by a function of the order parameter, and then showing by a volumetric argument that for Lebesgue almost every data the order parameter cannot stay below and be bounded away from 1 for all time; this is a Winfree model counterpart of the analysis of Ha and the author (2020) performed for the Kuramoto model. Using concentration of measure and the aforementioned volumetric argument, we show that, except possibly on a set of very small measure, oscillator death is observed in finite time; this rigorously demonstrates the existence of the oscillator death regime numerically observed by Ariaratnam and Strogatz (2001). These results are robust under many other choices of interaction functions often considered for the Winfree model. We demonstrate that the asymptotic dynamics described in this paper are sharp by analyzing the equilibria of the Winfree model, and we bound the total number of equilibria using a polynomial description. The sequel to this paper will deal with the Winfree model on higher-dimensional manifolds, such as the Euclidean spheres and the unitary groups.
Key words and phrases:
Complete oscillator death, Winfree model, order parameter, partial oscillator death, divergence, concentration of measure
1991 Mathematics Subject Classification:
34C11, 34C15, 34D06
Acknowledgements: Partially supported by the Korea Foundation for Advanced Studies. Part of this work was conducted under the auspices of the Hyperbolic and Kinetic Equations research group at Seoul National University in the academic year 2020-2021. The primary impetus for writing this article was discussions with Myeongju Kang and Bora Moon, who have indicated that a version of [HPR15] that is more robust to generic initial data would be useful in their work on the infinite-$N$ limit of the Winfree model.
1. Introduction
The Winfree model is given as the initial value problem
$\begin{cases}\dot{\theta}_{i}(t)=\omega_{i}+\frac{\kappa}{N}\sum_{j=1}^{N}I(\theta_{j}(t))S(\theta_{i}(t)),\quad t>0,\\ \theta_{i}(0)=\theta_{i}^{0},\end{cases}\quad i=1,\cdots,N,$ (1)
for the $N\geq 1$ real variables $\theta_{1},\cdots,\theta_{N}$, where $\omega_{1},\cdots,\omega_{N}\in\mathbb{R}$ (the intrinsic frequencies), $\kappa\in\mathbb{R}$ (the coupling strength), and $\theta_{1}^{0},\cdots,\theta_{N}^{0}\in\mathbb{R}$ (the initial data) are fixed parameters, and $I:\mathbb{R}\to\mathbb{R}$ (the influence function) and $S:\mathbb{R}\to\mathbb{R}$ (the sensitivity function) are $2\pi$-periodic Lipschitz functions. Uniqueness and global existence of solutions to (1) follow from the standard Cauchy-Lipschitz theory.
The purpose of this paper is to rigorously establish criteria for boundedness of solutions to (1), which we will refer to as the oscillator death phenomenon.
In the prototypical case
$S(\theta)=-\sin\theta,\quad I(\theta)=1+\cos\theta,$ (2)
the initial value problem (1) becomes
$\begin{cases}\dot{\theta}_{i}(t)=\omega_{i}-\frac{\kappa}{N}\sum_{j=1}^{N}(1+\cos\theta_{j}(t))\sin\theta_{i}(t),\quad t>0,\\ \theta_{i}(0)=\theta_{i}^{0},\end{cases}\quad i=1,\cdots,N.$ (3)
We define the order parameter
$R\coloneqq\frac{1}{N}\sum_{j=1}^{N}(1+\cos\theta_{j}),$ (4)
with which one can rewrite the ordinary differential equation of (3) in the mean-field form
$\dot{\theta}_{i}=\omega_{i}-\kappa R\sin\theta_{i}.$ (5)
The main result of this paper stated in the special case (2) is the following theorem.
Theorem 1.
Fix parameters $\{\omega_{i}\}_{i=1}^{N}\in\mathbb{R}^{N}$ and $\kappa>0$ such that
$\kappa>2\max_{i=1,\cdots,N}|\omega_{i}|.$
Then for Lebesgue almost every initial data $\{\theta_{i}^{0}\}_{i=1}^{N}$, the solution $\{\theta_{i}(t)\}_{i=1}^{N}$ to (3) satisfies the following properties.
  1. (a)
    (Oscillator death) For all $i=1,\cdots,N$, the limit $\lim_{t\to\infty}\theta_{i}(t)$ exists and $\lim_{t\to\infty}\dot{\theta}_{i}(t)=0$.
  2. (b)
    (Uniform lower bound on the order parameter) We have $R_{\infty}\coloneqq\lim_{t\to\infty}R(t)\geq\sqrt{\frac{1}{2}+\sqrt{\frac{1}{4}-\frac{\max_{i}|\omega_{i}|^{2}}{\kappa^{2}}}}$.
Remark 2.
It is clear from (5) and Theorem 1 that the structure of the equilibrium in Theorem 1 is given as $\lim_{t\to\infty}\theta_{i}(t)\equiv\sin^{-1}\frac{\omega_{i}}{\kappa R_{\infty}}\mathrm{~or~}\pi-\sin^{-1}\frac{\omega_{i}}{\kappa R_{\infty}}\mod 2\pi$ for each $i=1,\cdots,N$.
Corollary 3.
Fix parameters $\{\omega_{i}\}_{i=1}^{N}\in\mathbb{R}^{N}$, and for each coupling strength $\kappa>0$ and initial data $\{\theta_{i}^{0}\}_{i=1}^{N}\in\mathbb{R}^{N}$ let $\{\theta_{i}(t)\}_{i=1}^{N}$ be the solution to (3). Let $R(t)$ denote the order parameter of the solution. Then
$\liminf_{\kappa\to\infty}\underset{\{\theta_{i}^{0}\}_{i=1}^{N}\in\mathbb{R}^{N}}{\operatorname{ess~inf}}\liminf_{t\to\infty}R(t)\geq 1.$
Numerical simulations suggest that the limiting equilibrium for almost every initial data is that of [HPR15], which is confined in a small neighborhood of 0 modulo $2\pi$, which becomes arbitrarily small as one takes $\kappa\to\infty$. We thus pose the following question.
Question 4.
In the setting of Corollary 3, is the stronger statement
$\lim_{\kappa\to\infty}\underset{\{\theta_{i}^{0}\}_{i=1}^{N}\in\mathbb{R}^{N}}{\operatorname{ess~inf}}\liminf_{t\to\infty}R(t)=2$
true?
Remark 5.
By Theorem 34 below, Corollary 3 is false if we take $\inf$ instead of $\operatorname{ess~inf}$, because it tells us that
$\limsup_{\kappa\to\infty}\inf_{\{\theta_{i}^{0}\}_{i=1}^{N}\in\mathbb{R}^{N}}\limsup_{t\to\infty}R(t)\leq\frac{3}{N}.$
On the other hand, [HPR15, Theorem 2.2] (stated below in Theorem 18) tells us that a variant of Question 4 where we interchange the order of quantifiers is affirmative:
$\underset{\{\theta_{i}^{0}\}_{i=1}^{N}\in\mathbb{R}^{N}}{\operatorname{ess~inf}}\liminf_{\kappa\to\infty}\liminf_{t\to\infty}R(t)=2.$
For general interaction functions $I$ and $S$, we will show that as long as $I$ and $S$ mimic (2) in a certain sense, then with a sufficiently large $\kappa$ we have oscillator death, i.e., boundedness of solutions, for Lebesgue almost every data. See subsection 2.3 for the statements and further discussion.
Winfree proposed the model (1) in [Win67] to model the behavior of biological oscillators as linearly perturbative periodic solutions to (1), when $\kappa$ is very small (i.e. the system is “weakly coupled”) and the $\omega_{i}$ are concentrated around a single positive value (i.e. the system has “small bandwidth”). Ariaratnam and Strogatz [AS01] observed that in the mathematically tractable special case (2), which is representative of the behavior of some biological oscillators, if one strengthens the parameters $\kappa$ and widens the distribution of the $\omega_{i}$’s, the system exhibits a variety of asymptotic behavior. In particular, if $\kappa$ is sufficiently large compared to the distribution of the $\omega_{i}$’s, then their numerical simulations suggest that $\lim_{t\to\infty}\theta_{i}(t)/t=0$ for all $i=1,\cdots,N$, regardless of the initial value.
Here is the exact setup of [AS01]. They take $N=800$, a sample of $\theta_{i}^{0}\in[-\pi,\pi]$, a (nonrandom) sample of $\omega_{i}$ from the uniform distribution on $1-\gamma,1+\gamma]$, and some fixed $\kappa\geq 0$, and then simulate (3) with these parameters up to time $t=500$ (sometimes longer times were required). Then they observed that the qualitative profile of the “rotation numbers”
$\rho_{i}\coloneqq\lim_{t\to\infty}\theta_{i}(t)/t,\quad i=1,\cdots,N$
does not depend on the initial conditions $\theta_{i}^{0}$, and they drew a phase diagram depicting the configuration of the rotation numbers for different choices of the parameters $\kappa\in[0,1]$ and $\gamma\in[0,1]$. The large $\kappa$ regime of this phase diagram corresponds to the “death” regime, where $\rho_{i}=0$ for all $i=1,\cdots,N$.
Note that oscillator death implies $\lim_{t\to\infty}\theta_{i}(t)/t=0$ for all $i=1,\cdots,N$. Theorem 1 implies the oscillator death phenomenon for the model (3) when $\kappa>2\max_{i=1,\cdots,N}|\omega_{i}|$, thus rigorously confirming the numerical observation of the death regime by Ariaratnam and Strogatz.
To be more precise, in the following theorem, we can compute an upper bound on the measure of the exceptional set that does not exhibit bounded behavior starting at time $t=0$, which is in a sense a quantitative version of Theorem 1. Below, we introduce the notation
$\|\Omega\|_{\infty}\coloneqq\max_{i=1,\cdots,N}|\omega_{i}|.$
Theorem 6.
Fix parameters $\{\omega_{i}\}_{i=1}^{N}\in\mathbb{R}^{N}$, $\kappa>0$, and $\varepsilon\in(0,1]$ such that
$\kappa>(2+\varepsilon)\|\Omega\|_{\infty}.$
For each initial data $\{\theta_{i}^{0}\}_{i=1}^{N}$, denote by $\{\theta_{i}(t)\}_{i=1}^{N}$ the solution to (3). Then, denoting by $m$ the normalized Lebesgue measure on $\pi,\pi]^{N}$ such that $m([-\pi,\pi]^{N})=1$, one has
$m\Big\{\{\theta_{i}^{0}\}_{i=1}^{N}\in[-\pi,\pi]^{N}:~\forall t\geq 0~R(t)\geq\frac{1}{\sqrt{2}}-\frac{3\varepsilon}{20},$
$\qquad\forall i=1,\cdots,N~\exists\lim_{t\to\infty}\theta_{i}(t),~\sup_{t\geq 0}\theta_{i}(t)-\inf_{t\geq 0}\theta_{i}(t)<2\pi,~\lim_{t\to\infty}\dot{\theta}_{i}(t)=0$
$\qquad\mathrm{and~}\forall i\in\{1,\cdots,N\}\forall t_{0}\geq 0\mathrm{~if~}\cos\theta_{i}(t_{0})\geq-\frac{1}{\sqrt{2}}\mathrm{~then}$
$\qquad\cos\theta_{i}(t_{0})\geq-\frac{1}{\sqrt{2}}\mathrm{~for~}t\geq t_{0}\mathrm{~and~}\cos\theta_{i}(t)\geq\frac{1}{\sqrt{2}}\mathrm{~for~}t\geq t_{0}+\frac{(2+\varepsilon)\pi}{\kappa-(2+\varepsilon)\|\Omega\|_{\infty}}\Big\}$
$\geq 1-\exp\left(-\frac{\varepsilon^{2}N}{25}\right).$
If, in addition, $\kappa>\left(\frac{\pi e}{2}\right)^{3/2}\|\Omega\|_{\infty}\approx 8.8231\|\Omega\|_{\infty},$ then
$m\Big\{\{\theta_{i}^{0}\}_{i=1}^{N}\in[-\pi,\pi]^{N}:\exists\lim_{t\to\infty}\theta_{i}(t),~\sup_{t\geq 0}\theta_{i}(t)-\inf_{t\geq 0}\theta_{i}(t)<2\pi,~\lim_{t\to\infty}\dot{\theta}_{i}(t)=0~\mbox{for all }i=1,\cdots,N\Big\}$
$\geq 1-\left(\sqrt{\frac{\pi e}{2}}\frac{\|\Omega\|_{\infty}^{1/3}}{\kappa^{1/3}}\right)^{N}.$
Furthermore, for any time $T>0$, we can compute an upper bound on the measure of the exceptional set that does not exhibit bounded behavior up to time $t=T$, strengthening Theorem 6.
Theorem 7.
Fix parameters [^1] The reason we don’t consider the case $N=1$ is because oscillator death is trivial; see Proposition 46 with $N=1$. $N\geq 2$, $\{\omega_{i}\}_{i=1}^{N}\in\mathbb{R}^{N}$, $\kappa>0$, and $\varepsilon\in(0,1]$ such that
$\kappa>(2+\varepsilon)\|\Omega\|_{\infty}.$
For each initial data $\{\theta_{i}^{0}\}_{i=1}^{N}$, denote by $\{\theta_{i}(t)\}_{i=1}^{N}$ the solution to (3). Then, denoting by $m$ the normalized Lebesgue measure on $\pi,\pi]^{N}$ such that $m([-\pi,\pi]^{N})=1$, one has, for each time $T>0$,
$m\Big\{\{\theta_{i}^{0}\}_{i=1}^{N}\in[-\pi,\pi]^{N}:~\forall t\geq T~R(t)\geq\frac{1}{\sqrt{2}}-\frac{3\varepsilon}{20},$
$\qquad\forall i=1,\cdots,N~\exists\lim_{t\to\infty}\theta_{i}(t),~\sup_{t\geq T}\theta_{i}(t)-\inf_{t\geq T}\theta_{i}(t)<2\pi,~\lim_{t\to\infty}\dot{\theta}_{i}(t)=0$
$\qquad\mathrm{and~}\forall i\in\{1,\cdots,N\}\forall t_{0}\geq T\mathrm{~if~}\cos\theta_{i}(t_{0})\geq-\frac{1}{\sqrt{2}}\mathrm{~then}$
$\qquad\cos\theta_{i}(t_{0})\geq-\frac{1}{\sqrt{2}}\mathrm{~for~}t\geq t_{0}\mathrm{~and~}\cos\theta_{i}(t)\geq\frac{1}{\sqrt{2}}\mathrm{~for~}t\geq t_{0}+\frac{(2+\varepsilon)\pi}{\kappa-(2+\varepsilon)\|\Omega\|_{\infty}}\Big\}$
$\geq\begin{cases}1-\frac{1}{N/2+1}\left(\frac{\sqrt{\pi e\varepsilon}}{2\sqrt{5}}\right)^{N}\left(1-\exp\left(-\frac{\kappa N\varepsilon(5-\varepsilon)T}{25}\right)\right)+\exp\left(-\frac{N\varepsilon^{2}}{25}-\frac{\kappa N\varepsilon(5-\varepsilon)T}{25}\right),&\mathrm{if~}T\leq T_{0},\\ 1-\left(\frac{20}{\pi e\varepsilon}+\frac{4\kappa(5-\varepsilon)}{5\pi e}(T-T_{0})\right)^{-\frac{N}{2}},&\mathrm{if~}T>T_{0},\end{cases}$
where $T_{0}=\frac{25}{\kappa N\varepsilon(5-\varepsilon)}\log\left(\left(1+\frac{2}{N}\right)\left(\frac{2\sqrt{5}}{\sqrt{\pi\varepsilon}e^{1/2+\varepsilon^{2}/25}}\right)^{N}-\frac{2}{N}\right)$.
Furthermore, if $\kappa>8\|\Omega\|_{\infty},$ then [^2] If $\|\Omega\|_{\infty}=0$ so that a fraction in the expression below is undefined, then the stated probability is 1.
$m\Big\{\{\theta_{i}^{0}\}_{i=1}^{N}\in[-\pi,\pi]^{N}:\exists\lim_{t\to\infty}\theta_{i}(t),~\sup_{t\geq T}\theta_{i}(t)-\inf_{t\geq T}\theta_{i}(t)<2\pi,~\lim_{t\to\infty}\dot{\theta}_{i}(t)=0~\mbox{for all }i=1,\cdots,N\Big\}$
$\geq\begin{cases}1-\left(e^{1/2}+\frac{3\kappa T}{\pi e}\right)^{-\frac{N}{2}}&\mathrm{if~}8\|\Omega\|_{\infty}<\kappa\leq 16\sqrt{2}\|\Omega\|_{\infty},\\ 1-\left(\frac{4}{\pi e}\left(\frac{\kappa}{2\sqrt{2}\|\Omega\|_{\infty}}\right)^{2/3}+\frac{3\kappa T}{\pi e}\right)^{-\frac{N}{2}}&\mathrm{if~}\kappa>16\sqrt{2}\|\Omega\|_{\infty}.\end{cases}$
Theorems 6 and 7 rigorously confirm the experimental observations of [AS01] in a more precise manner. Namely, Theorem 6 tells us that if $\kappa>(2+\varepsilon)\max_{i}|\omega_{i}|$, then if we independently sample $\theta_{i}^{0}$ uniformly from $\pi,\pi]$ and simulate (3), we will then observe that the oscillators stay within $2\pi$ of its initial value for all time with probability at least $1-\exp\left(-\frac{\varepsilon^{2}N}{25}\right)$, which tends rapidly to $1$ as $N\to\infty$. For example, if we have $N=800$ and $\kappa>3\max_{i}|\omega_{i}|$, then this probability is at least $1-\exp\left(-800/25\right)\approx 1-1.266\times 10^{-14}$. Theorem 7 goes one step further and tells us that oscillator death behavior will be observed within some large time $T$ with probability at least $1-\left(\frac{20}{\pi e\varepsilon}+\frac{4\kappa(5-\varepsilon)}{5\pi e}(T-T_{0})\right)^{-\frac{N}{2}}$ if $T$ is large enough. For example, if $N=800$, $\max_{i}|\omega_{i}|=2$, $\kappa=6$ (so that $\varepsilon=1$), and $T=500$, then $T_{0}=\frac{25}{4\kappa N}\log\left(\left(1+\frac{2}{N}\right)\left(\frac{2\sqrt{5}}{\sqrt{\pi}e^{1/2+1/25}}\right)^{N}-\frac{2}{N}\right)=0.40156699\cdots$ and the possibility of observing oscillator death within time $T=500$ is at least
$1-\left(\frac{20}{\pi e}+\frac{16\kappa}{5\pi e}(T-T_{0})\right)^{-\frac{N}{2}}\approx 1-2.7990\times 10^{-1221}.$
Although the statements of Theorems 1 and 6 treat $2\max_{i}|\omega_{i}|$ as if it were a critical threshold, this is likely an artifact of our proof, as the phase diagram of [AS01] indicates that oscillator death occurs even for smaller values of $\kappa$. To be precise, we make the following definitions.
Definition 8.
Fix a frequency vector $\Omega=\{\omega_{1},\cdots,\omega_{N}\}\in\mathbb{R}^{N}$.
  1. (a)
    The critical coupling strength is
[…]
  2. (b)
    For an initial phase vector $\Theta^{0}=\{\theta^{0}_{1},\cdots,\theta^{0}_{N}\}\in\mathbb{R}^{N}$, the pathwise critical coupling strength is
[…]
We have a fortiori $\kappa_{\mathrm{c}}(\Omega)\leq\kappa_{\mathrm{pc}}(\Theta^{0},\Omega)$, while by Proposition 63 we have
$\frac{2N}{4N-1}\max_{i=1,\cdots,N}|\omega_{i}|\leq\kappa_{\mathrm{c}}(\Omega)\leq\frac{4}{3\sqrt{3}}\max_{i=1,\cdots,N}|\omega_{i}|.$
Also, statement (a) of Theorem 1 can be summarized as follows.
Theorem 9.
For fixed $\Omega=\{\omega_{1},\cdots,\omega_{N}\}\in\mathbb{R}^{N}$, we have
$\kappa_{\mathrm{pc}}(\Theta^{0},\Omega)\leq 2\max_{i=1,\cdots,N}|\omega_{i}|,\quad\mathrm{~for~a.e.~}\Theta^{0}=\{\theta^{0}_{1},\cdots,\theta^{0}_{N}\}\in\mathbb{R}^{N}.$
We ask whether this holds for all $\Theta^{0}$.
Conjecture 10.
There exists a universal constant $c>0$ such that
$\kappa_{\mathrm{pc}}(\Theta^{0},\Omega)\leq c\max_{i=1,\cdots,N}|\omega_{i}|,\quad\forall\Theta^{0}=\{\theta^{0}_{1},\cdots,\theta^{0}_{N}\}\in\mathbb{R}^{N},~\forall\Omega=\{\omega_{1},\cdots,\omega_{N}\}\in\mathbb{R}^{N}.$
By Theorem 9, $\kappa_{\mathrm{c}}(\Omega)$ and $\kappa_{\mathrm{pc}}(\Theta^{0},\Omega)$ are within universal constant multiples of each other for a.e. $\Theta^{0}$. However, the phase diagram of [AS01] seems to indicate the stronger statement that $\kappa_{\mathrm{c}}(\Omega)$ and $\kappa_{\mathrm{pc}}(\Theta^{0},\Omega)$ coincide. We leave this as a conjecture.
Conjecture 11.
Let $\Omega=\{\omega_{1},\cdots,\omega_{N}\}\in\mathbb{R}^{N}$ be a fixed frequency vector.
  1. (a)
    (Weak version) $\kappa_{\mathrm{c}}(\Omega)=\kappa_{\mathrm{pc}}(\Theta^{0},\Omega)$ for a.e. $\Theta^{0}=\{\theta_{1}^{0},\cdots,\theta_{N}^{0}\}\in\mathbb{R}^{N}$.
  2. (b)
    (Strong version) $\kappa_{\mathrm{c}}(\Omega)=\kappa_{\mathrm{pc}}(\Theta^{0},\Omega)$ for all $\Theta^{0}=\{\theta_{1}^{0},\cdots,\theta_{N}^{0}\}\in\mathbb{R}^{N}$.
Of course, Conjecture 11(b) implies Conjecture 10.
It seems likely that the results obtained here for the ODE level carry on to the infinite-$N$ limits, but this needs to be checked. See Ha, Kang, and Moon [HKM21b] for a description of the continuum limit of (3).
Many of the results of this paper can be extended to the manifold setting, particularly the Euclidean sphere $\mathbb{S}^{d-1}$ and the unitary group $\mathrm{U}(d)$. This will be dealt with in a sequel to this paper.
Notation. We will use capital Greek letters to refer to the tuples of the corresponding lower Greek letters:
$\Omega\coloneqq(\omega_{1},\cdots,\omega_{N}),\quad\Theta\coloneqq(\theta_{1},\cdots,\theta_{N}),$
We will use $\|\cdot\|_{\infty}$ to denote the $\ell^{\infty}$-norm:
$\|\Omega\|_{\infty}\coloneqq\max_{i=1,\cdots,N}|\omega_{i}|,\quad\|\Theta\|_{\infty}\coloneqq\max_{i=1,\cdots,N}|\theta_{i}|.$
Definition 12.
Let $\Theta=(\theta_{1},\cdots,\theta_{N})$ be a solution to the Winfree model (1). We say that the ensemble $\Theta$ exhibits (complete) oscillator death if
$\sup_{t\geq 0}\|\Theta(t)\|_{\infty}<\infty.$
Remark 13.
By Proposition 19 and its following remarks, in the prototypical model (3), oscillator death for $\Theta$ implies the stronger condition
$\exists\lim_{t\to\infty}\theta_{i}(t)\mbox{ and }\lim_{t\to\infty}\dot{\theta}_{i}(t)=0,\quad\forall i=1,\cdots,N.$
Remark 14.
Our notion of oscillator death is stronger than that usually used in the literature, which typically defines oscillator death to be the phenomenon where the rotation numbers
$\rho_{i}\coloneqq\lim_{t\to\infty}\frac{\theta_{i}(t)}{t},\quad i=1,\cdots,N,$
defined whenever the limit exists, equal zero for all $i=1,\cdots,N$.
We will use the following (standard) asymptotic notation. For P,Q>0P,~Q>0, the notations $P\lesssim Q$, $Q\gtrsim P$, $P=O(Q)$, and $Q=\Omega(P)$ mean that $P\leq KQ$ for a universal constant $K\in(0,\infty)$, and $P\asymp Q$ means $(P\lesssim Q)\wedge(Q\lesssim P)$. If we need to allow for dependence on parameters, we indicate this by subscripts. For example, in the presence of auxiliary parameters $\psi,\xi$, the notations $P\lesssim_{\psi,\xi}Q$, $Q\gtrsim_{\psi,\xi}P$, $P=O_{\psi,\xi}(Q)$, $Q=\Omega_{\psi,\xi}(P)$ mean that $P\leq K(\psi,\xi)Q$ where $K(\psi,\xi)\in(0,\infty)$ may depend only on $\psi$ and $\xi$, and $P\asymp_{\psi,\xi}Q$ means $(P\lesssim_{\psi,\xi}Q)\wedge(Q\lesssim_{\psi,\xi}P)$.
Roadmap. The rest of this paper is organized as follows. In Section 2, we list our new technical results. In subsection 2.1 we will state a key intermediate Corollary 16 in proving Theorems 1 and 6. For more general interaction functions $I$ and $S$, we will not be able to conclude the convergence of system (1) unless under very restrictive conditions, and to make this point clear, we will discuss the Łojasiewicz gradient theorem, which is a powerful tool for discussing convergence of (3), in subsection 2.2. In subsection 2.3, we will see how to extend the statements of Theorems 1 and 6 and Corollary 16 to more general interaction functions $I$ and $S$. After that, in subsection 2.4 we will formalize the notion of partial oscillator death and provide criteria for it to occur, and in subsection 2.5 we will discuss the equilibria of (5) to show the sharpness of the asymptotic dynamics of Corollary 16.
In Section 3, we briefly review the relevant literature on the Winfree model, starting from Winfree’s original paper [Win67] and describing the subsequent developments. We will also compare the Winfree model to the Kuramoto model, which was inspired by the Winfree model and which exhibits analogous but different synchronization behavior. We will suggest some lines of possible future research along the way.
Section 4 is devoted to proving Corollary 16 and its variant Theorem 21 via a bootstrapping argument which also gives the sufficient conditions of subsection 2.4 that guarantee partial oscillator death.
In Section 5, we calculate the divergence of the vector field associated to system (1) and use an argument on the rate of change of volume to prove Theorem 1. By applying some large deviations theory and capitalizing on the positivity of the divergence, we will prove Theorems 6 and 7, plus many other new theorems mentioned in Section 2.
In Section 6, we prove the statements of subsection 2.5, namely we obtain some existence results for the equilibria of (3) and demonstrate a polynomial description of the equilibria of (3) that gives us a uniform bound on the total number of equilibria. We also compute the critical coupling strength $\kappa_{\mathrm{c}}$.
Finally, in Appendix A we prove some elementary calculus inequalities that are needed in this paper.
2. Gallery of new results
In this section, we outline our new results regarding the models (1) and (3).
2.1. The pathwise critical coupling strength is bounded by a function of the order parameter
The first step in the proof of Theorem 1 is the following intermediate result on the convergence of system (3).
Theorem 15.
Let $\Theta=\Theta(t)$ be the solution to (3) with parameters $\{\omega_{i}\}_{i=1}^{N}$, $\kappa$, and $\{\theta_{i}^{0}\}_{i=1}^{N}$. Suppose ${R_{0}}>0$, fix $\mu\in(0,\min\{R_{0},1\})$, and suppose
$\kappa>\frac{\|\Omega\|_{\infty}}{({R_{0}}-\mu)\sqrt{\mu(2-\mu)}}.$ (6)
Then
  1. (a)
    $R(t)\geq R_{0}-\mu$ for all $t\geq 0$,
  2. (b)
    For all $i=1,\cdots,N$, the limit $\lim_{t\to\infty}\theta_{i}(t)$ exists, and
    $\sup_{t\geq 0}\theta_{i}(t)-\inf_{t\geq 0}\theta_{i}(t)<2\pi,\quad\lim_{t\to\infty}\dot{\theta}_{i}(t)=0.$
  3. (c)
    for any $i\in\{1,\cdots,N\}$ and $t_{0}\geq 0$ with $\cos\theta_{i}(t_{0})\geq-1+\mu$, we have $\cos\theta_{i}(t)\geq-1+\mu$ for $t\geq t_{0}$ and $\cos\theta_{i}(t)\geq 1-\mu$ for $t\geq t_{0}+\frac{\pi}{\kappa(R_{0}-\mu)\sqrt{\mu(2-\mu)}-\|\Omega\|_{\infty}}$,
  4. (d)
    for any $i\in\{1,\cdots,N\}$, either $\sup_{t\geq 0}\cos\theta_{i}(t)<-1+\mu$, in which case $\lim_{t\to\infty}\theta_{i}(t)=\pi-\sin^{-1}\frac{\omega_{i}}{\kappa\lim_{t\to\infty}R(t)}\mod 2\pi$, or $\sup_{t\geq 0}\cos\theta_{i}(t)\geq-1+\mu$, in which case $\lim_{t\to\infty}\theta_{i}(t)=\sin^{-1}\frac{\omega_{i}}{\kappa\lim_{t\to\infty}R(t)}\mod 2\pi$.
  5. (e)
    For fixed $i,j\in\{1,\cdots,N\}$, suppose there is a time $t_{0}\geq 0$ such that $\cos\theta_{i}(t_{0})\geq-1+\mu$ and $\cos\theta_{j}(t_{0})\geq-1+\mu$. If $\omega_{i}>\omega_{j}$, then after modulo $2\pi$ translations,
    $\theta_{i}(t)-\theta_{j}(t)\leq\frac{\omega_{i}-\omega_{j}}{\kappa(R_{0}-\mu)(1-\mu)}+\pi\exp\left(-\kappa(R_{0}-\mu)(1-\mu)\left(t-t_{0}-\frac{\pi}{\kappa(R_{0}-\mu)\sqrt{\mu(2-\mu)}-{\|\Omega\|_{\infty}}}\right)\right)$
    $\mathrm{for~}t\geq t_{0}+\frac{\pi}{\kappa(R_{0}-\mu)\sqrt{\mu(2-\mu)}-{\|\Omega\|_{\infty}}},$
    and
    $\theta_{i}(t)-\theta_{j}(t)\geq\frac{\omega_{i}-\omega_{j}}{2\kappa}-\frac{\omega_{i}-\omega_{j}}{2\kappa}\exp\left(-2\kappa\left(t-t_{0}-\frac{\pi}{\kappa(R_{0}-\mu)\sqrt{\mu(2-\mu)}-{\|\Omega\|_{\infty}}}-\frac{\pi}{\omega_{i}-\omega_{j}}\right)\right)$
    $\mathrm{for~}t\geq t_{0}+\frac{\pi}{\kappa(R_{0}-\mu)\sqrt{\mu(2-\mu)}-{\|\Omega\|_{\infty}}}+\frac{\pi}{\omega_{i}-\omega_{j}}.$
    If $\omega_{i}=\omega_{j}$, then after modulo $2\pi$ translations, $\lim_{t\to\infty}\left(\theta_{i}(t)-\theta_{j}(t)\right)=0$ exponentially:
    $|\theta_{i}(t)-\theta_{j}(t)|\leq\pi\exp\left(-\kappa(R_{0}-\mu)(1-\mu)\left(t-t_{0}-\frac{\pi}{\kappa(R_{0}-\mu)\sqrt{\mu(2-\mu)}-{\|\Omega\|_{\infty}}}\right)\right)$
    $\mathrm{for~}t\geq t_{0}+\frac{\pi}{\kappa(R_{0}-\mu)\sqrt{\mu(2-\mu)}-{\|\Omega\|_{\infty}}}.$
Applying special values of $\mu$ gives the following Corollary.
Corollary 16.
Let $\Theta=\Theta(t)$ be the solution to (3) with parameters $\{\omega_{i}\}_{i=1}^{N}$, $\kappa$, and $\{\theta_{i}^{0}\}_{i=1}^{N}$. Suppose ${R_{0}}>0$ and
$\kappa>K_{c}\|\Omega\|_{\infty},$ (7)
where
$K_{c}\coloneqq\begin{cases}\frac{2}{{R_{0}}^{3/2}},&\mbox{if }{R_{0}}\leq 1,\\ 2(2-R_{0})+\frac{4}{3\sqrt{3}}(R_{0}-1),&\mbox{if }1<{R_{0}}\leq 2.\end{cases}$ (8)
Then
  1. (a)
    $R(t)\geq\frac{3{R_{0}}-3+\sqrt{{R_{0}}^{2}-2{R_{0}}+9}}{4}\geq\frac{2}{3}{R_{0}}$ for all $t\geq 0$,
  2. (b)
    For all $i=1,\cdots,N$, the limit $\lim_{t\to\infty}\theta_{i}(t)$ exists, and
    $\sup_{t\geq 0}\theta_{i}(t)-\inf_{t\geq 0}\theta_{i}(t)<2\pi,\quad\lim_{t\to\infty}\dot{\theta}_{i}(t)=0.$
  3. (c)
    for any $i\in\{1,\cdots,N\}$ and $t_{0}\geq 0$ with $\cos\theta_{i}(t_{0})\geq\frac{-1+{R_{0}}-\sqrt{{R_{0}}^{2}-2{R_{0}}+9}}{4}$, we have $\cos\theta_{i}(t)\geq\frac{-1+{R_{0}}-\sqrt{{R_{0}}^{2}-2{R_{0}}+9}}{4}$ for $t\geq t_{0}$ and $\cos\theta_{i}(t)\geq\frac{1-{R_{0}}+\sqrt{{R_{0}}^{2}-2{R_{0}}+9}}{4}$ for $t\geq t_{0}+\frac{\pi K_{c}}{\kappa-K_{c}\|\Omega\|_{\infty}}$; for example, for all $i=1,\cdots,N$ and $t_{0}\geq 0$ such that $\cos\theta_{i}(t_{0})\geq-1+\frac{1}{3}{R_{0}}$, one has $\cos\theta_{i}(t)\geq-1+\frac{1}{3}{R_{0}}$ for $t\geq t_{0}$ and $\cos\theta_{i}(t)\geq 1-\frac{R_{0}}{3}$ for $t\geq t_{0}+\frac{\pi K_{c}}{\kappa-K_{c}\max_{i}|\omega_{i}|}$.
  4. (d)
    for any $i\in\{1,\cdots,N\}$, either $\sup_{t\geq 0}\cos\theta_{i}(t)<\frac{-1+{R_{0}}-\sqrt{{R_{0}}^{2}-2{R_{0}}+9}}{4}$, in which case $\lim_{t\to\infty}\theta_{i}(t)=\pi-\sin^{-1}\frac{\omega_{i}}{\kappa\lim_{t\to\infty}R(t)}\mod 2\pi$, or $\sup_{t\geq 0}\cos\theta_{i}(t)\geq\frac{-1+{R_{0}}-\sqrt{{R_{0}}^{2}-2{R_{0}}+9}}{4}$, in which case $\lim_{t\to\infty}\theta_{i}(t)=\sin^{-1}\frac{\omega_{i}}{\kappa\lim_{t\to\infty}R(t)}\mod 2\pi$.
  5. (e)
    For fixed $i,j\in\{1,\cdots,N\}$, suppose there is a time $t_{0}\geq 0$ such that $\cos\theta_{i}(t_{0})\geq-1+\frac{R_{0}}{3}$ and $\cos\theta_{j}(t_{0})\geq-1+\frac{R_{0}}{3}$. If $\omega_{i}>\omega_{j}$, then after modulo $2\pi$ translations,
    $\theta_{i}(t)-\theta_{j}(t)\leq\frac{9(\omega_{i}-\omega_{j})}{2\kappa R_{0}(3-R_{0})}+\pi\exp\left(-\frac{2\kappa R_{0}(3-R_{0})}{9}\left(t-t_{0}-\frac{\pi K_{c}}{\kappa-K_{c}\|\Omega\|_{\infty}}\right)\right)$
    $\mathrm{for~}t\geq t_{0}+\frac{\pi K_{c}}{\kappa-K_{c}{\|\Omega\|_{\infty}}},$
    and
    $\theta_{i}(t)-\theta_{j}(t)\geq\frac{\omega_{i}-\omega_{j}}{2\kappa}-\frac{\omega_{i}-\omega_{j}}{2\kappa}\exp\left(-2\kappa\left(t-t_{0}-\frac{\pi K_{c}}{\kappa-K_{c}{\|\Omega\|_{\infty}}}-\frac{\pi}{\omega_{i}-\omega_{j}}\right)\right)$
    $\mathrm{for~}t\geq t_{0}+\frac{\pi K_{c}}{\kappa-K_{c}\|\Omega\|_{\infty}}+\frac{\pi}{\omega_{i}-\omega_{j}}.$
    If $\omega_{i}=\omega_{j}$, then after modulo $2\pi$ translations, $\lim_{t\to\infty}\left(\theta_{i}(t)-\theta_{j}(t)\right)=0$ exponentially:
    $|\theta_{i}(t)-\theta_{j}(t)|\leq\pi\exp\left(-\frac{2\kappa R_{0}(3-R_{0})}{9}\left(t-t_{0}-\frac{\pi K_{c}}{\kappa-K_{c}{\|\Omega\|_{\infty}}}\right)\right)\mathrm{~for~}t\geq t_{0}+\frac{\pi K_{c}}{\kappa-K_{c}{\|\Omega\|_{\infty}}}.$
Furthermore, if $\kappa>\frac{2\sqrt{2}\|\Omega\|_{\infty}}{R_{0}^{3/2}},$ then
  1. (f)
    $R(t)\geq R_{0}-\frac{4\|\Omega\|_{\infty}^{2}}{\kappa^{2}R_{0}^{2}}$ for all $t\geq 0$,
  2. (g)
    for any $i\in\{1,\cdots,N\}$ and $t_{0}\geq 0$ with $\cos\theta_{i}(t_{0})\geq-1+\frac{4\|\Omega\|_{\infty}^{2}}{\kappa^{2}R_{0}^{2}}$, we have $\cos\theta_{i}(t)\geq-1+\frac{4\|\Omega\|_{\infty}^{2}}{\kappa^{2}R_{0}^{2}}$ for $t\geq t_{0}$ and $\cos\theta_{i}(t)\geq 1-\frac{4\|\Omega\|_{\infty}^{2}}{\kappa^{2}R_{0}^{2}}$ for $t\geq t_{0}+\frac{\pi}{\|\Omega\|_{\infty}\left(1-\frac{8\|\Omega\|_{\infty}^{2}}{\kappa^{2}R_{0}^{3}}\right)}$.
  3. (h)
    for any $i\in\{1,\cdots,N\}$, either $\sup_{t\geq 0}\cos\theta_{i}(t)<-1+\frac{4\|\Omega\|_{\infty}^{2}}{\kappa^{2}R_{0}^{2}}$, in which case $\lim_{t\to\infty}\theta_{i}(t)=\pi-\sin^{-1}\frac{\omega_{i}}{\kappa\lim_{t\to\infty}R(t)}\mod 2\pi$, or $\sup_{t\geq 0}\cos\theta_{i}(t)\geq-1+\frac{4\|\Omega\|_{\infty}^{2}}{\kappa^{2}R_{0}^{2}}$, in which case $\lim_{t\to\infty}\theta_{i}(t)=\sin^{-1}\frac{\omega_{i}}{\kappa\lim_{t\to\infty}R(t)}\mod 2\pi$.
  4. (i)
    For fixed $i,j\in\{1,\cdots,N\}$, suppose there is a time $t_{0}\geq 0$ such that $\cos\theta_{i}(t_{0})\geq-1+\frac{4\|\Omega\|_{\infty}^{2}}{\kappa^{2}R_{0}^{2}}$ and $\cos\theta_{j}(t_{0})\geq-1+\frac{4\|\Omega\|_{\infty}^{2}}{\kappa^{2}R_{0}^{2}}$. If $\omega_{i}>\omega_{j}$, then after modulo $2\pi$ translations,
    $\theta_{i}(t)-\theta_{j}(t)\leq\frac{2(\omega_{i}-\omega_{j})}{\kappa R_{0}\left(1-\frac{4\|\Omega\|_{\infty}^{2}}{\kappa^{2}R_{0}^{2}}\right)}+\pi\exp\left(-\frac{\kappa R_{0}}{2}\left(1-\frac{4\|\Omega\|_{\infty}^{2}}{\kappa^{2}R_{0}^{2}}\right)\left(t-t_{0}-\frac{\pi}{\|\Omega\|_{\infty}\left(1-\frac{8\|\Omega\|_{\infty}^{2}}{\kappa^{2}R_{0}^{3}}\right)}\right)\right)$
    $\mathrm{for~}t\geq t_{0}+\frac{\pi}{\|\Omega\|_{\infty}\left(1-\frac{8\|\Omega\|_{\infty}^{2}}{\kappa^{2}R_{0}^{3}}\right)},$
    and
    $\theta_{i}(t)-\theta_{j}(t)\geq\frac{\omega_{i}-\omega_{j}}{2\kappa}-\frac{\omega_{i}-\omega_{j}}{2\kappa}\exp\left(-2\kappa\left(t-t_{0}-\frac{\pi}{\|\Omega\|_{\infty}\left(1-\frac{8\|\Omega\|_{\infty}^{2}}{\kappa^{2}R_{0}^{3}}\right)}-\frac{\pi}{\omega_{i}-\omega_{j}}\right)\right)$
    $\mathrm{for~}t\geq t_{0}+\frac{\pi}{\|\Omega\|_{\infty}\left(1-\frac{8\|\Omega\|_{\infty}^{2}}{\kappa^{2}R_{0}^{3}}\right)}+\frac{\pi}{\omega_{i}-\omega_{j}}.$
    If $\omega_{i}=\omega_{j}$, then after modulo $2\pi$ translations, $\lim_{t\to\infty}\left(\theta_{i}(t)-\theta_{j}(t)\right)=0$ exponentially:
    $|\theta_{i}(t)-\theta_{j}(t)|\leq\pi\exp\left(-\frac{\kappa R_{0}}{2}\left(1-\frac{4\|\Omega\|_{\infty}^{2}}{\kappa^{2}R_{0}^{2}}\right)\left(t-t_{0}-\frac{\pi}{\|\Omega\|_{\infty}\left(1-\frac{8\|\Omega\|_{\infty}^{2}}{\kappa^{2}R_{0}^{3}}\right)}\right)\right)$
    $\mathrm{for~}t\geq t_{0}+\frac{\pi}{\|\Omega\|_{\infty}\left(1-\frac{8\|\Omega\|_{\infty}^{2}}{\kappa^{2}R_{0}^{3}}\right)}.$
Simply put, we have the following.
Theorem 17.
$\kappa_{\mathrm{pc}}(\Theta^{0},\Omega)\leq\begin{cases}\frac{2}{{R_{0}}^{3/2}}\|\Omega\|_{\infty},&\mbox{if }{R_{0}}\leq 1,\\ [2(2-{R_{0}})+\frac{4}{3\sqrt{3}}({R_{0}}-1)]\|\Omega\|_{\infty},&\mbox{if }1<{R_{0}}\leq 2.\end{cases}$
It is conceptually convenient to identify each $\theta_{j}$ with the corresponding point $e^{\mathrm{i}\theta_{j}}$ on the unit circle and consider (3) to be a dynamical system on $\mathbb{S}^{1}$. Then, the order parameter $R$ measures how far away and how close the points $e^{\mathrm{i}\theta_{j}}$ are from 1-1 and to $1$, respectively. Under this interpretation, Corollary 16 roughly says that if $\kappa\gtrsim{R_{0}}^{-3/2}\max_{i}|\omega_{i}|$, then (a) there is a uniform-in-time lower bound on the macroscopic quantity $R(t)$, (b) each particle $e^{\mathrm{i}\theta_{j}}$ will converge while not making a full revolution around the unit circle, and (c) any particle $e^{\mathrm{i}\theta_{j}}$ which is far away from 1-1 will be pulled towards and trapped near $1$. Thus, it is easier to constrain the behavior of (3) when $R_{0}$} is large, and when $R_{0}$} is small, we cannot say much about the long-term dynamics of (3) using Corollary 16 unless we pay the high price of $\kappa\gtrsim{R_{0}}^{-3/2}\max_{i}|\omega_{i}|$. Furthermore, if we are allowed to take $\kappa$ to be arbitrarily large, then (f) we can guarantee $R(t)$ to not go below its initial value by any prescribed amount, and (g) for any neighborhood of 1-1 in $\mathbb{S}^{1}$ we may bring any oscillator not in that neighborhood close to $1$ with arbitrary accuracy. All this happens while (e,i) the oscillators $e^{\mathrm{i}\theta_{j}}$ close to 1 are linearly ordered according to their natural frequencies $\omega_{j}$.
The threshold $K_{c}\max_{i=1,\cdots,N}|\omega_{i}|$ is far from sharp in light of Theorem 9; rather, we conjecture that $K_{c}$ can be taken to be a universal constant (Conjecture 10). However, the asymptotic dynamics described by Corollary 16 are sharp up to universal constants; see Theorem 34 and Remark 35.
Theorems 1 and 6 are derived from Corollary 16. Intuitively, the derivation is as follows: for most initial data $\{\theta_{i}^{0}\}_{i=1}^{N}$, the order parameter $R_{0}$} is close to $1$ by a simple Chernoff bound (this explains the factor $\exp\left(-\frac{\varepsilon^{2}N}{25}\right)$ in Theorem 6), and that even for initial data $R_{0}$ much smaller than $1$ the evolution of (3) will lead to an increase of $R_{0}$ close towards $1$ by a volumetric argument involving the divergence (this explains the factor $\left(\sqrt{\frac{\pi e}{2}}\frac{\|\Omega\|_{\infty}^{1/3}}{\kappa^{1/3}}\right)^{N}$ in Theorem 6). Hence we need only take $\kappa$ large enough to cover the case $R_{0}=1$, i.e., we need only $\kappa>2\max_{i}|\omega_{i}|$ in Theorems 1 and 6.
Although the statement of Theorem 6 treats $2\max_{i}|\omega_{i}|$ as if it were a critical threshold for $\kappa$, due to the above argument of inserting $R_{0}=1$ into Corollary 16, this is likely an artifact of our proof, as the phase diagram of [AS01] indicates that complete oscillator death happens even for smaller values of $\kappa$. The value $2\max_{i}|\omega_{i}|$ for $\kappa$ comes from the case $R_{0}=1$ of (8), but note that in the endpoint case $R_{0}=2$ of (8), the value $\frac{4}{3\sqrt{3}}\|\Omega\|_{\infty}$ for $\kappa$ is the maximum of $\kappa_{c}(\Omega)$ ranging over all $\Omega$ with a fixed $\|\Omega\|_{\infty}$. In other words, for a fixed $\|\Omega\|_{\infty}=\max_{i}|\omega_{i}|$, $\kappa_{\mathrm{crit}}\coloneqq\frac{4}{3\sqrt{3}}\|\Omega\|_{\infty}$ is a critical threshold for the existence of equilibria in the sense that if $\kappa>\kappa_{\mathrm{crit}}$ then (3), by [HPR15] (see Theorem 18 below), has an asymptotically stable equilibrium (i.e., an equilibrium with an open neighborhood such that for any initial data in that open set the solution to (3) converges to that equilibrium), while if $|\kappa|<\kappa_{\mathrm{crit}}$ and $\omega_{1}=\cdots=\omega_{N}=\|\Omega\|_{\infty}$ then, by [HKPR17, OKT17] not only does (3) fail to have any equilibria, but there is a positive measure of solutions to (3) satisfying $\theta_{i}(t)\to\infty$ as $t\to\infty$, $i=1,\cdots,N$.
Here is the theorem mentioned above.
Theorem 18 ([HPR15, special case of Theorem 2.2]).
For $\alpha\in\left(\frac{\pi}{3},\pi\right)$, let $\alpha^{\infty}\in\left(0,\frac{\pi}{3}\right)$ be the unique solution to $\sin\alpha^{\infty}(1+\cos\alpha^{\infty})=\sin\alpha(1+\cos\alpha)$.
  1. (a)
    (Existence and uniqueness of equilibrium) Then system (3) with parameters $\{\omega_{i}\}_{i=1}^{N}$ and $\kappa$ satisfying
    $\kappa>\frac{\|\Omega\|_{\infty}}{\sin\alpha(1+\cos\alpha)}$
    has a unique equilibrium $\Theta^{\infty}$ in $\alpha,\alpha]^{N}$. Furthermore, $\Theta^{\infty}\in(-\alpha^{\infty},\alpha^{\infty})^{N}$.
Moreover, let $\Theta(t)=\{\theta_{i}(t)\}_{i=1}^{N}$ be the solution to (3) with initial data $\{\theta_{i}^{0}\}_{i=1}^{N}$ such that $\theta_{i}^{0}\in[-\alpha,\alpha]$, $i=1,\cdots,N$. Then $\Theta(t)\to\Theta^{\infty}$ exponentially as $t\to\infty$. More precisely,
  1. (b)
    (Finite-time entrance into stable region) there exists a time $T\leq\frac{\pi}{\kappa\sin\alpha(1+\cos\alpha)-\|\Omega\|_{\infty}}$ such that $\Theta(t)\in(-\alpha^{\infty},\alpha^{\infty})^{N}$ for $t\geq T$, and we can take $T=0$ if $\Theta^{0}\in(-\alpha^{\infty},\alpha^{\infty})^{N}$; also,
  2. (c)
    (Exponential convergence to equilibrium) we have, for $q\geq 1$, that
    $\|\Theta(t)-\Theta^{\infty}\|_{\ell_{q}^{N}}\leq\|\Theta(T)-\Theta^{\infty}\|_{\ell_{q}^{N}}\exp\left[-\kappa(2\cos\alpha^{\infty}-1)(\cos\alpha^{\infty}+1)(t-T)\right],\quad t\geq T.$
In particular, if $\kappa>\frac{4}{3\sqrt{3}}\|\Omega\|_{\infty}$, we may find $\alpha\in(\frac{\pi}{3},\pi)$ such that $\kappa>\frac{\|\Omega\|_{\infty}}{\sin\alpha(1+\cos\alpha)}$, and Theorem 18 tells us that system (3) has a stable equilibrium in $(-\frac{\pi}{3},\frac{\pi}{3})^{N}$.
In light of this, the smallest constant $c$ we can take in Conjecture 10 is $c=\frac{4}{3\sqrt{3}}$, and Conjecture 11 (b) would imply this is attainable. Currently, Corollary 16 tells us that when $R_{0}=\Omega(1)$, $\kappa>\kappa_{\mathrm{crit}}+\Omega(2-R_{0})$ (which approaches the threshold $\kappa_{\mathrm{crit}}$ as $R_{0}$ approaches $2$) guarantees the oscillator death phenomenon. See subsubection 3.1.3 for a detailed discussion of possible future work in this direction.
2.2. The Łojasiewicz gradient theorem
Before we start discussing the results we have for general interaction functions $I$ and $S$, we should discuss the Łojasiewicz gradient theorem. One immediate advantage of specializing to the equation (3) is that the interaction functions $I$ and $S$ are real-analytic with $S=I^{\prime}$, so that (1) can be written as the gradient flow of a real analytic potential: if we define $V:\mathbb{R}^{N}\to\mathbb{R}$ as
$V(\Theta)\coloneqq-\sum_{i=1}^{N}\omega_{i}\theta_{i}-\frac{\kappa}{2N}\left[\sum_{i=1}^{N}I(\theta_{i})\right]^{2},$
then $V$ is a real analytic function and the ODE of (1) is equivalent to
$\dot{\Theta}=-\nabla_{\Theta}V(\Theta).$
The following is a special case of the well-known Łojasiewicz gradient flow theorem [Loj82], which in turn is a corollary of the Łojasiewicz gradient inequality [Loj63]:
Proposition 19.
Let $\Theta=\Theta(t)$ be the solution to (1) with parameters $\{\omega_{i}\}_{i=1}^{N}$ and $\kappa$, initial data $\{\theta_{i}^{0}\}_{i=1}^{N}$, and real-analytic interaction functions $I$ and $S$ such that $S=I^{\prime}$. If
$\sup_{t\geq 0}\|\Theta(t)\|_{\infty}<\infty,$
then there exists $\Theta^{\infty}\in\mathbb{R}^{N}$ such that
$\lim_{t\to\infty}\|\Theta(t)-\Theta^{\infty}\|_{\infty}=0\quad and\quad\lim_{t\to\infty}\|\dot{\Theta}(t)\|_{\infty}=0.$
Remark 20.
  1. (a)
    Proposition 19 was inspired by [DX13, HLX13] who applied the Łojasiewicz gradient theorem to the Kuramoto model, which is closely related to the Winfree model (see subsection 3.2) and whose gradient flow formulation was given by [VHW93]. The gradient flow formulation of (3) and consequently Proposition 19 is likely well-known, but we could not find them explicitly stated in the literature.
  2. (b)
    To know about the speed of the convergence toward the equilibrium guaranteed by the Łojasiewicz gradient theorem, one needs to know the exact exponent in the Łojasiewicz gradient inequality, which tells us whether the convergence is exponential or algebraic [BDL07, Theorem 4.7]. In [LXY15], the Łojasiewicz exponent was computed for certain equilibria of the Kuramoto model; one might be able to compute the Łojasiewicz exponent for the Winfree model (3) using similar methods.
Thus, it follows from Proposition 19 that statement (b) of Corollary 16 is equivalent to the weaker statement
$\sup_{t\geq 0}\theta_{i}(t)-\inf_{t\geq 0}\theta_{i}(t)<\infty,\quad i=1,\cdots,N,$
so later on we will only prove the above weaker statement when proving statement (b) of Corollary 16. This is the reasoning behind Remark 13.
On the other hand, when the assumption that $I$ and $S$ are real analytic with $S=I^{\prime}$ is unavailable, then we will not be able to deduce the convergence of the system (1). That is why, in the next subsection, we will have to content ourselves with proving only boundedness of solutions.
2.3. General interaction functions
We have stated Theorems 1, 6, and 16 in the case of the special Winfree model (3) for concreteness. The proof methods are robust enough to allow for many other interaction functions $I$ and $S$ considered in the literature.
In the case where $I$ is everywhere strictly positive and $S$ changes sign at least once, it will be straightforward to rigorously prove the oscillator death phenomenon for all initial data with a uniformly bounded $\kappa$; see Proposition 47. (This will serve as a toy model for our bootstrapping proof method in subsection 4.1.) Most of the interaction functions $I$ and $S$ considered in the literature are such that $I$ is nonnegative with a unique zero and $S$ changes sign around that zero, as in the standard choice (2).
Thus, except in subsection 4.1, we will assume that there exists an angle $\alpha_{0}\in(0,\pi)$, exponents $p,q\geq 1$, and constants $c_{1},c_{2}>0$ and $c_{3}\in(0,1]$ such that
[…]
$0\leq I(\theta)\leq c_{2}(\pi-|\theta|)^{q}\mbox{ for }\theta\in[-\pi,\pi],$ (10)
and
$\min_{|\phi|\leq\max\{|\theta|,\alpha_{0}\}}I(\phi)\geq c_{3}I(\theta)\mbox{ for }\theta\in[-\pi,\pi].$ (11)
Condition (10) says that $I$ has a zero at $\pi$ and has a power-type upper bound at $\pi$ of exponent $q$. Similarly, (9) says that $S$ has a zero at $\pi$ and has a power-type lower bound on its absolute value of exponent $p$ at $\pi$. Also, note that (11) is satisfied with $c_{3}=\frac{I(\alpha_{0})}{I(0)}$ if $I$ is an even function strictly positive on $(-\pi,\pi)$ and is monotone decreasing on $0,\pi]$; thus we may consider (11) to be a “weak monotonicity condition” on $I$. There is nothing special about the location of the zero $\pi$; it can be at any other location, and as long as we have power-type behavior as in (9) and (10) and the weak monotonicity condition (11) around that zero, the argument of this paper will follow through.
These assumptions are robust under various modifications one may wish to make for $S$ and $I$. For example, in the prototypical case $S(\theta)=-\sin\theta$ and $I(\theta)=1+\cos\theta$, one would take $p=1$, $q=2$, $\alpha_{0}=\frac{\pi}{2}$, and $c_{1}$, $c_{2}$, and $c_{3}$ to be universal constants. In [AS01] one also considers the examples $I_{n}(\theta)=(1+\cos\theta)^{n}$ for $n\geq 1$, which satisfies (10) with $q=2n$ and (11). Gallego, Montbrió, and Pazó [GMP17] consider the “rectified Poisson kernel” $I_{r}(\theta)=\frac{(1-r)(1+\cos\theta)}{1-2r\cos\theta+r^{2}}$, $r\in(-1,1)$, which satisfies (10) with $q=2$ and (11), and many functions $I$ that are compactly supported in $(-\pi,\pi)$, which satisfy (10) with any $q\geq 1$. Likewise, many natural kernels that approximate $\delta_{\{0\}}$ and which have a zero at $\pi$ satisfy (10) and (11).
As in the prototypical case, we introduce an average influence:
$R\coloneqq\frac{1}{N}\sum_{j=1}^{N}I(\theta_{j}),$ (12)
with which the ordinary differential equation of (1) can be rewritten in the mean-field form
$\dot{\theta}_{i}=\omega_{i}+\kappa RS(\theta_{i}).$ (13)
The analogue of Corollary 16 for the general model (1) is the following theorem.
Theorem 21.
Assume that the $2\pi$-periodic Lipschitz functions $S,I:\mathbb{R}\to\mathbb{R}$ satisfy the constraints (9), (10), and (11). Let $\{\theta_{i}(t)\}_{i=1}^{N}$ be the solution to (1) with parameters $\{\omega_{i}\}_{i=1}^{N}$, $\kappa$, and $\{\theta_{i}^{0}\}_{i=1}^{N}$, and interaction functions $S$ and $I$. Suppose ${R_{0}}\coloneqq R(0)>0$ and
$\kappa>\max\left\{\frac{2(2c_{2})^{p/q}}{c_{1}c_{3}}\cdot\frac{\|\Omega\|_{\infty}}{{R_{0}}^{1+\frac{p}{q}}},\frac{2}{c_{1}c_{3}(\pi-\alpha_{0})^{p}}\cdot\frac{\|\Omega\|_{\infty}}{{R_{0}}}\right\}.$ (14)
Then
  1. (a)
    $R(t)>\frac{c_{3}{R_{0}}}{2}$ for all $t\geq 0$, and
  2. (b)
    $\sup_{t\geq 0}\theta_{i}(t)-\inf_{t\geq 0}\theta_{i}(t)<2\pi$ for all $i=1,\cdots,N$.
Remark 22.
By Proposition 19, we can deduce further that if $S$ and $I$ are real-analytic with $S=I^{\prime}$, then for all $i=1,\cdots,N$, the limit $\lim_{t\to\infty}\theta_{i}(t)$ exists and $\lim_{t\to\infty}\dot{\theta}_{i}(t)=0$. We will omit this type of observation in the following discussion.
This theorem tells us that as long as $\kappa\gtrsim\frac{\max_{i}|\omega_{i}|}{{R_{0}}^{1+\frac{p}{q}}}$ we can guarantee boundedness of the solution $\{\theta_{i}(t)\}_{i=1}^{N}$. This theorem also tells us why we get the exponent $\frac{3}{2}$ in Corollary 16, namely that it is because $p=1$ and $q=2$ in that case.
The reason why we have stated Corollary 16 separately from Theorem 21 was to exhibit that it is possible to optimize the proof so as to obtain reasonable constants in the expression for $\kappa$, namely $2\cdot\frac{\max_{i}|\omega_{i}|}{{R_{0}}^{3/2}}$ for $0<R_{0}\leq 1$; that being said, the focus of Theorem 21 is to obtain the power type dependence on $R_{0}$} without caring much about the constants.
The analogue of Theorem 18 for general interaction functions $I$ and $S$ is given as follows.
Theorem 23 ([HPR15, Theorem 2.2]).
Consider the setting of the Winfree model (1). Assume that the interaction functions $S$ and $I$ are $2\pi$-periodic and $C^{2}$, that $S$ is odd and $I$ is even, and that there are angles $0<\theta_{*}<\theta^{*}\leq\pi$ such that
$S\leq 0\mbox{ on }[0,\theta^{*}],\quad S^{\prime}\leq 0,~S^{\prime\prime}\geq 0\mbox{ on }[0,\theta_{*}],$
[…]
Fix $\alpha\in(0,\theta^{*})$, and let $\alpha^{\infty}$ be the unique solution to $(SI)(\alpha^{\infty})=(SI)(\alpha)$, $\alpha^{\infty}\in(0,\theta_{*}]$. Assume
$\kappa>-\frac{\|\Omega\|_{\infty}}{S(\alpha)I(\alpha)}.$
Then the system (3) has a unique equilibrium in the set $\alpha,\alpha]^{N}$, with the equilibrium lying in the smaller open set $(-\alpha^{\infty},\alpha^{\infty})^{N}$, and for any initial data $\Theta^{0}\in[-\alpha,\alpha]^{N}$ in that open set, the solution $\Theta=\Theta(t)$ to (1) converges exponentially to that unique equilibrium.
The conditions on $S$ and $I$ roughly mean that $S$ and $I$ mimic (2).
Given the differences between the special case (2) and the more general case of (9), (10), and (11), we can redefine the critical and pathwise critical coupling strengths as follows.
Definition 24.
Fix a frequency vector $\Omega=\{\omega_{1},\cdots,\omega_{N}\}\in\mathbb{R}^{N}$ and $2\pi$-periodic Lipschitz interaction functions $I$ and $S$.
  1. (a)
    The critical coupling strength is
[…]
  2. (b)
    For an initial phase vector $\Theta^{0}=\{\theta^{0}_{1},\cdots,\theta^{0}_{N}\}\in\mathbb{R}^{N}$, the pathwise critical coupling strength is
[…]
Then Theorem 21 can be summarized as follows.
Theorem 25.
For a fixed initial position vector $\Theta^{0}$ with $R_{0}>0$, a natural frequency vector $\Omega$, and $2\pi$-periodic Lipschitz interaction functions $I$ and $S$ satisfying (9), (10), and (11), we have
$\kappa_{\mathrm{pc}}(\Theta^{0},\Omega,I,S)\leq\max\left\{\frac{2(2c_{2})^{p/q}}{c_{1}c_{3}}\cdot\frac{\|\Omega\|_{\infty}}{{R_{0}}^{1+\frac{p}{q}}},\frac{2}{c_{1}c_{3}(\pi-\alpha_{0})^{p}}\cdot\frac{\|\Omega\|_{\infty}}{{R_{0}}}\right\}.$
One can deduce the following theorem from Theorem 21 similarly to the deduction of Theorem 6 from Corollary 16. Again, it is based on the intuition that if we are sampling the initial data $\theta_{i}^{0}$ independently from a fixed distribution on $\pi,\pi]$, then by a simple Chernoff bound, the initial average influence $R_{0}$ will be not too far from its expected value.
Theorem 26.
Assume we are given parameters $\{\omega_{i}\}_{i=1}^{N}$ and $\kappa$, and $2\pi$-periodic Lipschitz interaction functions $S$ and $I$ that satisfy the constraints (9), (10), and (11). For each initial data $\{\theta_{i}^{0}\}_{i=1}^{N}$, denote by $\{\theta_{i}(t)\}_{i=1}^{N}$ the solution to (1).
Let $\mu$ be a probability measure on $\pi,\pi]$, and denote and assume
$R^{*}\coloneqq\int_{-\pi}^{\pi}I(\theta)d\mu(\theta)>0.$ (15)
Assume $\kappa$ is large enough so that
$\kappa>\max\left\{\frac{4(4c_{2})^{p/q}}{c_{1}c_{3}}\cdot\frac{\|\Omega\|_{\infty}}{(R^{*})^{1+\frac{p}{q}}},\frac{4}{c_{1}c_{3}(\pi-\alpha_{0})^{p}}\cdot\frac{\|\Omega\|_{\infty}}{R^{*}}\right\}.$ (16)
Then
$\mu^{\otimes N}\Big\{\{\theta_{i}^{0}\}_{i=1}^{N}\in[-\pi,\pi]^{N}:R(t)>\frac{c_{3}R^{*}}{4}~\forall t\geq 0\mbox{ and}$
$\qquad\qquad\qquad\qquad\qquad\qquad\sup_{t\geq 0}\theta_{i}(t)-\inf_{t\geq 0}\theta_{i}(t)<2\pi~\forall i=1,\cdots,N\Big\}$
$\geq 1-\exp\left(-\frac{(R^{*})^{2}N}{2\|I\|_{\sup}^{2}}\right),$
where $\mu^{\otimes N}$ denotes the product measure on $\pi,\pi]^{N}$.
In words, given a distribution $\mu$ satisfying the non-degeneracy condition (15), if the coupling strength is large as in (16) and if one samples $\theta_{i}^{0}\sim\mu$ independently for each $i$ and simulates (1), then, with high probability, each $\theta_{i}$ will fail to make a full revolution around the unit circle, and we will have a uniform lower bound on the macroscopic quantity $R$. We again consider this to be a near confirmation of the existence of the oscillator death regime for the interaction functions $I$ and $S$.
Under more assumptions, we can prove an improvement as $\kappa$ becomes large.
Theorem 27.
Assume we are given parameters $\{\omega_{i}\}_{i=1}^{N}$ and $\kappa$, and $2\pi$-periodic Lipschitz interaction functions $S$ and $I$ that satisfy the constraints (9), (10), and (11). In addition, let $\mu$ be a probability measure on $\pi,\pi]$ under which the function $I$ satisfies
$\int_{-\pi}^{\pi}e^{-\lambda I(\theta)}d\mu(\theta)\leq\frac{C_{\mu}}{\lambda^{\beta}},\quad\lambda>0,$ (17)
for some $C_{\mu},\beta>0$. For each initial data $\{\theta_{i}^{0}\}_{i=1}^{N}$, denote by $\{\theta_{i}(t)\}_{i=1}^{N}$ the solution to (1).
Assume $\kappa$ is large enough so that
$\kappa>\max\left\{\frac{2(2c_{2})^{p/q}}{c_{1}c_{3}}\left(\frac{e}{\beta}\right)^{1+\frac{p}{q}}C_{\mu}^{\frac{1}{\beta}(1+\frac{p}{q})}\cdot\|\Omega\|_{\infty},\frac{2eC_{\mu}^{1/\beta}}{c_{1}c_{3}(\pi-\alpha_{0})^{p}\beta}\cdot\frac{\|\Omega\|_{\infty}}{{R_{0}}}\right\}.$ (18)
Then
$\mu^{\otimes N}\Big\{\{\theta_{i}^{0}\}_{i=1}^{N}\in[-\pi,\pi]^{N}:\sup_{t\geq 0}\theta_{i}(t)-\inf_{t\geq 0}\theta_{i}(t)<2\pi~\forall i=1,\cdots,N\Big\}$
$\geq 1-C_{\mu}^{N}\left(e/\beta\right)^{\beta N}\max\left\{\left(\frac{2}{c_{1}c_{3}}\frac{\|\Omega\|_{\infty}}{\kappa}\right)^{\frac{1}{1+p/q}}(2c_{2})^{\frac{p/q}{1+p/q}},\frac{2}{c_{1}c_{3}(\pi-\alpha_{0})^{p}}\frac{\|\Omega\|_{\infty}}{\kappa}\right\}^{\beta N},$
where $\mu^{\otimes N}$ denotes the product measure on $\pi,\pi]^{N}$.
Condition (17) roughly means that when $\theta\sim\mu$, $I(\theta)$ is not too concentrated near 0.
Under stronger conditions on $I$ and $S$, we can take advantage of the time flow and can deduce the following uniform oscillator death result which is an analogue of Theorem 1.
Theorem 28.
Assume that the $2\pi$-periodic Lipschitz functions $S,I:\mathbb{R}\to\mathbb{R}$ satisfy the constraints (9), (10), and (11), along with the additional constraints
$S^{\prime}(\theta)\geq c_{4}(I_{*}-I(\theta)),\quad\theta\in\mathbb{R},$ (19)
for some $I_{*}>0$ and $c_{4}>0$,
$I^{\prime}(\theta)S(\theta)\geq 0,\quad\theta\in\mathbb{R},$ (20)
and the nondegeneracy condition
$I(\theta)>0,\quad\forall\theta\in(-\pi,\pi).$ (21)
Assume
$\kappa>\max\left\{\frac{2(2c_{2})^{p/q}}{c_{1}c_{3}}\cdot\frac{\|\Omega\|_{\infty}}{{I_{*}}^{1+\frac{p}{q}}},\frac{2}{c_{1}c_{3}(\pi-\alpha_{0})^{p}}\cdot\frac{\|\Omega\|_{\infty}}{{I_{*}}}\right\}.$
Then for Lebesgue almost every initial data $\Theta^{0}$, the solution $\Theta=\Theta(t)$ to (1) satisfies the following assertions.
  1. (a)
    (Uniform lower bound on the average influence) $\liminf_{t\to\infty}R(t)>\frac{c_{3}I_{*}}{2}$.
  2. (b)
    (Oscillator death) $\limsup_{t\to\infty}\theta_{i}(t)-\liminf_{t\to\infty}\theta_{i}(t)<2\pi$ for all $i=1,\cdots,N$.
For example, the interaction functions considered above, namely $S(\theta)=-\sin\theta$ and $I(\theta)=I_{n}(\theta)$ or $I_{r}(\theta)$ where $I_{n}(\theta)=(1+\cos\theta)^{n}$ for $n\geq 1$ and $I_{r}(\theta)=\frac{(1-r)(1+\cos\theta)}{1-2r\cos\theta+r^{2}}$ for $r\in(-1,1)$, all satisfy (9), (10), (11), (19), (20), and (21).
Theorem 28 can be summarized as follows.
Theorem 29.
For a fixed frequency vector $\Omega\in\mathbb{R}^{N}$ and $2\pi$-periodic Lipschitz interaction functions $I$ and $S$ satisfying (9), (10), (11), (19), (20), and (21), we have
$\kappa_{\mathrm{pc}}(\Theta^{0},\Omega,I,S)\leq\max\left\{\frac{2(2c_{2})^{p/q}}{c_{1}c_{3}}\cdot\frac{\|\Omega\|_{\infty}}{{I_{*}}^{1+\frac{p}{q}}},\frac{2}{c_{1}c_{3}(\pi-\alpha_{0})^{p}}\cdot\frac{\|\Omega\|_{\infty}}{{I_{*}}}\right\},\quad\mathrm{~for~a.e.~}\Theta^{0}=\{\theta^{0}_{1},\cdots,\theta^{0}_{N}\}\in\mathbb{R}^{N}.$
With a further assumption on $I$, we can prove the following quantitative improvement as time flows, which is an analogue of Theorem 7.
Theorem 30.
Assume we are given parameters $\{\omega_{i}\}_{i=1}^{N}$ and $\kappa$ and $2\pi$-periodic Lipschitz interaction functions $S$ and $I$ that satisfy the constraints (9), (10), (11), (19), (20), and the following strenghtening of (21):
$I(\theta)\geq c_{5}(\pi-|\theta|)^{r}\mbox{ for }\theta\in[-\pi,\pi],$ (22)
For each initial data $\{\theta_{i}^{0}\}_{i=1}^{N}$, denote by $\{\theta_{i}(t)\}_{i=1}^{N}$ the solution to (1).
If $\kappa$ is large enough so that
$\kappa>\max\left\{\frac{4(4c_{2})^{p/q}}{c_{1}c_{3}}\cdot\frac{\|\Omega\|_{\infty}}{{I_{*}}^{1+\frac{p}{q}}},\frac{4}{c_{1}c_{3}(\pi-\alpha_{0})^{p}}\cdot\frac{\|\Omega\|_{\infty}}{{I_{*}}}\right\},$ (23)
then
$m\left\{\{\theta_{i}^{0}\}_{i=1}^{N}\in[-\pi,\pi]^{N}:R(t)>\frac{c_{3}I_{*}}{4}~\forall t\geq T\mbox{ and }\sup_{t\geq T}\theta_{i}(t)-\inf_{t\geq T}\theta_{i}(t)<2\pi~\forall i=1,\cdots,N\right\}$
$\geq 1-\left(\exp\left(\frac{rI_{*}^{2}}{2\|I\|^{2}_{\sup}}\right)+\frac{c_{4}c_{5}\pi^{r}r^{r-1}}{\Gamma(1/r)^{r}(1+r/N)}\kappa N(I_{*}-\delta)T\right)^{-N/r}.$
If $\kappa$ is large enough so that
$\kappa>\max\left\{\frac{4e(4ec_{2})^{p/q}}{c_{1}c_{3}}\left(\frac{\Gamma(1/r)^{r}}{c_{5}\pi^{r}r^{r-1}}\right)^{1+p/q},\frac{4e\Gamma(1/r)^{r}}{c_{1}c_{3}c_{5}(\pi-\alpha_{0})^{p}\pi^{r}r^{r-1}}\right\}\|\Omega\|_{\infty},$ (24)
then
$m\left\{\{\theta_{i}^{0}\}_{i=1}^{N}\in[-\pi,\pi]^{N}:\sup_{t\geq T}\theta_{i}(t)-\inf_{t\geq T}\theta_{i}(t)<2\pi~\forall i=1,\cdots,N\right\}$
$\geq 1-\left(\frac{\Gamma(1/r)}{c_{5}^{1/r}\pi r^{1-1/r}}\right)^{N}\left(e^{-1}\min\left\{\left(\frac{c_{1}c_{3}}{2(2c_{2})^{p/q}}\right)^{\frac{1}{1+p/q}}\left(\frac{\kappa}{\|\Omega\|_{\infty}}\right)^{\frac{1}{1+p/q}},\frac{c_{1}c_{3}(\pi-\alpha_{0})^{p}}{2}\frac{\kappa}{\|\Omega\|_{\infty}}\right\}+\frac{c_{4}\kappa NI_{*}}{2(1+r/N)}T\right)^{-N/r}.$
2.4. Partial oscillator death
In the proofs of Theorems 15 and 21, we will be able to deduce statements about the behavior of subpopulations. For $\mathcal{B}\subseteq\{1,\cdots,N\}$ we denote
$\Omega_{\mathcal{B}}\coloneqq(\omega_{i})_{i\in\mathcal{B}},\quad\Theta_{\mathcal{B}}\coloneqq(\theta_{i})_{i\in\mathcal{B}}.$
Likewise, we use $\|\cdot\|_{\infty}$ to denote the $\ell^{\infty}$-norm:
$\|\Omega_{\mathcal{B}}\|_{\infty}\coloneqq\max_{i\in\mathcal{B}}|\omega_{i}|,\quad\|\Theta_{\mathcal{B}}\|_{\infty}\coloneqq\max_{i\in\mathcal{B}}|\theta_{i}|.$
Definition 31.
Let $\Theta=(\theta_{1},\cdots,\theta_{N})$ be a solution to the Winfree model (1). For $\mathcal{B}\subseteq\{1,\cdots,N\}$, we say that the ensemble $\Theta$ exhibits $\mathcal{B}$-partial oscillator death if the subensemble $\Theta_{\mathcal{B}}$ is bounded:
$\sup_{t\geq 0}\|\Theta_{\mathcal{B}}(t)\|_{\infty}<\infty.$
We will use $\mathcal{A}$ and $\mathcal{B}$ to denote subsets of $\{1,\cdots,N\}$. By convention, we will have $\mathcal{A}\subset\mathcal{B}$, where $\mathcal{A}$ will denote some collection of oscillators that we intend to trap within a proper subinterval $(-\pi+\delta,\pi-\delta)$ of $\pi,\pi]$, and $\mathcal{B}$ will denote a larger collection of oscillators for which we wish to prove $\mathcal{B}$-partial oscillator death.
Our results on partial oscillator death are stated as follows.
Proposition 32 (Criterion for partial oscillator death in (3)).
Let $\Theta=\Theta(t)$ be the solution to (3) with parameters $\{\omega_{i}\}_{i=1}^{N}$, $\kappa$, and $\{\theta_{i}^{0}\}_{i=1}^{N}$. Suppose $0<\rho\leq 2$ satisfies
$\kappa>\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}}{\rho},$ (25)
and $\mathcal{A}\subseteq\mathcal{B}\subseteq\{1,\cdots,N\}$ satisfy
$\frac{1}{N}\sum_{i\in\mathcal{A}}(1+\cos\theta_{i}^{0})>\frac{2\rho}{1+\sqrt{1-\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}^{2}}{\kappa^{2}\rho^{2}}}}$ (26)
and
$\cos\theta_{i}^{0}\geq-\sqrt{1-\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}^{2}}{\kappa^{2}\rho^{2}}},\quad\forall i\in\mathcal{A}.$ (27)
Then
  1. (a)
    $R(t)\geq\rho$ for all $t\geq 0$,
  2. (b)
    $\cos\theta_{i}(t)\geq-\sqrt{1-\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}^{2}}{\kappa^{2}\rho^{2}}}$ for all $i\in\mathcal{A}$ and $t\geq 0$,
  3. (c)
    Let $i\in\mathcal{B}$, $0<\alpha<\sqrt{1-\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}^{2}}{\kappa^{2}\rho^{2}}}$, and suppose there is a time $t_{0}\geq 0$ such that $\cos\theta_{i}(t_{0})\geq-\alpha$. Then $\cos\theta_{i}(t)\geq-\alpha$ for $t\geq t_{0}$, and $\cos\theta_{i}(t)\geq\alpha$ for $t\geq t_{0}+\frac{\pi}{\kappa\rho\sqrt{1-\alpha^{2}}-{\|\Omega_{\mathcal{B}}\|_{\infty}}}$. In particular, $\liminf_{t\to\infty}\cos\theta_{i}(t)\geq\sqrt{1-\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}^{2}}{\kappa^{2}\rho^{2}}}$.
  4. (d)
    For all $i\in\mathcal{B}$,
    $\sup_{t\geq 0}\theta_{i}(t)-\inf_{t\geq 0}\theta_{i}(t)<2\pi.$
  5. (e)
    Let $i,j\in\mathcal{B}$, $0<\alpha<\sqrt{1-\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}^{2}}{\kappa^{2}\rho^{2}}}$, and suppose there is a time $t_{0}\geq 0$ such that $\cos\theta_{i}(t_{0})\geq-\alpha$ and $\cos\theta_{j}(t_{0})\geq-\alpha$. If $\omega_{i}>\omega_{j}$, then after modulo $2\pi$ translations,
    $\theta_{i}(t)-\theta_{j}(t)\leq\frac{\omega_{i}-\omega_{j}}{\kappa\rho\alpha}+\pi\exp\left(-\kappa\rho\alpha\left(t-t_{0}-\frac{\pi}{\kappa\rho\sqrt{1-\alpha^{2}}-{\|\Omega_{\mathcal{B}}\|_{\infty}}}\right)\right)\mbox{ for }t\geq t_{0}+\frac{\pi}{\kappa\rho\sqrt{1-\alpha^{2}}-{\|\Omega_{\mathcal{B}}\|_{\infty}}},$
    and
    $\theta_{i}(t)-\theta_{j}(t)\geq\frac{\omega_{i}-\omega_{j}}{2\kappa}-\frac{\omega_{i}-\omega_{j}}{2\kappa}\exp\left(-2\kappa\left(t-t_{0}-\frac{\pi}{\kappa\rho\sqrt{1-\alpha^{2}}-{\|\Omega_{\mathcal{B}}\|_{\infty}}}-\frac{\pi}{\omega_{i}-\omega_{j}}\right)\right)$
    $\mbox{for }t\geq t_{0}+\frac{\pi}{\kappa\rho\sqrt{1-\alpha^{2}}-{\|\Omega_{\mathcal{B}}\|_{\infty}}}+\frac{\pi}{\omega_{i}-\omega_{j}}.$
    If $\omega_{i}=\omega_{j}$, then after modulo $2\pi$ translations, $\lim_{t\to\infty}\left(\theta_{i}(t)-\theta_{j}(t)\right)=0$ exponentially:
    $|\theta_{i}(t)-\theta_{j}(t)|\leq\pi\exp\left(-\kappa\rho\alpha\left(t-t_{0}-\frac{\pi}{\kappa\rho\sqrt{1-\alpha^{2}}-{\|\Omega_{\mathcal{B}}\|_{\infty}}}\right)\right)\mbox{ for }t\geq t_{0}+\frac{\pi}{\kappa\rho\sqrt{1-\alpha^{2}}-{\|\Omega_{\mathcal{B}}\|_{\infty}}}.$
  6. (f)
    If $\mathcal{B}=\{1,\cdots,N\}$, then for all $i=1,\cdots,N$, the limit $\lim_{t\to\infty}\theta_{i}(t)$ exists, and
    $\lim_{t\to\infty}\dot{\theta}_{i}(t)=0.$
Proposition 33 (Criterion for partial oscillator death in (1)).
Let $\Theta=\Theta(t)$ be the solution to (1) with parameters $\{\omega_{i}\}_{i=1}^{N}$, $\kappa$, and initial data $\{\theta_{i}^{0}\}_{i=1}^{N}$, and let the $2\pi$-periodic Lipschitz functions $I$ and $S$ satisfy (9), (10), and (11). Suppose $0<\rho\leq\|I\|_{\mathrm{sup}}$ and $\mathcal{A}\subset\mathcal{B}\subset\{1,\cdots,N\}$ satisfy
$\frac{1}{N}\sum_{i\in\mathcal{A}}I(\theta_{i}^{0})>\frac{\rho}{c_{3}},\quad\kappa>\frac{\|\Omega_{\mathcal{B}}\|_{\infty}}{\rho c_{1}(\pi-\alpha_{0})^{p}}$ (28)
and
$\theta_{i}^{0}\in\left[-\pi+\left(\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}}{\kappa\rho c_{1}}\right)^{1/p},\pi-\left(\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}}{\kappa\rho c_{1}}\right)^{1/p}\right],\quad\forall i\in\mathcal{A}.$ (29)
Then
  1. (a)
    $R(t)>\rho$ for $t\in[0,\infty)$,
  2. (b)
    $\theta_{i}(t)\in[-\pi+\left(\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}}{\kappa\rho c_{1}}\right)^{1/p},\pi-\left(\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}}{\kappa\rho c_{1}}\right)^{1/p}]$ for $i\in\mathcal{A}$ and $t\geq 0$,
  3. (c)
    $\sup_{t\geq 0}\theta_{i}(t)-\inf_{t\geq 0}\theta_{i}(t)<2\pi$ for $i\in\mathcal{B}$.
The proofs of Propositions 32 and $3$ will be based on a bootstrapping method to prove a global lower bound on the order parameter or average influence $R$. We will prove Theorems 16 and 21 by seeing that the conditions of Propositions 32 and $3$ are satisfied with $\mathcal{B}=\{1,\cdots,N\}$ for a suitable choice of $\mathcal{A}$.
2.5. Equilibria of the prototypical Winfree model
The asymptotic dynamics described in Corollary 16 are optimal in the sense that for any $0<\rho\leq 2$ there exists an equilibrium state of (3) whose order parameter is roughly $\rho$.
Theorem 34.
Fix any $\rho\in(0,2]$. For any $N\geq 1$, $\kappa\neq 0$ and $\{\omega_{i}\}_{i=1}^{N}$ such that
$N\geq\frac{2}{\rho},\quad\frac{\max_{i}|\omega_{i}|}{|\kappa|}<\frac{\rho^{1.5}}{16},$
there exists an equilibrium initial data $\{\theta_{i}^{0}\}_{i=1}^{N}$ to (3) with order parameter $\in[\frac{1}{4}\rho,\frac{3}{2}\rho]$.
Moreover, we can take $\{\theta_{i}^{0}\}_{i=1}^{N}$ so that there exists $1\leq m\leq N$ with $\frac{\rho}{4}<\frac{m}{N}\leq\frac{\rho}{2}$ such that
$\frac{2|\omega_{i}|}{3\rho|\kappa|}\leq|\theta_{i}^{0}|\leq\frac{2\pi|\omega_{i}|}{\rho|\kappa|},\quad\mbox{for }i=1,\cdots,m,$
and
$\frac{2|\omega_{i}|}{3\rho|\kappa|}\leq|\theta_{i}^{0}-\pi|\leq\frac{2\pi|\omega_{i}|}{\rho|\kappa|},\quad\mbox{for }i=m+1,\cdots,N.$
Remark 35.
By Theorem 34, Corollary 16 is optimal in the sense that even if we raise $\kappa$ to arbitrary high values, there exist configurations such that (a) we cannot make the limiting value of $R(t)$ to be much larger than $R_{0}$, (c) we cannot bring closer to 0 the oscillators initially close to $\pi$, which may make up a certain proportion of the oscillators, and (d) we cannot make the other oscillators closer to 0 than the order of $\frac{|\omega_{i}|}{\kappa R_{0}}$. Furthermore, it is not hard to construct fixed inital configurations that exhibit the aforementioned properties (a), (c), and (d) as we raise $\kappa$ to higher values; simply take a set $\mathcal{C}\subset\{1,\cdots,N\}$ large enough with
$\theta_{i}^{0}=\pi,~\omega_{i}=0,\quad i\in\mathcal{C},$
so that $\theta_{i}(t)=\pi$ for $t\geq 0$ and $i\in\mathcal{C}$, and $\sup_{t\geq 0}R(t)\leq\frac{2(N-|\mathcal{C}|)}{N}$. It is not clear to us what happens when $\omega_{i}\neq 0$ for $i\in\mathcal{C}$; understanding this seems to be the main challenge in understanding oscillator death for generic initial data. We remark that for a fixed generic initial data where $\theta_{i}^{0}\not\equiv\pi\mod 2\pi$, as we raise $\kappa$ to higher values more and more oscillators will satisfy the conditions of Corollary 16 (e) and thus will satisfy $\liminf_{\kappa\to\infty}\liminf_{t\to\infty}\cos\theta_{i}(t)=1$.
In the proof of Theorem 34 we will analyze the basic structure of the equilibria of (3). From this we will construct a polynomial $W(r)$ of degree at most $2^{N+1}$ with the property that if $r_{0}$ is a root of $W(r)$ of order $m$, then there can only be at most $m$ distinct equilibria of (3) with order parameter $r_{0}$, where we say that two equilibria are distinct if they are distinct modulo $2\pi$. From this we obtain the following theorem.
Theorem 36.
For fixed data $(N,\{\omega_{i}\}_{i=1}^{N},\kappa)$ with $\kappa\neq 0$, there are at most $2^{N+1}$ distinct (modulo $2\pi$) equilibria of the system (3).
For ease of reference, we will call the above proof method a polynomial description of the equilibria of (3).
3. Previous work and directions for possible future research
In this section, we discuss previous works on the Winfree model (1) from the standpoint of dynamical systems of ordinary differential equations, and highlight the context of our theorems and their significance. We then relate the Winfree model to the Kuramoto model, in which Ha and the author [HR20] obtained results analogous to Theorem 1 and Corollary 16. During these discussions, we will suggest some questions for future research.
Those who are interested only in the proof of our results may skip to the next section.
3.1. Oscillator death versus oscillator locking
3.1.1. Winfree’s original work: a perturbative linear periodic system
In 1967, Winfree [Win67] proposed the mathematical model (1) to describe the behavior of weakly coupled biological oscillators such as the pacemaker cells in the human heart [^3] Under this interpretation, oscillator death can be viewed as cardiac arrest. or the synchronous flashing of fireflies [BB66]; see [PRK01, Win02] for more real-world applications. The modeling process can be described as follows. Individually, under the absence of other oscillators, each ($2\pi$-periodic) oscillator $\theta_{i}$ would undergo uniform motion at the rate of its intrinsic frequency $\omega_{i}$. Under the presence of other oscillators, however, each oscillator feels a frequency perturbation determined by the other oscillators’ phases. The interaction between the oscillators happens via a mean-field $R$ where each oscillator contributes by the amount $I(\theta)$ determined by its phase $\theta$, and the mean-field $R$ is given as the average of all contributions:
$R(\Theta)=\frac{1}{N}\sum_{j=1}^{N}I(\theta_{j}).$
The strength of the mean-field is denoted by the multiplicative factor of the coupling strength $\kappa$. Each oscillator reacts to the mean-field with sensitivity $S(\theta)$ depending on its phase $\theta$. From this we derive the Winfree model (1). One could modify (1) to incorporate more realistic behavior. For example, one could consider different connection networks (other than the all-to-all uniform connection of (1)), time-delayed interactions, or frustrations, which could arise from the attenuation of the influence $I$ due to some spatial distribution of the oscillators. We will not pursue these directions in this paper.
Winfree considers (1) to be a perturbation of a periodic system, by assuming that the frequencies $\omega_{i}$ are concentrated around a positive value $\omega_{\mathrm{average}}$, $\kappa>0$ is very small compared to $\omega_{\mathrm{average}}$, and the phases $\theta_{i}$ are also concentrated around a moving average. Winfree informally takes the infinite $N$ limit to assume the phases and frequencies have a distribution. Under these assumptions, he describes the structure of the solution as a perturbation of the periodic solution $\theta_{i}(t)=\omega_{\mathrm{average}}t$ and computes some macroscopic quantities associated to the solution.
3.1.2. Ariaratnam and Strogatz: rich nonlinear dynamics
Ariaratnam and Strogatz [AS01] realized that the Winfree model in the form of the nonlinear ordinary differential equation of (1) is much more versatile than Winfree’s intention as a linear perturbation of a periodic system. As mentioned in the first section, after fixing $\kappa>0$ and $\gamma\in[0,1]$, they (nonrandomly) sampled $\omega_{i}$ according to the uniform distribution on $1-\gamma,1+\gamma]$ and numerically simulated the prototypical model (3). They classified the asymptotic behavior of the model based on the rotation numbers $\rho_{i}=\lim_{t\to\infty}\theta_{i}(t)/t$, and observed that the shape of the distribution of the $\rho_{i}$’s did not depend on the initial data $\theta_{i}^{0}$. Based on their data, they drew a phase diagram for different $\gamma$ and $\kappa$ depicting different asymptotic behavior. They observed not only a complete locking regime ($\rho_{i}=\rho>0$ for all $i$ for some $\rho$) for small $\kappa$ and $\gamma$, which corresponds to Winfree’s heuristics, but also a complete death regime ($\rho_{i}=0$ for all $i$) for large $\kappa$, a partial death regime ($\rho_{i}=0$ for some $i$) for large $\gamma$ and intermediate $\kappa$, a partial locking regime ($\rho_{i}=\rho$ for some $i$ for some $\rho$) for small $\kappa$ and intermediate $\gamma$, and an incoherence regime (other behavior) for small $\kappa$.
Intuitively, in the ordinary differential equation (5), there is a competition between the linear first-order force $\omega_{i}$ that drives the system towards periodic motion, and the nonlinear first-order interaction force $\kappa R\sin\theta_{i}$ that tries to keep each $\theta_{i}$ near 0 modulo $2\pi$. Winfree’s heuristics is that for small $\kappa>0$ and for $\omega_{i}$ concentrated near a positive quantity, the linear term $\omega_{i}$ dominates, while the nonlinear interaction term $\kappa R\sin\theta_{i}$ simply moderates the small difference of the $\omega_{i}$’s among the oscillators, so that the collection of the oscillators rotates with a common frequency. If the $\kappa$ is very large so that the nonlinear interaction term dominates, each oscillator would stay bounded near 0 and its intrinsic frequency $\omega_{i}$ would simply tell us how far away the oscillator stays away from 0. If the frequency distribution is wide enough compared to the magnitude of either $\kappa$ or the mean frequency, then different asymptotic behavior will happen among the oscillators, and we will observe either partial death or partial locking, or other incoherent behavior.
According to the phase diagram of [AS01], for a fixed distribution of oscillators, as one increases $\kappa$, the system (3) experiences phase transitions between the regimes such as
$\begin{aligned} \mbox{Incoherence}&\quad\Longrightarrow\quad\mbox{Partial death}\quad\Longrightarrow\quad\mbox{Complete death}\\ \mbox{Incoherence}&\quad\Longrightarrow\quad\mbox{Partial locking}\quad\Longrightarrow\quad\mbox{Complete death}\\ \mbox{Incoherence}&\quad\Longrightarrow\quad\mbox{Partial locking}\quad\Longrightarrow\quad\mbox{Complete locking}\quad\Longrightarrow\quad\mbox{Complete death}.\end{aligned}$ (30)
From this standpoint, this paper presents sufficient conditions for the transition towards partial death and complete death.
3.1.3. Rigorous nonlinear dynamics: oscillator death
Most of the well-known literature on the Winfree model seem to be either based on numerical simulations of the Winfree model or its variants such as [AS01, GMP17], or describe the infinite-$N$ limit [ALM+04, BU09, GMP17, QRS07, PG20] (there are also similar works on variants of the Winfree model, such as [HKM21a, OS16, PG20, GMN+07]). The numerical simulations are significant since they show that diverse asymptotic dynamics emerge for the Winfree model, while also showing that generally the asymptotic dynamics do not depend on the initial data. The works on the infinite-$N$ limit are significant since for example they tell us exactly where the phase transitions happen between the different regimes. However, there are only a handful of rigorous results in the finite-$N$ case. Ermentrout and Kopell [EK90] extensively studied the oscillator death phenomenon for $N=2$, while the works [HPR15, HKPR16, HKPR17, OKT17, OTK19], although treating arbitrarily large $N$, possess the limitation that they require the initial data to be well-controlled.
A rigorous analysis and proof of oscillator death in the finite-$N$ case was given by Ha, Park, and the author in [HPR15]; see Theorem 23. This work rigorously proves that for a fixed generic initial data the oscillator death phenomenon happens for large $\kappa$, and establishes a nonlinear stability statement for an equilibrium for (1) and identifies its basin of attraction.
We remark that Theorem 23 deduces convergence of the system (1) while only requiring $C^{2}$-regularity on $S$ and $I$, and not even requiring that $S=I^{\prime}$. This suggests that it might be possible to deduce the convergence statement of Theorems 1 and 16 without appealing to the Łojasiewicz gradient theorem, and that under suitable assumptions we might be able to extend this to general interaction functions.
Question 37.
Suppose $S$ and $I$ satisfy the conditions of Theorems 21 and 23, and possibly some other regularity conditions, but not necessarily real analyticity. Assume (14). Can we conclude the convergence of $\Theta(t)$?
The subsequent work [HKPR16] showed that a similar framework holds with general symmetric connectivity graphs. However, these works [HPR15, HKPR16] have the drawback that they require all of the initial data to lie on some proper subinterval of $\pi,\pi]$, on which the sufficient coupling strength $\kappa$ depends. Thus, when one is sampling $\theta_{i}^{0}$ uniformly randomly from the entire interval $\pi,\pi]$, the $\kappa$ typically required by Theorem 23 will grow arbitrarily large as $N\to\infty$.
The significance of our Theorems 16 and 21 is that they prove oscillator death while only requiring control on the order parameter. This is good news since the order parameter is a macroscopic quantity that is easier to control, as shown in Theorems 1, 6, 7, 26, 27, 28, and 30.
Although Corollary 16 and Theorem 21 fail to address the case of initial data with very small order parameter, the initial data $\Theta^{0}$ such that $R^{0}$ is close to 0 is unstable in the sense that the divergence of the flow (3) is positive. Thus, most solutions to (3) escape the small $R$ regime in finite time, after which we to apply Corollary 16 or Theorem 21 to obtain Theorems 1, 6, 7, 26, 27, 28, and 30, namely oscillator death for Lebesgue a.e. initial data.
Remark 38.
The reason we do not have a positive resolution of Conjecture 10 in this paper is that we use volumetric arguments to obtain Theorem 1. The following is a naïve argument using Corollary 16 which fails. If an $\Omega(1)$ proportion of the oscillators satisfy $|\omega_{i}|\gtrsim\|\Omega\|_{\infty}$, then we can see that there exists a time $t_{0}\geq 0$ at which $R(t_{0})\gtrsim\frac{\|\Omega\|_{\infty}}{\kappa}$. Indeed, otherwise $R(t)\lesssim\frac{\|\Omega\|_{\infty}}{\kappa}$ for $t\geq 0$ and an $\Omega(1)$ proportion of the oscillators satisfy $\operatorname{sgn}\omega_{i}\cdot\dot{\theta_{i}}=|\omega_{i}|+O(\|\Omega\|_{\infty})\gtrsim\|\Omega\|_{\infty}$, so that the time-average of $R(t)$ is $\Omega(1)$, a contradiction. However, the lower bound $R\gtrsim\frac{\|\Omega\|_{\infty}}{\kappa}$ falls short of the stronger requirement $R\gtrsim\left(\frac{\|\Omega\|_{\infty}}{\kappa}\right)^{2/3}$ to use Corollary 16.
Remark 39.
Conjecture 10 is true if we relax the condition that $c$ should be a universal constant. Indeed, we can take $c=\frac{4N}{3\sqrt{3}}$ (see Proposition 46), in which case it is easy to see that $\kappa>c\max_{i}|\omega_{i}|$ implies $\dot{\theta_{i}}\leq 0$ if $\theta_{i}\equiv\frac{\pi}{3}\mod 2\pi$ and $\dot{\theta_{i}}\geq 0$ if $\theta_{i}\equiv-\frac{\pi}{3}\mod 2\pi$, so that oscillator death occurs. Thus we know that such a $c$ exists; the question is whether it can be a universal constant independent of $N$.
3.1.4. Rigorous nonlinear dynamics: oscillator locking
There are also rigorous works [HKPR17, OKT17, OTK19] on oscillator locking, which verify Winfree’s original heuristics. These works prove the existence and nonlinear stability of traveling solutions to (3) in the finite-$N$ regime, when $\kappa$ is small and the intrinsic frequencies are tightly concentrated around a fixed positive value. To be precise, we consider the following definitions.
Definition 40.
Let $\Theta=(\theta_{1},\cdots,\theta_{N})$ be a solution to the Winfree model (1).
  1. (a)
    We say that the ensemble $\Theta$ exhibits (complete) oscillator locking if
    $\sup_{t\geq 0}|\theta_{i}(t)-\theta_{j}(t)|<\infty,\quad\forall i,j=1,\cdots,N,$
    and
    $\lim_{t\to\infty}\frac{\theta_{i}(t)}{t}\neq 0,\quad\forall i=1,\cdots,N,$
    where the latter statement does not necessarily assume the existence of the limit.
  2. (b)
    For $\mathcal{B}\subseteq\{1,\cdots,N\}$, we say that the ensemble $\Theta$ exhibits $\mathcal{B}$-partial oscillator locking if the subensemble $\Theta_{\mathcal{B}}$ exhibits oscillator locking:
    $\sup_{t\geq 0}|\theta_{i}(t)-\theta_{j}(t)|<\infty,\quad\forall i,j\in\mathcal{B},$
    and
    $\lim_{t\to\infty}\frac{\theta_{i}(t)}{t}\neq 0,\quad\forall i\in\mathcal{B}.$
The mechanism of oscillator locking in [OKT17, OTK19, HKPR17] is as follows. Suppose $\kappa$ is small, the $\omega_{i}$ are concentrated around a fixed positive quantity, say $\omega_{i}\in[1-\gamma,1+\gamma]$, and the $\theta_{i}^{0}$’s are concentrated on a small arc. Then the term $\omega_{i}\approx 1$ dominates in (1), and the entire population will rotate around the circle. One can compute how the infinitesimal distance between two oscillators change after one revolution, namely, if all the $\omega_{i}$’s are $1$, then the log of the ratio between the infinitesimal distances between $\theta_{i}$ and $\theta_{j}$ before and after one revolution is $\int_{0}^{2\pi}\frac{I(\theta)S^{\prime}(\theta)}{1+\kappa I(\theta)S(\theta)}d\theta$. Thus, if
$\int_{0}^{2\pi}\frac{I(\theta)S^{\prime}(\theta)}{1+\kappa I(\theta)S(\theta)}d\theta<0$ (31)
(this is called the “synchronization hypothesis” in [OKT17]), the oscillators will become more concentrated after each revolution, and hence will stay close together for all time, exhibiting oscillator locking. The point of the works [OKT17, OTK19, HKPR17] is to “nonlinearize” the above “linear” analysis when $\kappa$, $\gamma$, and the diameter of the initial data are small.
More precisely, [OKT17] states that if (31) holds, $\kappa$ and $\gamma$ are small, and the oscillators are initially close to each other, then they will stay close to each other for all time. In their sequel work [OTK19] they obtain more refined stability properties of the solution operator under the same hypotheses. On the other hand, the work [HKPR17] shows in the special case $S(\theta)=-\sin\theta$ and $I(\theta)=1+\cos\theta$ that if at least $\frac{4}{4+\pi}$ of the oscillators are concentrated on a small arc, then partial locking happens. It seems very likely that a version of this partial locking statement holds for general interaction functions $S$ and $I$ which satisfy (31).
Question 41 (Rigorous partial locking for general $S$ and $I$).
Suppose the $2\pi$-periodic Lipschitz functions $S$ and $I$ satisfy (31). Do there exist $\rho,\gamma\in(0,1]$, and $\kappa_{0},\Delta_{1},\Delta_{2}>0$ such that if $\mathcal{A}\subset\mathcal{B}\subset\{1,\cdots,N\}$, $|\mathcal{A}|/N\geq\rho$, $\omega_{i}\in[1-\gamma,1+\gamma]$ for all $i\in\mathcal{B}$, $\operatorname{diam}\{\theta_{i}^{0}\}_{i\in\mathcal{A}}<\Delta_{1}$, and $0<\kappa<\kappa_{0}$, then $\operatorname{diam}\{\theta_{i}(t)\}_{i\in\mathcal{A}}<\Delta_{2}$ for all $t\geq 0$ and the ensemble $\Theta$ exhibits $\mathcal{B}$-partial oscillator locking?
There still remains the question whether (partial) locking states may emerge from generic initial data. For this, Winfree [Win67] provides the following insight:
… if $I(\theta)$] adds constructively to an entraining rhythm which is the sum of all the other $I(\theta_{i})$], and if the spectral density of oscillators is sufficiently great, then any infinitesimal periodicity in [$R$] is able to make a few oscillators coherent, which then sufficiently augment the influence rhythm to entrain a few more of nearby frequency. Like spontaneous combustion, the growth of $R(\theta)$] is autocatalytic. The phase-locked nucleus expands along the frequency axis until it has engulfed a bandwidth under $N(f)$ that cannot further expand except by conscripting more oscillators to boost $|R(\phi)|]$ than are available in the rarified population just beyond its frequency axis boundaries… [Win67, p.27]
It would seem that new ideas are needed to elaborate on this intuition to show the emergence of (partial) locking from generic initial data.
Using the ideas of this paper along with the ideas from [OKT17, OTK19, HKPR17], it may be possible to rigorously demonstrate the emergence of a mixture of partial death and partial locking, described below.
Question 42 (Rigorous partial death with partial locking).
For simplicity, let $S$ and $I$ be given as in (2). Do there exist $\rho_{1},\rho_{2},\gamma\in(0,1]$ with $\rho_{1}+\rho_{2}\leq 1$, and $\kappa>0$ such that if $\mathcal{B},\mathcal{C}\subset\{1,\cdots,N\}$ are mutually disjoint subsets with $|\mathcal{B}|/N\geq\rho_{1}$, $|\mathcal{C}|/N\geq\rho_{2}$, $\omega_{i}\in[-\gamma,\gamma]$ for all $i\in\mathcal{B}$, and $\omega_{i}\in[1-\gamma,1+\gamma]$ for all $i\in\mathcal{C}$, then for some initial data $\Theta^{0}$, the solution $\Theta=\Theta(t)$ to (3) exhibits $\mathcal{B}$-partial oscillator death and $\mathcal{C}$-partial oscillator locking?
Here is one idea to answer Question 42 in the affirmative. As the ensemble $\mathcal{C}$ rotates around the unit circle at more or less the same speed, the average contribution of $\mathcal{C}$ to $R(t)$ will be positive. However, the proof of Corollary 16 tells us that a pointwise lower bound on $R$ can be exploited to prove oscillator death (see Lemmas 48 and 49). One might be able to adapt the methodology of this paper to exploit an average lower bound on $R$ (as opposed to a pointwise one) to show $\mathcal{B}$-partial oscillator death.
3.2. Analogy with the Kuramoto model
The Kuramoto model is the following initial value problem for the $N\geq 1$ variables $\theta_{1},\cdots,\theta_{N}\in\mathbb{R}$, with the same parameters $\omega_{1},\cdots,\omega_{N}\in\mathbb{R}$, $\kappa\in\mathbb{R}$, and $\theta_{1}^{0},\cdots,\theta_{N}^{0}\in\mathbb{R}$ as in (3):
$\begin{cases}\dot{\theta}_{i}(t)=\omega_{i}+\frac{\kappa}{N}\sum_{j=1}^{N}\sin(\theta_{j}(t)-\theta_{i}(t)),\quad t>0,\\ \theta_{i}(0)=\theta_{i}^{0},\end{cases}\quad i=1,\cdots,N.$ (32)
This model was proposed by Kuramoto in [Kur75, Kur03] as a variant of the Winfree model (3) that is more amenable to mathematical analysis, namely he was able to demonstrate a phase-transition phenomenon when increasing $\kappa$ from 0 to some finite value. This model has since found applications in Josephson junction arrays [WS94]; see [ABV+05, PR15, Str00] for an exposition on this topic.
Unlike the Winfree model, the Kuramoto model (32) has the conservation law
$\frac{d}{dt}\left[\sum_{i=1}^{N}\theta_{i}(t)-\left(\sum_{i=1}^{N}\omega_{i}\right)t\right]=0,$
and admits a corresponding $\mathrm{U}(1)$-symmetry, namely (32) is invariant under the Galilean transformation
$\omega_{i}\mapsto\omega_{i}-\nu,\quad\theta_{i}^{0}\mapsto\theta_{i}^{0}-\vartheta,\quad\theta_{i}(t)\mapsto\theta_{i}(t)-\nu t-\vartheta,$
for fixed $\nu,\vartheta\in\mathbb{R}$. Therefore, compared to the Winfree model (3), in the analysis of the Kuramoto model (32), the diameter of the intrinsic frequencies $\omega_{i}$,
$D(\Omega)\coloneqq\max_{1\leq i,j\leq N}|\omega_{i}-\omega_{j}|,$
becomes the correct analogue of the maximum frequency $\max_{i}|\omega_{i}|$, and the notion of relative equilibria, namely solutions $\{\theta_{i}(t)\}_{i=1}^{N}$ such that
$\theta_{i}(t)-\theta_{j}(t)=\theta_{i}^{0}-\theta_{j}^{0},\quad\forall i,j=1,\cdots,N,\quad\forall t\geq 0,$
becomes the correct analogue of the notion of (absolute) equilibria.
One can also introduce order parameters, namely $R\in[0,1]$, $\phi\in\mathbb{R}$ such that
$Re^{\mathrm{i}\phi}=\frac{1}{N}\sum_{j=1}^{N}e^{\mathrm{i}\theta_{j}},$
which simplify the ODE of (32) to
$\dot{\theta}_{i}=\omega_{i}-\kappa R\sin(\theta_{i}-\phi)$
and which satisfy
$R=\frac{1}{N}\sum_{j=1}^{N}\cos(\theta_{j}-\phi).$ (33)
One of the main motivations for this paper was to create analogues for the Winfree model of some results that were previously known for the Kuramoto model, stated as follows. More precisely, Corollary 16 is the analogue of Theorem 43, and Theorem 1 is the analogue of Theorem 44.
Theorem 43 ([HR20, Theorem 3.3]).
Suppose that $\Theta^{0}$, $\Omega$ and $\kappa$ satisfy
$R_{0}\coloneqq R(\Theta^{0})>0,\quad\kappa>1.6\frac{D(\Omega)}{R_{0}^{2}},$
and let $\Theta=\Theta(t)$ be the solution to (32). Then, the following assertions hold.
  1. (a)
    Asymptotic complete phase-locking occurs, namely
    $\exists\lim_{t\to\infty}\left(\theta_{i}(t)-\theta_{j}(t)\right),~\mathrm{and}~\lim_{t\to\infty}\left(\dot{\theta}_{i}(t)-\dot{\theta}_{j}(t)\right)=0,\quad\forall i,j=1,\cdots,N.$
  2. (b)
    Let
    $\gamma(R_{0})=\begin{cases}0.5+\frac{0.35}{0.94}R_{0}&0<R_{0}\leq 0.94,\\ 1-2.5(1-R_{0})&0.94<R_{0}\leq 1.\end{cases}$
    Then there exists $\mathcal{A}\subset\{1,\cdots,N\}$ with $|\mathcal{A}|/N\geq\gamma(R_{0})$ such that for all large enough time $t$, $\Theta_{\mathcal{A}}(t)$ lies on an arc of length $\leq\frac{3\pi}{4(2\gamma(R_{0})-1)}\frac{D(\Omega)}{\kappa}$, up to modulo $2\pi$-translations.
  3. (c)
    The aforementioned stable ensemble $\Theta_{\mathcal{A}}$ becomes ordered in accordance with its natural frequencies: if $i,j\in\mathcal{A}$ and $\omega_{i}\geq\omega_{j}$, then
    $\frac{\omega_{i}-\omega_{j}}{\kappa}\leq\liminf_{t\rightarrow\infty}[\theta_{i}(t)-\theta_{j}(t)]\leq\limsup_{t\rightarrow\infty}[\theta_{i}(t)-\theta_{j}(t)]\leq c\frac{\omega_{i}-\omega_{j}}{\kappa}$
    after modulo $2\pi$-translations, where the constant $c$ depends only on $\gamma(R_{0})$ and $D(\Omega)/\kappa$.
Theorem 44 ([HR20, Theorem 3.2]).
Suppose that $\Omega$ and $\kappa$ satisfy
$\kappa>1.6ND(\Omega),$
Then for Lebesgue almost every inital data $\Theta^{0}$, the solution $\Theta=\Theta(t)$ to (32) with initial data $\Theta^{0}$. Then, the following assertions hold.
  1. (a)
    Asymptotic complete phase-locking occurs, namely
    $\exists\lim_{t\to\infty}\left(\theta_{i}(t)-\theta_{j}(t)\right),~\mathrm{and}~\lim_{t\to\infty}\left(\dot{\theta}_{i}(t)-\dot{\theta}_{j}(t)\right)=0,\quad\forall i,j=1,\cdots,N.$
  2. (b)
    Let
    $\gamma_{N}=0.5+\frac{0.35}{0.94\sqrt{N}}$
    Then there exists $\mathcal{A}\subset\{1,\cdots,N\}$ with $|\mathcal{A}|/N\geq\gamma_{N}$ such that for all large enough time $t$, $\Theta_{\mathcal{A}}(t)$ lies on an arc of length $\leq\frac{3\pi}{4(2\gamma_{N}-1)}\frac{D(\Omega)}{\kappa}$, up to modulo $2\pi$-translations.
  3. (c)
    The aforementioned stable ensemble $\Theta_{\mathcal{A}}$ becomes ordered in accordance with its natural frequencies: if $i,j\in\mathcal{A}$ and $\nu_{i}\geq\nu_{j}$, then
    $\frac{\nu_{i}-\nu_{j}}{\kappa}\leq\liminf_{t\rightarrow\infty}[\theta_{i}(t)-\theta_{j}(t)]\leq\limsup_{t\rightarrow\infty}[\theta_{i}(t)-\theta_{j}(t)]\leq c\frac{\nu_{i}-\nu_{j}}{\kappa}$
    after modulo $2\pi$-translations, where the constant $c$ depends only on $\gamma_{N}$ and $D(\Omega)/\kappa$.
The connection between the Winfree model and the Kuramoto model is that their ODEs are both equations of the form
$\dot{\theta}_{i}=\omega_{i}+A\cos\theta_{i}+B\sin\theta_{i},\quad i=1,\cdots,N,$ (34)
or alternatively of the form
$\begin{cases}\dot{\theta}_{i}(t)=\omega_{i}(\Theta(t))+\kappa R(t)\sin(\phi(t)-\theta_{i}(t)),\quad t>0,\\ \theta_{i}(0)=\theta_{i}^{0},\end{cases}\quad i=1,\cdots,N.$ (35)
with $\omega_{i}(t)=\omega(\Theta(t))$, $A(t)=A(\Theta(t))$, $B(t)=B(\Theta(t))$, $R(t)=R(\Theta(t))$, $\phi=\phi(\Theta(t))$ being functions of $\Theta$.
The equations (34) and (35) were shown to be highly integrable by Watanabe and Strogatz [WS94] in the case $\omega_{i}$ are identical. See also Chen, Engelbrecht, and Mirollo [CEM17] for a formulation of (35) in terms of a flow on the hyperbolic disc, and the work of Ha, Kim, Park and the author [HKPR21] for higher-dimensional generalizations of the Watanabe-Strogatz constants of motion.
The sinusoidal Winfree model (3) can be viewed as a special case of the general Kuramoto model (35), where we consider $2N$ oscillators $\theta_{1},\cdots,\theta_{2N}$ with
$\omega_{N+1}=\cdots=\omega_{2N}=0,\quad\theta_{N+1}^{0}=\cdots=\theta_{2N}^{0}=0$
and
$\phi=0,\quad R=\frac{1}{N}\sum_{j=1}^{2N}\cos(\phi-\theta_{j}).$
Note that $R$ is of the form (33), up to the factor $2$. Exactly half of the oscillators are fixed at 0 and are “beckoning” the other half to come close to 0.
Compared to the Kuramoto model, the Winfree model (3) has the $\mathrm{U}(1)$ symmetry broken; it appears that this makes the analysis of complete oscillator death relatively easier than that of asymptotic complete phase-locking in the Kuramoto model, but makes the analysis of complete locking harder. In fact, when comparing Corollary 16 versus Theorem 43, we see that Corollary 16 requires the weaker lower bound ${R_{0}}^{-1.5}$ on $\kappa$ than ${R_{0}}^{-2}$ of Theorem 43, while controlling a comparable number of oscillators (Corollary 16 guarantees the behavior of $O(R_{0})$ of the oscillators and $N$ “virtual” oscillators described above, and so in total guarantees the behavior of $\frac{1}{2}+O(R_{0})$ oscillators, just as in Theorem 43). Also, comparing Theorem 1 versus Theorem 44, we can see that we only require a weaker condition $\kappa>2\|\Omega\|_{\infty}$ in Theorem 1 while we require the somewhat unreasonably large condition $\kappa>1.6ND(\Omega)$ in Theorem 44; this is because the volumetric argument using the divergence which gives Theorem 1 from Corollary 16 and Theorem 44 from Theorem 43 tells us that the order parameter $R$ reaches the universal constant $1$ for the Winfree model while it reaches the small amount $\frac{1}{\sqrt{N}}$ for the Kuramoto model.
Due to the similarity of Corollary 16 and Theorem 43, it seems natural to pose the following question.
Question 45.
For which functions $\phi$ and $R$ with (33) do there exist a constant $c_{1}$ and exponent $p>0$ such that if $\Theta=\Theta(t)$ is the solution to (35) with $\omega_{i}(\Theta)=\omega_{i}\in\mathbb{R}$, $R_{0}=R(\Theta^{0})>0$, and $\kappa>c_{1}\frac{\max_{i}|\omega_{i}|}{R_{0}^{p}}$, then
$\sup_{t\geq 0}|\theta_{i}(t)-\theta_{j}(t)|<\infty,\quad i,j=1,\cdots,N$
holds? If so, is there a constant $c_{2}$ and an exponent $q>0$ such that there exists a subset $\mathcal{A}\subset\{1,\cdots,N\}$ with $|\mathcal{A}|/N\geq\frac{1}{2}+c_{2}R_{0}^{q}$ such that
$\limsup_{t\to\infty}|\theta_{i}(t)-\theta_{j}(t)|\lesssim_{R_{0}}\frac{\max_{i}|\omega_{i}|}{\kappa}$
holds after modulo $2\pi$ translations? What about $\phi$ and $R$ which do not satisfy (33)?
4. An order parameter bootstrapping argument
The purpose of this section is to prove Theorems 15 and 21, Corollary 16, and Propositions 32 and 33. Roughly, the proof idea is as follows. Choosing a suitable $\rho\in(0,R_{0})$, we will find a subpopulation $\Theta_{\mathcal{B}}$ that we have some control over, in the sense that for all sufficiently small times $t\geq 0$ we have $R(t)\geq\frac{1}{N}\sum_{i\in\mathcal{B}}I(\theta_{i})>\rho$. By “continuous induction”, we will be able to guarantee this property for all times $t\geq 0$.
We first prove toy versions of Theorem 21, which tell us why it is important to have a uniform-in-time lower bound on the order parameter when proving oscillator death. We then prove Theorem 21 itself, which is a version of Corollary 16 (which in turn is a consequence of Theorem 15) for more general interaction functions, before proving Theorem 15, because the argument is simpler in this more general setting and so it is easier to see how the bootstrapping proof works and why we get the exponent $1+\frac{p}{q}$ in Theorem 21. We will then prove Theorem 15 by specializing this proof method to the system (3) while optimizing the argument to gain some control on the constants involved.
We will also obtain criteria for partial oscillator death in this section, namely we will prove Propositions 32 and 33. Complete oscillator death will be proven by an application of these criteria.
We will prove the toy versions of Theorem 21 in subsection 4.1, Theorem 21 and Proposition 33 in subsection 4.2, and Theorem 15, Corollary 16, and Proposition 32 in subsection 4.3.
4.1. Two toy models
4.1.1. Each oscillator constrains itself
Suppose
$I\geq 0,\quad\min_{\theta\in[-\pi,\pi]}I(\theta)S(\theta)<0<\max_{\theta\in[-\pi,\pi]}I(\theta)S(\theta).$ (36)
In the single oscillator case $N=1$ the ODE becomes
$\dot{\theta}_{1}=\omega_{1}+\kappa I(\theta)S(\theta)$
and it is easy to see that $\kappa>\frac{|\omega_{1}|}{\min\{-\min IS,\max IS\}}$ implies boundedness of $\theta_{1}$. It is not hard to extend this argument to higher $N$ if we are willing to pay the price of a $\kappa$ proportional to $N$.
Proposition 46 (Criterion for partial oscillator death under (36)).
Let $\Theta=\Theta(t)$ be the solution to (1) with parameters $\{\omega_{i}\}_{i=1}^{N}$, $\kappa$, and initial data $\{\theta_{i}^{0}\}_{i=1}^{N}$, and with the $2\pi$-periodic Lipschitz interaction functions $I$ and $S$ satisfying (36).
Fix $\mathcal{B}\subset\{1,\cdots,N\}$. If
$\kappa>\frac{N\|\Omega_{\mathcal{B}}\|_{\infty}}{\min\{-\min IS,\max IS\}},$ (37)
then
$\sup_{t\geq 0}\theta_{i}(t)-\inf_{t\geq 0}\theta_{i}(t)<2\pi\quad\forall i\in\mathcal{B}.$
Proof.
Let $\alpha_{0},\beta_{0}\in[-\pi,\pi]$ such that
$\alpha_{0}\in\arg\min IS,\quad\beta_{0}\in\arg\max IS,$
so that $(IS)(\alpha_{0})<0$ and $(IS)(\beta_{0})>0$. In particular, $I(\alpha_{0})>0$, $S(\alpha_{0})<0$, $I(\beta_{0})>0$, and $S(\beta_{0})>0$. For any $i\in\mathcal{B}$ we see by (37) that $\theta_{i}(t_{0})\equiv\alpha_{0}\mod 2\pi$ implies
[…]
and similarly that $\theta_{i}(t_{0})\equiv\beta_{0}\mod 2\pi$ implies $\dot{\theta}_{i}(t_{0})>0$. Thus, if we choose integers $k_{1},k_{2}\in\mathbb{Z}$ such that $\theta_{i}^{0}\in[\beta_{0}+2k_{1}\pi,\alpha_{0}+2k_{2}\pi]$ and $\mathrm{length}[\beta_{0}+2k_{1}\pi,\alpha_{0}+2k_{2}\pi]<2\pi$, then $\theta_{i}(t)\in[\beta_{0}+2k_{1}\pi,\alpha_{0}+2k_{2}\pi]$ for all $t\geq 0$. ∎
In particular, Proposition 46 tells us that the pathwise critical coupling strength is finite everywhere:
$\kappa_{\mathrm{pc}}(\Theta^{0},\Omega,I,S)\leq\frac{N\|\Omega\|_{\infty}}{\min\{-\min IS,\max IS\}}.$
However, this is not an effective bound for large $N$.
4.1.2. Importance of a lower bound on the average influence
Note that in (38) we have wasted many nonpositive terms. The main idea of the bootstrapping argument for Theorems 15 and 21 is to guarantee a uniform-in-time lower bound on the average influence $R$. To see why we need such a condition, we examine the case when the interaction functions $I$ and $S$ satisfy the following:
$I(\theta)\geq\rho\quad\forall\theta\in[-\pi,\pi]$ (39)
for some $\rho>0$ and
$\min_{\theta\in[-\pi,\pi]}S(\theta)<0<\max_{\theta\in[-\pi,\pi]}S(\theta).$ (40)
Proposition 47 (Criterion for partial oscillator death under (39)-(40)).
Let $\Theta=\Theta(t)$ be the solution to (1) with parameters $\{\omega_{i}\}_{i=1}^{N}$, $\kappa$, and initial data $\{\theta_{i}^{0}\}_{i=1}^{N}$, and with the $2\pi$-periodic Lipschitz interaction functions $I$ and $S$ satisfying (39) and (40).
Fix $\mathcal{B}\subset\{1,\cdots,N\}$. If
$\kappa>\frac{\|\Omega_{\mathcal{B}}\|_{\infty}}{\rho\min\{-\min S,\max S\}},$ (41)
then
$\sup_{t\geq 0}\theta_{i}(t)-\inf_{t\geq 0}\theta_{i}(t)<2\pi\quad\forall i\in\mathcal{B}.$
Proof.
Let $\alpha_{0},\beta_{0}\in[-\pi,\pi]$ be such that
$\alpha_{0}\in\arg\min S,\quad\beta_{0}\in\arg\max S,$
so that $S(\alpha_{0})<0$ and $S(\beta_{0})>0$. For any $i\in\mathcal{B}$ we see by (13) and (41) (and the observation that $R\geq\rho$) that $\theta_{i}(t_{0})\equiv\alpha_{0}\mod 2\pi$ implies $\theta_{i}(t_{0})<0$ and that $\theta_{i}(t_{0})\equiv\beta_{0}\mod 2\pi$ implies $\dot{\theta}_{i}(t_{0})>0$. Thus, if we choose integers $k_{1},k_{2}\in\mathbb{Z}$ such that $\theta_{i}^{0}\in[\beta_{0}+2k_{1}\pi,\alpha_{0}+2k_{2}\pi]$ and $\mathrm{length}[\beta_{0}+2k_{1}\pi,\alpha_{0}+2k_{2}\pi]<2\pi$, then $\theta_{i}(t)\in[\beta_{0}+2k_{1}\pi,\alpha_{0}+2k_{2}\pi]$ for all $t\geq 0$. ∎
Therefore, for $I$ and $S$ satisfying (39) and (40), complete oscillator death happens for any initial data whenever $\kappa$ exceeds the fixed value given in (41) with $\mathcal{B}=\{1,\cdots,N\}$:
$\kappa>\frac{\|\Omega\|_{\infty}}{\rho\min\{-\min S,\max S\}}.$
In other words, the pathwise critical coupling strength (recall Definition 24) is uniformly bounded as follows:
$\kappa_{\mathrm{pc}}(\Theta^{0},\Omega,I,S)\leq\frac{\|\Omega\|_{\infty}}{\rho\min\{-\min S,\max S\}}.$
Moreover, by choosing different $\mathcal{B}$, we obtain various sufficient coupling strengths for partial oscillator death. We have a full proof of the existence of the oscillator death regime (and partial oscillator death regime) in this case.
The rest of this section is an imitation of this ideal case. We will find conditions that serve as a proxy for (39) and thus guarantee oscillator death. The price for assuming these conditions will ultimately be paid either in the form of a large coupling strength depending on the initial data as in Corollary 16 and Theorem 21, or by sacrificing a certain singular set of initial data (which, to our advantage, turns out to have asymptotically small measure) as in Theorems 1, 6, 7, 26, 27, 28, and 30.
4.2. The general interaction function case
The purpose of this subsection is to prove Theorem 21 and Proposition 33.
Recall conditions (9), (10), and (11): there is a constant $\alpha_{0}\in(0,\pi)$ such that
[…]
$0\leq I(\theta)\leq c_{2}(\pi-|\theta|)^{q}\mbox{ for }\theta\in[-\pi,\pi],$
$\min_{|\phi|\leq\max\{|\theta|,\alpha_{0}\}}I(\phi)\geq c_{3}I(\theta),\mbox{ for }\theta\in[-\pi,\pi].$
The starting point of the analysis is the observation that, if we make the a priori assumption that the order parameter $R(t)$ is uniformly bounded below in time, then we have some control on the behavior of some of the variables.
Lemma 48 (A priori estimate).
Let $\Theta=\Theta(t)$ be the solution to (1) with parameters $\{\omega_{i}\}_{i=1}^{N}$, $\kappa$, and initial data $\{\theta_{i}^{0}\}_{i=1}^{N}$, and let II and $S$ satisfy (9), (10), and (11). Suppose $T>0$, $0<\rho\leq\|I\|_{\sup}$, and $\mathcal{B}\subseteq\{1,\cdots,N\}$ satisfy
$R(t)\geq\rho,~\forall t\in[0,T),\quad and\quad\kappa>\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}}{\rho c_{1}(\pi-\alpha_{0})^{p}}.$
Let $i\in\mathcal{B}$ be such that $\theta_{i}^{0}\in\left[-\pi+\left(\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}}{\kappa\rho c_{1}}\right)^{1/p},\pi-\left(\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}}{\kappa\rho c_{1}}\right)^{1/p}\right]$. Then $|\theta_{i}(t)|\leq\max\{|\theta_{i}^{0}|,\alpha_{0}\}$ and $I(\theta_{i}(t))\geq c_{3}I(\theta_{i}^{0})$ for $t\in[0,T)$.
Proof.
We compute that for any $t_{0}\in[0,T)$, if $\theta_{i}(t_{0})\in\left[\alpha_{0},\pi-\left(\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}}{\kappa\rho c_{1}}\right)^{1/p}\right]$ (note that our choice of $\kappa$ makes this a nonempty interval) then
[…]
and similarly if $\theta_{i}(t_{0})\in\left[-\pi+\left(\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}}{\kappa\rho c_{1}}\right)^{1/p},-\alpha_{0}\right]$ then $\dot{\theta}_{i}(t_{0})\geq 0$. By a standard exit-time argument, we have $|\theta_{i}(t)|\leq\max\{|\theta_{i}^{0}|,\alpha_{0}\}$ for $t\in[0,T)$. This then implies $I(\theta_{i}(t))\geq c_{3}I(\theta_{i}^{0})$ by (11). ∎
All in all, Lemma 48 says that if we have an a priori uniform-in-time lower bound for the average influence $R=\frac{1}{N}\sum_{j=1}^{N}I(\theta_{j})$, then for those $\theta_{i}$ that are far from $\pi$ (modulo $2\pi$) we have a lower bound on $I(\theta_{i})$. If we can find a large enough number of such $\theta_{i}$’s, we would be able to deduce a lower bound on their average $R$. The main idea of our proof of Theorem 21 is to formalize this into a bootstrapping argument so as to guarantee a lower bound on $R(t)$ for all $t\geq 0$.
Proof of Proposition 33.
Define
$\mathcal{T}\coloneqq\{T>0:R(t)>\rho\mbox{ for }t\in[0,T)\}.$
Noting from $c_{3}\leq 1$ and (28)-(i) that
${R(0)}\geq\frac{1}{N}\sum_{i\in\mathcal{A}}I(\theta_{i}^{0})>\frac{\rho}{c_{3}}\geq\rho,$
by continuity of $R(t)$ we may find a small $\delta>0$ such that $R(t)>\rho$ for $t\in[0,\delta)$. It then follows that $\delta\in\mathcal{T}$, so $T^{*}\coloneqq\sup\mathcal{T}>0$. It is immediate that $\mathcal{T}\supset(0,T^{*})$.
Now by (28) and (29), Lemma 48 applies up to time $T^{*}$, so we have $I(\theta_{i}(t))\geq c_{3}I(\theta_{i}^{0})$ for all $i\in\mathcal{A}$ and all $t\in[0,T^{*})$, which gives
$R(t)\geq\frac{1}{N}\sum_{i\in\mathcal{A}}I(\theta_{i}(t))\geq c_{3}\cdot\frac{1}{N}\sum_{i\in\mathcal{A}}I(\theta_{i}^{0}),\quad t\in[0,T^{*}).$
If we had $T^{*}<\infty$, then by continuity of $R$ we would have
[…]
and again by continuity of $R$ there would exist a $\delta>0$ such that $R(t)>\rho$ for $t\in[T^{*},T^{*}+\delta)$, contradicting the definition of $T^{*}$. Thus, we must have $T^{*}=\infty$, i.e., $R(t)>\rho$ for all $t\geq 0$. This is assertion (a) of Proposition 33.
Since Lemma 48 applies up to time $T^{*}=\infty$, we have assertion (b) of Proposition 33.
It remains to prove assertion (c) of Proposition 33, i.e., $\sup_{t\geq 0}\theta_{i}(t)-\inf_{t\geq 0}\theta_{i}(t)<2\pi$ for all $i\in\mathcal{B}$. We need to separate into two cases. Fix an $i\in\mathcal{B}$.
  • Case 1.
    $\theta_{i}^{0}\in\left[-\pi+\left(\frac{\|\Omega_{\mathcal{B}}\|_{\infty}}{\kappa\rho c_{1}}\right)^{1/p},\pi-\left(\frac{\|\Omega_{\mathcal{B}}\|_{\infty}}{\kappa\rho c_{1}}\right)^{1/p}\right]\mod 2\pi$.
    Then, by Lemma 48, we have $\theta_{i}(t)\in\left[-\max\{|\theta_{i}^{0}|,\alpha_{0}\},\max\{|\theta_{i}^{0}|,\alpha_{0}\}\right]\mod 2\pi$ for all $t\geq 0$, from which the assertion $\sup_{t\geq 0}\theta_{i}(t)-\inf_{t\geq 0}\theta_{i}(t)<2\pi$ easily follows.
  • Case 2.
    $\theta_{i}^{0}\notin\left[-\pi+\left(\frac{\|\Omega_{\mathcal{B}}\|_{\infty}}{\kappa\rho c_{1}}\right)^{1/p},\pi-\left(\frac{\|\Omega_{\mathcal{B}}\|_{\infty}}{\kappa\rho c_{1}}\right)^{1/p}\right]\mod 2\pi$.
    Then either $\theta_{i}(t)\notin\left[-\pi+\left(\frac{\|\Omega_{\mathcal{B}}\|_{\infty}}{\kappa\rho c_{1}}\right)^{1/p},\pi-\left(\frac{\|\Omega_{\mathcal{B}}\|_{\infty}}{\kappa\rho c_{1}}\right)^{1/p}\right]\mod 2\pi$ for all $t\geq 0$, in which case $\theta_{i}(t)$ is trapped in an interval of length $2\pi$ for all time and the assertion follows, or there will be the smallest time $t^{*}>0$ at which $\theta_{i}(t^{*})=-\pi+\left(\frac{\|\Omega_{\mathcal{B}}\|_{\infty}}{\kappa\rho c_{1}}\right)^{1/p}\mbox{ or }\pi-\left(\frac{\|\Omega_{\mathcal{B}}\|_{\infty}}{\kappa\rho c_{1}}\right)^{1/p}\mod 2\pi$. In the latter case, we can apply Lemma 48 starting at time $t^{*}$ (this is possible since (3) is a time-autonomous system) and deduce that the set $\{\theta_{i}(t):t\geq 0\}$ will be contained in an interval either of the form $\left(-\pi-\left(\frac{\|\Omega_{\mathcal{B}}\|_{\infty}}{\kappa\rho c_{1}}\right)^{1/p},\pi-\left(\frac{\|\Omega_{\mathcal{B}}\|_{\infty}}{\kappa\rho c_{1}}\right)^{1/p}\right]\mod 2\pi$ or $\left[-\pi+\left(\frac{\|\Omega_{\mathcal{B}}\|_{\infty}}{\kappa\rho c_{1}}\right)^{1/p},\pi+\left(\frac{\|\Omega_{\mathcal{B}}\|_{\infty}}{\kappa\rho c_{1}}\right)^{1/p}\right)\mod 2\pi$, and the assertion easily follows in both of these cases.
We are now ready to prove Theorem 21. We will apply our criterion for partial oscillator death given in Proposition 33. By choosing the coupling strength $\kappa$ to be sufficiently large depending on the initial data, we will be able to find a subensemble $\mathcal{A}$ which satisfies the conditions (28)-(i) and (29).
Proof of Theorem 21.
After modulo $2\pi$ shifts, we may assume that $\theta_{i}^{0}\in[-\pi,\pi)$ for all $i=1,\cdots,N$. Define
$\mathcal{A}=\{i=1,\cdots,N:I(\theta_{i}^{0})\geq\frac{{R_{0}}}{2}\}.$
Then
[…]
We can also see by (10) that
[…]
We may now see that (42) and (43) satisfies the hypotheses of Proposition 33 with $\rho=\frac{c_{3}R_{0}}{2}$ and $\mathcal{B}=\{1,\cdots,N\}$, with (14)-(ii) becoming (28)-(ii). Now Proposition 33 (a,c) give statements (a) and (b). ∎
4.3. The special interaction function case
The purpose of this subsection is to prove Theorem 15 and derive Corollary 16 from it. We will prove Proposition 32 along the way. Note that, with Theorem 21 at hand, we already have a version of Corollary 16 with worse constants. We are presenting a refined version of the proof of Theorem 21 to show how one may optimize the analysis to obtain a reasonable coupling strength that is closer to the range considered in [AS01]. Thus, in this subsection, we will try to be precise in our analysis to get good constants.
As in the previous subsection, the starting point of the analysis is the observation that, if we make the a priori assumption that the order parameter $R(t)$ is uniformly bounded below in time, then we have some control on the dynamics of some of the variables. The following is the analogue of Lemma 48 for system (3).
Lemma 49 (A priori estimate).
Let $\Theta=\Theta(t)$ be the solution to (3) with parameters $\{\omega_{i}\}_{i=1}^{N}$, $\kappa$, and $\{\theta_{i}^{0}\}_{i=1}^{N}$. Suppose $T>0$, $0<\rho\leq 2$, and $\mathcal{B}\subseteq\{1,\cdots,N\}$ satisfy
$R(t)\geq\rho,~\forall t\in[0,T),\quad\mathrm{and}\quad\kappa>\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}}{\rho}.$
Let $i\in\mathcal{B}$ be such that $\cos\theta^{0}_{i}>-\sqrt{1-\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}^{2}}{\kappa^{2}\rho^{2}}}$. Then
  1. (a)
    $\cos\theta_{i}(t)\geq\cos\theta_{i}^{0}$ for $t\in[0,T)$ if $\cos\theta^{0}_{i}\leq\sqrt{1-\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}^{2}}{\kappa^{2}\rho^{2}}}$,
  2. (b)
    $\cos\theta_{i}(t)\geq\sqrt{1-\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}^{2}}{\kappa^{2}\rho^{2}}}$ for $t\in[0,T)$ if $\cos\theta^{0}_{i}>\sqrt{1-\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}^{2}}{\kappa^{2}\rho^{2}}}$,
  3. (c)
    $1+\cos\theta_{i}(t)\geq\frac{1}{2}\left(1+\sqrt{1-\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}^{2}}{\kappa^{2}\rho^{2}}}\right)(1+\cos\theta_{i}^{0})$ for all $t\in[0,T)$,
  4. (d)
    for a given $0<\alpha<\sqrt{1-\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}^{2}}{\kappa^{2}\rho^{2}}}$, if $\cos\theta_{i}^{0}\geq-\alpha$, then $\cos\theta_{i}(t)\geq-\alpha$ for $t\in[0,T)$,
  5. (e)
    for a given $0<\alpha<\sqrt{1-\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}^{2}}{\kappa^{2}\rho^{2}}}$, if $T>\frac{\pi}{\kappa\rho\sqrt{1-\alpha^{2}}-{\|\Omega_{\mathcal{B}}\|_{\infty}}}$ and if $\cos\theta_{i}^{0}\geq-\alpha$, then $\cos\theta_{i}(t)\geq\alpha$ for $t\in\left[\frac{\pi}{\kappa\rho\sqrt{1-\alpha^{2}}-{\|\Omega_{\mathcal{B}}\|_{\infty}}},T\right)$.
Furthermore, let $0<\alpha<\sqrt{1-\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}^{2}}{\kappa^{2}\rho^{2}}}$ be such that $\cos\theta_{i}^{0}>-\alpha$, and suppose we chose another $j\in\mathcal{B}$ such that $\cos\theta^{0}_{j}>-\alpha$. Suppose $T>\frac{\pi}{\kappa\rho\sqrt{1-\alpha^{2}}-{\|\Omega_{\mathcal{B}}\|_{\infty}}}$. Then
  1. (f)
    if $\omega_{i}>\omega_{j}$, then after modulo $2\pi$ translations,
    $\theta_{i}(t)-\theta_{j}(t)\leq\frac{\omega_{i}-\omega_{j}}{\kappa\rho\alpha}+\pi\exp\left(-\kappa\rho\alpha\left(t-\frac{\pi}{\kappa\rho\sqrt{1-\alpha^{2}}-{\|\Omega_{\mathcal{B}}\|_{\infty}}}\right)\right)\mbox{ for }t\in\left[\frac{\pi}{\kappa\rho\sqrt{1-\alpha^{2}}-{\|\Omega_{\mathcal{B}}\|_{\infty}}},T\right),$
    and if $T>\frac{\pi}{\kappa\rho\sqrt{1-\alpha^{2}}-{\|\Omega_{\mathcal{B}}\|_{\infty}}}+\frac{\pi}{\omega_{i}-\omega_{j}}$,
    $\theta_{i}(t)-\theta_{j}(t)\geq\frac{\omega_{i}-\omega_{j}}{2\kappa}-\frac{\omega_{i}-\omega_{j}}{2\kappa}\exp\left(-2\kappa\left(t-\frac{\pi}{\kappa\rho\sqrt{1-\alpha^{2}}-{\|\Omega_{\mathcal{B}}\|_{\infty}}}-\frac{\pi}{\omega_{i}-\omega_{j}}\right)\right)$
    $\mbox{for }t\in\left[\frac{\pi}{\kappa\rho\sqrt{1-\alpha^{2}}-{\|\Omega_{\mathcal{B}}\|_{\infty}}}+\frac{\pi}{\omega_{i}-\omega_{j}},T\right),$
  2. (g)
    if $\omega_{i}=\omega_{j}$, then after modulo $2\pi$ translations,
    $|\theta_{i}(t)-\theta_{j}(t)|\leq\pi\exp\left(-\kappa\rho\alpha\left(t-\frac{\pi}{\kappa\rho\sqrt{1-\alpha^{2}}-{\|\Omega_{\mathcal{B}}\|_{\infty}}}\right)\right)\mbox{ for }t\in\left[\frac{\pi}{\kappa\rho\sqrt{1-\alpha^{2}}-{\|\Omega_{\mathcal{B}}\|_{\infty}}},T\right).$
Proof.
We first observe that for any $t_{0}\in[0,T)$ such that $\cos\theta_{i}(t_{0})\in\left[-\sqrt{1-\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}^{2}}{\kappa^{2}\rho^{2}}},\sqrt{1-\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}^{2}}{\kappa^{2}\rho^{2}}}\right]$, we have
$\left.\frac{d}{dt}\right|_{t=t_{0}}\cos\theta_{i}=$ $-\dot{\theta}_{i}(t_{0})\sin\theta_{i}(t_{0})=-\omega_{i}\sin\theta_{i}(t_{0})+\kappa R(t_{0})\sin^{2}\theta_{i}(t_{0})$
$\geq$ $(-{\|\Omega_{\mathcal{B}}\|_{\infty}}+\kappa\rho|\sin\theta_{i}(t_{0})|)|\sin\theta_{i}(t_{0})|\geq 0.$
In other words, $\cos\theta_{i}(t)$ behaves as a non-decreasing function of $t$ whenever its value lies in $\left[-\sqrt{1-\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}^{2}}{\kappa^{2}\rho^{2}}},\sqrt{1-\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}^{2}}{\kappa^{2}\rho^{2}}}\right]$. A simple comparison argument gives assertions (a) and (b).
Assertions (c) and (d) immediately follows from assertions (a) and (b).
To show assertion (e), we calculate that, for any $t_{0}\in[0,T)$ such that $\theta_{i}(t_{0})\in[\cos^{-1}\alpha,\pi-\cos^{-1}\alpha]\mod 2\pi$, we have
$\dot{\theta}_{i}(t_{0})=\Omega_{i}-\kappa R(t_{0})\sin\theta_{i}(t_{0})\leq{\|\Omega_{\mathcal{B}}\|_{\infty}}-\kappa\rho\sqrt{1-\alpha^{2}}<0.$
Similarly, if $\theta_{i}(t_{0})\in[-\pi+\cos^{-1}\alpha,-\cos^{-1}\alpha]\mod 2\pi$, we have
$\dot{\theta}_{i}(t_{0})\geq-{\|\Omega_{\mathcal{B}}\|_{\infty}}+\kappa\rho\sqrt{1-\alpha^{2}}>0.$
A simple comparison argument now gives assertion (e).
For assertions (f) and (g), note that for $t\in\left[\frac{\pi}{\kappa\rho\sqrt{1-\alpha^{2}}-{\|\Omega_{\mathcal{B}}\|_{\infty}}},T\right)$, we have by (e) that $\theta_{i}(t),\theta_{j}(t)\in[-\cos^{-1}\alpha,\cos^{-1}\alpha]$. Using $\dot{\theta}_{i}-\dot{\theta}_{j}=\omega_{i}-\omega_{j}-\kappa R(\sin\theta_{i}-\sin\theta_{j})$, we have
$\theta_{i}(t)\geq\theta_{j}(t)\Rightarrow\omega_{i}-\omega_{j}-2\kappa(\theta_{i}(t)-\theta_{j}(t))\leq\dot{\theta}_{i}(t)-\dot{\theta}_{j}(t)\leq\omega_{i}-\omega_{j}-\kappa\rho\alpha(\theta_{i}(t)-\theta_{j}(t))$ (44)
$\theta_{i}(t)\leq\theta_{j}(t)\Rightarrow\dot{\theta}_{i}(t)-\dot{\theta}_{j}(t)\geq\omega_{i}-\omega_{j}-\kappa\rho\alpha(\theta_{i}-\theta_{j})\geq\omega_{i}-\omega_{j}.$ (45)
The first part of statement (f) and statement (g) follow easily from the right-hand side of (44). On the other hand, for the second statement of (f), (45) tells us that $\theta_{i}(t)>\theta_{j}(t)$ for $t\in\left[\frac{\pi}{\kappa\rho\sqrt{1-\alpha^{2}}-{\|\Omega_{\mathcal{B}}\|_{\infty}}}+\frac{\pi}{\omega_{i}-\omega_{j}},T\right)$, on which we apply the left-hand side of (44) to obtain the desired statement. ∎
As a direct corollary, the Winfree model (3) is well-controlled if we have a positive lower bound on the order parameter for all time. Compared to the previous subsection, we state the following extra Lemma separately, since we know additional information about the asymptotic dynamics thanks to the Łojasiewicz gradient theorem (Proposition 19).
Lemma 50.
Let $\Theta=\Theta(t)$ be the solution to (3) with parameters $\{\omega_{i}\}_{i=1}^{N}$, $\kappa$, and $\{\theta_{i}^{0}\}_{i=1}^{N}$. Suppose $0<\rho\leq 2$ and $\mathcal{B}\subseteq\{1,\cdots,N\}$ satisfy
$R(t)\geq\rho,~\forall t\in[0,\infty),\quad and\quad\kappa>\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}}{\rho}.$
  1. (a)
    Let $i\in\mathcal{B}$ be such that $\cos\theta^{0}_{i}\geq-\sqrt{1-\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}^{2}}{\kappa^{2}\rho^{2}}}$. Then $\cos\theta_{i}(t)\geq-\sqrt{1-\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}^{2}}{\kappa^{2}\rho^{2}}}$ and $1+\cos\theta_{i}(t)\geq\frac{1}{2}\left(1+\sqrt{1-\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}^{2}}{\kappa^{2}\rho^{2}}}\right)(1+\cos\theta_{i}^{0})$ for all $t\in[0,\infty)$.
  2. (b)
    Let $i\in\mathcal{B}$ and $0<\alpha<\sqrt{1-\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}^{2}}{\kappa^{2}\rho^{2}}}$ be such that $\cos\theta^{0}_{i}\geq-\alpha$. Then $\cos\theta_{i}(t)\geq-\alpha$ for $t\geq 0$ and $\cos\theta_{i}(t)\geq\alpha$ for $t\geq\frac{\pi}{\kappa\rho\sqrt{1-\alpha^{2}}-{\|\Omega_{\mathcal{B}}\|_{\infty}}}$.
  3. (c)
    Let $i,j\in\mathcal{B}$ and $0<\alpha<\sqrt{1-\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}^{2}}{\kappa^{2}\rho^{2}}}$ be such that $\cos\theta^{0}_{i}\geq-\alpha$ and $\cos\theta^{0}_{j}\geq-\alpha$. If $\omega_{i}>\omega_{j}$, then after modulo $2\pi$ translations,
    $\theta_{i}(t)-\theta_{j}(t)\leq\frac{\omega_{i}-\omega_{j}}{\kappa\rho\alpha}+\pi\exp\left(-\kappa\rho\alpha\left(t-\frac{\pi}{\kappa\rho\sqrt{1-\alpha^{2}}-{\|\Omega_{\mathcal{B}}\|_{\infty}}}\right)\right)\mbox{ for }t\geq\frac{\pi}{\kappa\rho\sqrt{1-\alpha^{2}}-{\|\Omega_{\mathcal{B}}\|_{\infty}}},$
    and
    $\theta_{i}(t)-\theta_{j}(t)\geq\frac{\omega_{i}-\omega_{j}}{2\kappa}-\frac{\omega_{i}-\omega_{j}}{2\kappa}\exp\left(-2\kappa\left(t-\frac{\pi}{\kappa\rho\sqrt{1-\alpha^{2}}-{\|\Omega_{\mathcal{B}}\|_{\infty}}}-\frac{\pi}{\omega_{i}-\omega_{j}}\right)\right)$
    $\mbox{for }t\geq\frac{\pi}{\kappa\rho\sqrt{1-\alpha^{2}}-{\|\Omega_{\mathcal{B}}\|_{\infty}}}+\frac{\pi}{\omega_{i}-\omega_{j}}.$
    If $\omega_{i}=\omega_{j}$, then after modulo $2\pi$ translations, $\lim_{t\to\infty}\left(\theta_{i}(t)-\theta_{j}(t)\right)=0$ exponentially:
    $|\theta_{i}(t)-\theta_{j}(t)|\leq\pi\exp\left(-\kappa\rho\alpha\left(t-\frac{\pi}{\kappa\rho\sqrt{1-\alpha^{2}}-{\|\Omega_{\mathcal{B}}\|_{\infty}}}\right)\right)\mbox{ for }t\geq\frac{\pi}{\kappa\rho\sqrt{1-\alpha^{2}}-{\|\Omega_{\mathcal{B}}\|_{\infty}}}.$
  4. (d)
    For all $i\in\mathcal{B}$,
    $\sup_{t\geq 0}\theta_{i}(t)-\inf_{t\geq 0}\theta_{i}(t)<2\pi.$
  5. (e)
    If $\mathcal{B}=\{1,\cdots,N\}$, then for all $i=1,\cdots,N$, the limit $\lim_{t\to\infty}\theta_{i}(t)$ exists, and
    $\lim_{t\to\infty}\dot{\theta}_{i}(t)=0.$
Proof.
Assertions (a), (b), and (c) follow immediately from Lemma 49. To prove assertion (d), fix an $i\in\mathcal{B}$.
  • Case 1.
    $\theta_{i}^{0}\in\left[-\cos^{-1}\Big(-\sqrt{1-\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}^{2}}{\kappa^{2}\rho^{2}}}\Big),\cos^{-1}\Big(-\sqrt{1-\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}^{2}}{\kappa^{2}\rho^{2}}}\Big)\right]\mod 2\pi$.
    Then $\theta_{i}(t)\in\left[-\cos^{-1}\Big(-\sqrt{1-\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}^{2}}{\kappa^{2}\rho^{2}}}\Big),\cos^{-1}\Big(-\sqrt{1-\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}^{2}}{\kappa^{2}\rho^{2}}}\Big)\right]\mod 2\pi$ for all $t\geq 0$, and $\theta_{i}(t)$ will belong to the same connected component, which has length $2\pi$. Assertion (d) follows in this case.
  • Case 2.
    $\theta_{i}^{0}\in\left(\cos^{-1}\Big(-\sqrt{1-\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}^{2}}{\kappa^{2}\rho^{2}}}\Big),2\pi-\cos^{-1}\Big(-\sqrt{1-\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}^{2}}{\kappa^{2}\rho^{2}}}\Big)\right)\mod 2\pi$.
    Then either $\theta_{i}(t)\in\left(\cos^{-1}\Big(-\sqrt{1-\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}^{2}}{\kappa^{2}\rho^{2}}}\Big),2\pi-\cos^{-1}\Big(-\sqrt{1-\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}^{2}}{\kappa^{2}\rho^{2}}}\Big)\right)\mod 2\pi$ for all $t\geq 0$ while staying in the same connected component, in which case assertion (d) follows, or there will be the smallest time $t^{*}>0$ at which $\theta_{i}(t^{*})=\cos^{-1}\Big(-\sqrt{1-\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}^{2}}{\kappa^{2}\rho^{2}}}\Big)\mbox{ or }2\pi-\cos^{-1}\Big(-\sqrt{1-\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}^{2}}{\kappa^{2}\rho^{2}}}\Big)\mod 2\pi$. In the latter case, we can apply Lemma 49 starting at time $t^{*}$ (this is possible since (3) is a time-autonomous system) and deduce that the set $\{\theta_{i}(t):t\geq 0\}$ will be contained in an interval either of the form $\left(\cos^{-1}\Big(-\sqrt{1-\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}^{2}}{\kappa^{2}\rho^{2}}}\Big),2\pi+\cos^{-1}\Big(-\sqrt{1-\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}^{2}}{\kappa^{2}\rho^{2}}}\Big)\right]\mod 2\pi$ or $\left[-\cos^{-1}\Big(-\sqrt{1-\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}^{2}}{\kappa^{2}\rho^{2}}}\Big),2\pi-\cos^{-1}\Big(-\sqrt{1-\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}^{2}}{\kappa^{2}\rho^{2}}}\Big)\right)\mod 2\pi$, and assertion (d) follows in both of these cases.
Assertion (e) follows from assertion (d) by Proposition 19. ∎
To guarantee the a priori lower bound, we can use a bootstrapping argument as in the previous subsection, which allows us to deduce the criterion for partial oscillator death, namely Proposition 32.
Proof of Proposition 32.
It is clear, by Lemma 50, that (b)-(f) follow once we prove (a). To prove (a), define
$\mathcal{T}\coloneqq\{T>0:R(t)\geq\rho\mbox{ for }t\in[0,T)\}.$
Noting that
${R_{0}}\geq\frac{1}{N}\sum_{i\in\mathcal{A}}(1+\cos\theta_{i}^{0})>\frac{2\rho}{1+\sqrt{1-\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}^{2}}{\kappa^{2}\rho^{2}}}}\geq\rho,$
by continuity of $R(t)$ we may find a small $\delta>0$ such that $R(t)>\rho$ for $t\in[0,\delta)$. By Lemma 49 (a)-(b) and (27), we see that $\cos\theta_{i}(t)\geq\cos\theta_{i}^{0}\geq-\sqrt{1-\frac{{\|\Omega_{\mathcal{B}}\|_{\infty}}^{2}}{\kappa^{2}\rho^{2}}}$ for all $i\in\mathcal{A}$ and $t\in[0,\delta)$. This proves that $\delta\in\mathcal{T}$, i.e., $\mathcal{T}$ is not empty. It is then immediate that $\mathcal{T}\supset(0,T^{*})$, where $T^{*}\coloneqq\sup\mathcal{T}>0$. Suppose, for contradiction, that $T^{*}<\infty$. Then, for any $t\in[0,T^{*})$, we have
[…]
so that by continuity of $R(t)$, we have
[…]
Again by continuity of $R(t)$, we may find a small $\delta>0$ such that $R(t)>\rho$ for $t\in[0,T^{*}+\delta)$. This shows $T^{*}+\delta\in\mathcal{T}$, contradicting the definition of $T^{*}$. Hence, we must have $T^{*}=\infty$, and this completes the proof of assertion (a). ∎
We are now ready to prove Theorem 15.
Proof of Theorem 15.
Define
$\mathcal{A}\coloneqq\{i=1,\cdots,N:\cos\theta_{i}^{0}\geq-1+\mu\}.$ (46)
We claim that
$\frac{1}{N}\sum_{i\in\mathcal{A}}(1+\cos\theta_{i}^{0})\geq\frac{2({R_{0}}-\mu)}{2-\mu}.$ (47)
To prove the claim, we first observe by definition (4) that
$\frac{1}{N}\sum_{i\in\mathcal{A}}(1+\cos\theta_{i}^{0})+\frac{1}{N}\sum_{i\in\{1,\cdots,N\}\setminus\mathcal{A}}(1+\cos\theta_{i}^{0})={R_{0}}.$ (48)
Then, using
$\begin{cases}1+\cos\theta_{i}\leq 2,&\mbox{for }i\in\mathcal{A},\\ 1+\cos\theta_{i}\leq\mu,&\mbox{for }i\in\{1,\cdots,N\}\setminus\mathcal{A},\end{cases}$ (49)
we derive
[…]
from which the above claim (47) follows.
Now let
$\rho={R_{0}}-\mu>0.$
We easily observe that condition (6) implies
$1+\mu>-\sqrt{1-\frac{\|\Omega\|_{\infty}^{2}}{\kappa^{2}\rho^{2}}}.$ (50)
We now confirm that our choice of $\mathcal{A}$ and $\rho$ along with $\mathcal{B}=\{1,\cdots,N\}$ satisfy the hypotheses of Proposition 32. Indeed, (26) follows from
[…]
(25) follows from
[…]
and (27) follows from
[…]
Assertion (a) of Theorem 15 follows from assertion (a) of Proposition 32, assertion (b) of Theorem 15 follows from assertions (d) and (f) of Proposition 32. Assertion (c) of Theorem 15 follows from assertion (c) of Proposition 32 by setting $\alpha=1-\mu$, because of (50) and $\cos\theta_{i}(t_{0})\geq-1+\mu=-\alpha$ by definition, so $\cos\theta_{i}(t)\geq 1-\mu$ for
$t\geq t_{0}+\frac{\pi}{\kappa\rho\sqrt{1-\alpha^{2}}-\|\Omega\|_{\infty}}=t_{0}+\frac{\pi}{\kappa\rho\sqrt{\mu(2-\mu)}-\|\Omega\|_{\infty}}.$
Likewise assertion (e) of Theorem 15 follows from assertion (e) of Proposition 32.
For assertion (d) of Theorem 15, note that if we denote $R_{\infty}=\lim_{t\to\infty}R(t)$, then by assertion (b) of Theorem 15 and (5) we have $\lim_{t\to\infty}\sin\theta_{i}(t)=\frac{\omega_{i}}{\kappa R_{\infty}}$, so that
$\lim_{t\to\infty}\theta_{i}(t)=\sin^{-1}\frac{\omega_{i}}{\kappa R_{\infty}}\mathrm{~or~}\pi-\sin^{-1}\frac{\omega_{i}}{\kappa R_{\infty}}\mod 2\pi.$
[…]
We are now ready to prove Corollary 16, which turns out to be special cases of Theorem 15.
Proof of Corollary 16.
Assertions (a), (b), (c), (d) and (e) of Corollary 16 just follow from the case
$\mu=\frac{3+{R_{0}}-\sqrt{{R_{0}}^{2}-2{R_{0}}+9}}{4}\in(0,\min\{R_{0},1\})$
of Theorem 15 and from
$\sqrt{R_{0}^{2}-2{R_{0}}+9}\geq 3-\frac{1}{3}{R_{0}}.$ (51)
Indeed, we need to check that (6) holds. Because of (7) and (8), it is enough to check
$K_{c}^{2}\geq\frac{1}{\rho^{2}\mu(2-\mu)}.$ (52)
We will verify this in Appendix A using elementary one-variable calculus. Also, we verify that
[…]
and we verify the time that appears in statements (c) and (d) follows from the following upper bound:
[…]
Assertions (f), (g), (h) and (i) of Corollary 16 follow from the case $\mu=\frac{4\|\Omega\|_{\infty}^{2}}{\kappa^{2}R_{0}^{2}}$ of Theorem 15. Note that because $\kappa>\frac{2\sqrt{2}\|\Omega\|_{\infty}}{R_{0}^{3/2}}$, we have $\mu<\frac{R_{0}}{2}\leq\min\{R_{0},1\}$. Condition (6) follows from $2-\mu\geq 1$ and $R_{0}-\mu>R_{0}/2$. The time that appears in statement (i) follows from the following upper bound:
$\frac{\pi}{\kappa(R_{0}-\mu)\sqrt{\mu(2-\mu)}-\|\Omega\|_{\infty}}\leq\frac{\pi}{\kappa(R_{0}-4\|\Omega\|^{2}_{\infty}/\kappa^{2}R_{0}^{2})\cdot 2\|\Omega\|_{\infty}/\kappa R_{0}-\|\Omega\|_{\infty}}=\frac{\pi}{\|\Omega\|_{\infty}\left(1-\frac{8\|\Omega\|_{\infty}^{2}}{\kappa^{2}R_{0}^{3}}\right)}.$
5. Volumetric arguments
The purpose of this section is to prove Theorems 1, 6, 7, 26, 27, 28, and 30. First, in subsection 5.1, we will recall some basic large deviations theory to provide estimates on the volume of certain singular sets and then derive Theorems 26 and 27 from Theorem 21, and Theorem 6 from Corollary 16. Then, in subsection 5.2, we will use the divergence to examine the evolution of the volume of the singular sets, to prove Theorem 1 from Theorem 15, Theorem 7 from Corollary 16, and Theorems 28 and 30 from Theorem 21.
5.1. Some large deviations theory
We first begin by discussing some time-independent large deviations theory to prove Theorems 6, 26, and 27.
In this paragraph, we recall some standard results and notations in large deviations theory (the statements of this paragraph can all be found in [vH14, Section 3.1]). For a given $\sigma>0$, a real-valued random variable $X$ is said to be $\sigma^{2}$-subgaussian if $X$ is integrable and
$\mathbb{E}[\exp(\lambda(X-\mathbb{E}X))]\leq\exp\left(\frac{\lambda^{2}\sigma^{2}}{2}\right),\quad\forall\lambda\in\mathbb{R}.$
It is clear that if the independent random variables $X$ and $Y$ are $\sigma^{2}$-subgaussian and $\tau^{2}$-subgaussian, respectively, then $X+Y$ is $(\sigma^{2}+\tau^{2})$-subgaussian. The routine Chernoff bound states that if $X$ is a $\sigma^{2}$-subgaussian random variable, then
$\mathbb{P}[X\geq\mathbb{E}X+t],\mathbb{P}[X\leq\mathbb{E}X-t]\leq\exp\left(-\frac{t^{2}}{2\sigma^{2}}\right),\quad\forall t>0.$ (53)
Thus, if $X_{1},\cdots,X_{n}$ are independent random variables that are each $\sigma^{2}$-subgaussian, then $\frac{X_{1}+\cdots+X_{n}}{n}$ is $\sigma^{2}/n$-subgaussian so that
$\mathbb{P}\left[\frac{X_{1}+\cdots+X_{n}}{n}\geq\frac{\mathbb{E}X_{1}+\cdots+\mathbb{E}X_{n}}{n}+t\right],\mathbb{P}\left[\frac{X_{1}+\cdots+X_{n}}{n}\leq\frac{\mathbb{E}X_{1}+\cdots+\mathbb{E}X_{n}}{n}-t\right]\leq\exp\left(-\frac{nt^{2}}{2\sigma^{2}}\right),\quad\forall t>0.$
In addition, the well-known Hoeffding lemma states that if $X$ is a real-valued random variable, with $a\leq X\leq b$ almost surely for some $a,b\in\mathbb{R}$, then $X$ is $(b-a)^{2}/4$-subgaussian.
With these basic tools, we are ready to prove Theorem 26.
Proof of Theorem 26.
The distribution $\mu^{\otimes N}$ means that we are independently and identically sampling $\theta_{i}^{0}\sim\mu$, so ${R_{0}}=\frac{1}{N}\sum_{j=1}^{N}I(\theta_{j}^{0})$ is the sum of the independent random variables $\frac{1}{N}I(\theta_{j}^{0})$, $j=1,\cdots,N$. By the Hoeffding lemma, $\frac{1}{N}I(\theta_{j}^{0})$ is $\frac{\|I\|_{\sup}^{2}}{4N^{2}}$-subgaussian, so by independence, ${R_{0}}=\sum_{j=1}^{N}\frac{1}{N}I(\theta_{j}^{0})$ is $\frac{\|I\|_{\sup}^{2}}{4N}$-subgaussian. As $\mathbb{E}_{\mu^{\otimes N}}{R_{0}}=R^{*}$, we conclude
$\mu^{\otimes N}\Big\{\{\theta_{i}^{0}\}_{i=1}^{N}\in[-\pi,\pi]^{N}:\Theta(t)\mbox{ satisfies }R(t)>\frac{c_{3}R^{*}}{4}~\forall t\geq 0\mbox{ and}$
$\qquad\qquad\qquad\qquad\qquad\qquad\sup_{t\geq 0}\theta_{i}(t)-\inf_{t\geq 0}\theta_{i}(t)<2\pi~\forall i=1,\cdots,N\Big\}$
[…]
$=1-\mu^{\otimes N}\left\{\{\theta_{i}^{0}\}_{i=1}^{N}\in[-\pi,\pi]^{N}:{R_{0}}\leq R^{*}-\frac{R^{*}}{2}\right\}$
[…]
If the distribution $\mu$ is such that $I(\theta^{0})$, $\theta^{0}\sim\mu$ has better subgaussian concentration, we may improve the probability estimate of Theorem 26. We will exploit better concentration behavior in Theorem 6.
Proposition 51.
Suppose $X\in[-\pi,\pi]$ is a random variable that is uniform with respect to Lebesgue measure. Then $\cos X$ is $1/2$-subgaussian.
Proof.
It is well-known that for $d\geq 2$ and a vector $a\in\mathbb{R}^{d}$,
$\int_{\mathbb{S}^{d-1}}e^{\mathrm{i}\langle a,u\rangle}d\sigma(u)=(2\pi)^{d/2}\|a\|^{-\frac{d}{2}+1}J_{\frac{d}{2}-1}(\|a\|),$
where $\sigma$ is the Lebesgue measure on $\mathbb{S}^{d-1}$, and, for $\nu\geq 0$, $J_{\nu}$ is the Bessel function of the first kind, which admits the convergent expansion
$J_{\nu}(x)=\left(\frac{x}{2}\right)^{\nu}\sum_{m=0}^{\infty}\frac{(-1)^{m}(x^{2}/4)^{m}}{m!\Gamma(\nu+m+1)}.$
Thus
$\fint_{\mathbb{S}^{1}}e^{\mathrm{i}\langle a,u\rangle}d\sigma(u)$ $=\frac{1}{2\pi}\int_{\mathbb{S}^{1}}e^{\mathrm{i}\langle a,u\rangle}d\sigma(u)$
$=J_{0}(\|a\|)$
$=\sum_{m=0}^{\infty}\frac{(-1)^{m}(\|a\|^{2}/4)^{m}}{(m!)^{2}}.$
Formally writing $a=-\mathrm{i}\lambda e_{1}$, $\lambda\in\mathbb{R}$, we have
$\fint_{-\pi}^{\pi}e^{\lambda\cos\theta}d\theta=\fint_{\mathbb{S}^{d-1}}e^{\lambda\langle e_{1},u\rangle}d\sigma(u)$ $=\sum_{m=0}^{\infty}\frac{(\lambda^{2}/4)^{m}}{(m!)^{2}}\leq\sum_{m=0}^{\infty}\frac{(\lambda^{2}/4)^{m}}{m!}=\exp\left(\frac{\lambda^{2}}{2d}\right).$
On the other hand, given a random variable $X$ and a real number $M$, if we know the concrete upper bound $X\leq M$, what can we say about how close $X$ can be to $M$? Trivially $\mathbb{E}e^{\lambda X}\leq e^{\lambda M}$ for $\lambda>0$. If the distribution of $X$ is absolutely continuous with respect to Lebesgue measure near $M$, say the density function is of the power type $O((M-x)^{\alpha-1})$ for $x<M$ near $M$ for some $\alpha>0$, then we have the bound $\mathbb{E}e^{\lambda X}\leq O(e^{\lambda M}/{\lambda^{\alpha}})+e^{\lambda(M-\Omega(1))}$ which is stronger for large $\lambda$. Conversely, if it is the case that we have the bound $\mathbb{E}e^{\lambda X}\leq\frac{C}{\lambda^{\alpha}}e^{\lambda M}$ for some constants $C,\alpha>0$, then we know the distribution of $X$ has power-type behavior near $M$, as follows; this has implications for sums of i.i.d. bounded random variables of this power-type behavior.
Lemma 52.
Let $X$ be a random variable such that for some constants $C,\alpha>0$, we have $\mathbb{E}e^{\lambda X}\leq\frac{C}{\lambda^{\alpha}}e^{\lambda M}$ for all $\lambda>0$. Then
$\mathbb{P}[X\geq M-\varepsilon]\leq C\left(\frac{e\varepsilon}{\alpha}\right)^{\alpha},\quad\varepsilon>0.$
Furthermore, if $X_{1},\cdots,X_{N}$ are independent and identically distributed copies of $X$, where $N\in\mathbb{N}$, then
$\mathbb{P}\left[\frac{X_{1}+\cdots+X_{N}}{N}\geq M-\varepsilon\right]\leq\left(C\left(\frac{e\varepsilon}{\alpha}\right)^{\alpha}\right)^{N}.$
Proof.
The first assertion follows from observing that
$\mathbb{P}[X\geq M-\varepsilon]\leq e^{-\lambda(M-\varepsilon)}\mathbb{E}e^{\lambda X}\leq\frac{Ce^{\varepsilon\lambda}}{\lambda^{\alpha}},\quad\lambda>0$
and optimizing over $\lambda$ by choosing $\lambda=\frac{\alpha}{\varepsilon}>0$. The second assertion follows from the first assertion by observing that
$\mathbb{E}\exp\left(\lambda\cdot\frac{X_{1}+\cdots+X_{N}}{N}\right)=(\mathbb{E}e^{\lambda X_{1}/N})^{N}\leq\frac{C^{N}N^{N\alpha}}{\lambda^{N\alpha}}e^{\lambda M}.$
Proposition 53.
Suppose $X\in[-\pi,\pi]$ is a random variable that is uniform with respect to Lebesgue measure. Then [^4] The correct asymptotics as $\lambda\to\infty$ is $\frac{e^{\lambda}}{\sqrt{2\pi\lambda}}(1+o(1))$.
$\mathbb{E}e^{\lambda\cos X}\leq\frac{1}{2}\sqrt{\frac{\pi}{2\lambda}}e^{\lambda},\quad\lambda>0.$
Proof.
We are to estimate the integral
$\frac{1}{2\pi}\int_{-\pi}^{\pi}e^{\lambda\cos\theta}d\theta.$
Laplace’s method tells us that this is $\frac{e^{\lambda}}{\sqrt{2\pi\lambda}}(1+o(1))$ as $\lambda\to\infty$. We content ourselves with proving a weaker bound, using the fact that $\cos\theta\leq 1-\frac{2}{\pi^{2}}\theta^{2}$ for $\theta\in[-\pi,\pi]$:
$\frac{1}{2\pi}\int_{-\pi}^{\pi}e^{\lambda\cos\theta}d\theta\leq\frac{e^{\lambda}}{2\pi}\int_{-\pi}^{\pi}e^{-2\lambda\theta^{2}/\pi^{2}}d\theta\leq\frac{e^{\lambda}}{2\pi}\int_{-\infty}^{\infty}e^{-2\lambda\theta^{2}/\pi^{2}}d\theta=\frac{1}{2}\sqrt{\frac{\pi}{2\lambda}}e^{\lambda}.$
We thus obtain the following information on the distribution of the order parameter $R$.
Lemma 54.
For $t\in(0,1)$,
$m\Big\{\{\theta_{i}^{0}\}_{i=1}^{N}\in[-\pi,\pi]^{N}:~{R_{0}}\leq t\Big\}\leq\min\left\{\exp\left(-(1-t)^{2}N\right),\left(\frac{\sqrt{\pi et}}{2}\right)^{N}\right\}.$
Proof.
Under the measure $m$, ${R_{0}}=\frac{1}{N}\sum_{j=1}^{N}(1+\cos\theta_{j}^{0})$ is the sum of the independent random variables $\frac{1}{N}(1+\cos\theta_{j}^{0})$, and as each $\frac{1}{N}(1+\cos\theta_{j}^{0})$ is $\frac{1}{2N^{2}}$-subgaussian by Proposition 51, ${R_{0}}=\sum_{j=1}^{N}\frac{1}{N}(1+\cos\theta_{j}^{0})$ is $\frac{1}{2N}$-subgaussian with mean $1$. Thus
$m\Big\{\{\theta_{i}^{0}\}_{i=1}^{N}\in[-\pi,\pi]^{N}:~{R_{0}}\leq t\Big\}\leq\exp\left(-\frac{(1-t)^{2}}{1/N}\right)=\exp\left(-(1-t)^{2}N\right).$
On the other hand, from Proposition 53 we have
$\mathbb{E}e^{-\lambda\cos\theta_{j}^{0}}=\mathbb{E}e^{\lambda\cos\theta_{j}^{0}}\leq\frac{1}{2}\sqrt{\frac{\pi}{2\lambda}}e^{\lambda},\quad\lambda>0,$
so that by Lemma 52 with $M=1$,
$\mathbb{P}\left[R_{0}\leq t\right]=\mathbb{P}\left[\frac{-\cos\theta_{1}^{0}-\cdots-\cos\theta_{N}^{0}}{N}\geq 1-t\right]\leq\left(\frac{\sqrt{\pi et}}{2}\right)^{N}.$
Proof of Theorem 6.
By Corollary 16, if $R_{0}\geq 1-\frac{\varepsilon}{5}$, since
$K_{c}\leq\frac{2}{(1-\varepsilon/5)^{3/2}}\leq 2+\varepsilon\quad(\because\varepsilon\in(0,1)),$
the condition $\kappa>(2+\varepsilon)\|\Omega\|_{\infty}$ implies
  1. (a)
    $R(t)\geq\frac{3{R_{0}}-3+\sqrt{{R_{0}}^{2}-2{R_{0}}+9}}{4}\geq\frac{\sqrt{2}}{2}-\frac{3\varepsilon}{20}$ for all $t\geq 0$,
  2. (b)
    for all $i=1,\cdots,N$, the limit $\lim_{t\to\infty}\theta_{i}(t)$ exists, and
    $\sup_{t\geq 0}\theta_{i}(t)-\inf_{t\geq 0}\theta_{i}(t)<2\pi,\quad\lim_{t\to\infty}\dot{\theta}_{i}(t)=0.$
  3. (c)
    for any $i\in\{1,\cdots,N\}$ and $t_{0}\geq 0$ with $\cos\theta_{i}(t_{0})\geq-\frac{\sqrt{2}}{2}\geq\frac{-1+{R_{0}}-\sqrt{{R_{0}}^{2}-2{R_{0}}+9}}{4}$, we have
    $\cos\theta_{i}(t)\geq\frac{-1+{R_{0}}-\sqrt{{R_{0}}^{2}-2{R_{0}}+9}}{4}\geq\frac{\sqrt{2}}{2}\quad\mathrm{for~}t\geq t_{0}+\frac{(2+\varepsilon)\pi}{\kappa-(2+\varepsilon)\|\Omega\|_{\infty}}\geq t_{0}+\frac{\pi K_{c}}{\kappa-K_{c}\|\Omega\|_{\infty}}.$
Thus
$m\Big\{\{\theta_{i}^{0}\}_{i=1}^{N}\in[-\pi,\pi]^{N}:~\forall t\geq 0~R(t)\geq\frac{1}{\sqrt{2}}-\frac{3\varepsilon}{20},$
$\qquad\forall i=1,\cdots,N~\exists\lim_{t\to\infty}\theta_{i}(t),~\sup_{t\geq 0}\theta_{i}(t)-\inf_{t\geq 0}\theta_{i}(t)<2\pi,~\lim_{t\to\infty}\dot{\theta}_{i}(t)=0$
$\qquad\mathrm{and~}\forall i\in\{1,\cdots,N\}\forall t_{0}\geq 0\mathrm{~if~}\cos\theta_{i}(t_{0})\geq-\frac{1}{\sqrt{2}}\mathrm{~then}$
$\qquad\cos\theta_{i}(t_{0})\geq-\frac{1}{\sqrt{2}}\mathrm{~for~}t\geq t_{0}\mathrm{~and~}\cos\theta_{i}(t)\geq\frac{1}{\sqrt{2}}\mathrm{~for~}t\geq t_{0}+\frac{(2+\varepsilon)\pi}{\kappa-(2+\varepsilon)\|\Omega\|_{\infty}}\Big\}$
$\geq m\Big\{\{\theta_{i}^{0}\}_{i=1}^{N}\in[-\pi,\pi]^{N}:~{R_{0}}\geq 1-\frac{\varepsilon}{5}\Big\}.$
On the other hand, by Lemma 54,
$m\left\{\{\theta_{i}^{0}\}_{i=1}^{N}\in[-\pi,\pi]^{N}:~{R_{0}}\geq 1-\frac{\varepsilon}{5}\right\}=1-m\left\{\{\theta_{i}^{0}\}_{i=1}^{N}\in[-\pi,\pi]^{N}:~{R_{0}}<1-\frac{\varepsilon}{5}\right\}\geq 1-\exp\left(-\frac{\varepsilon^{2}N}{25}\right).$
Similarly, if $\kappa>\left(\frac{\pi e}{2}\right)^{3/2}\|\Omega\|_{\infty}>2\sqrt{2}\|\Omega\|_{\infty}$ and $R_{0}>\left(\frac{2\sqrt{2}\|\Omega\|_{\infty}}{\kappa}\right)^{2/3}$, then by Corollary 16,
  1. (1)
    For all $i=1,\cdots,N$, the limit $\lim_{t\to\infty}\theta_{i}(t)$ exists, and
    $\sup_{t\geq 0}\theta_{i}(t)-\inf_{t\geq 0}\theta_{i}(t)<2\pi,\quad\lim_{t\to\infty}\dot{\theta}_{i}(t)=0.$
  2. (2)
    for any $i\in\{1,\cdots,N\}$ and $t_{0}\geq 0$ with $\cos\theta_{i}(t_{0})\geq-1+\frac{\|\Omega\|_{\infty}^{2/3}}{\kappa^{2/3}}\geq-1+\frac{4\|\Omega\|_{\infty}^{2}}{\kappa^{2}R_{0}^{2}}$, we have
    $\cos\theta_{i}(t)\geq 1-\frac{4\|\Omega\|_{\infty}^{2}}{\kappa^{2}R_{0}^{2}}\geq 1-\frac{\|\Omega\|_{\infty}^{2/3}}{\kappa^{2/3}}\quad\mathrm{for~}t\geq t_{0}+\frac{\pi}{\|\Omega\|_{\infty}\left(1-\frac{8\|\Omega\|_{\infty}^{2}}{\kappa^{2}R_{0}^{3}}\right)}\geq t_{0}+\frac{\pi K_{c}}{\kappa-K_{c}\|\Omega\|_{\infty}}.$
So by Lemma 54,
$m\Big\{\{\theta_{i}^{0}\}_{i=1}^{N}\in[-\pi,\pi]^{N}:\exists\lim_{t\to\infty}\theta_{i}(t),~\sup_{t\geq 0}\theta_{i}(t)-\inf_{t\geq 0}\theta_{i}(t)<2\pi,~\lim_{t\to\infty}\dot{\theta}_{i}(t)=0~\mbox{for all }i=1,\cdots,N\Big\}$
$\geq m\left\{\{\theta_{i}^{0}\}_{i=1}^{N}\in[-\pi,\pi]^{N}:~R_{0}>\left(\frac{2\sqrt{2}\|\Omega\|_{\infty}}{\kappa}\right)^{2/3}\right\}\geq 1-\left(\sqrt{\frac{\pi e}{2}}\frac{\|\Omega\|_{\infty}^{1/3}}{\kappa^{1/3}}\right)^{N}.$
Proof of Theorem 27.
The distribution $\mu^{\otimes N}$ means that we are independently and identically sampling $\theta_{i}^{0}\sim\mu$, so ${R_{0}}=\frac{1}{N}\sum_{j=1}^{N}I(\theta_{j}^{0})$ is the sum of the independent random variables $\frac{1}{N}I(\theta_{j}^{0})$, $j=1,\cdots,N$. By (17) and Lemma 52 with $M=0$, we have that
$\mu^{\otimes N}\left\{\{\theta_{i}^{0}\}_{i=1}^{N}\in[-\pi,\pi]^{N}:{R_{0}}=\frac{1}{N}\sum_{j=1}^{N}I(\theta_{j}^{0})\leq\varepsilon\right\}\leq\left(C_{\mu}\left(\frac{e\varepsilon}{\beta}\right)^{\beta}\right)^{N}.$ (54)
By condition (18), we may find $\varepsilon>0$ such that $C_{\mu}\left(\frac{e\varepsilon}{\beta}\right)^{\beta}<1$,
$\varepsilon<\max\left\{\left(\frac{2}{c_{1}c_{3}}\frac{\|\Omega\|_{\infty}}{\kappa}\right)^{\frac{1}{1+p/q}}(2c_{2})^{\frac{p/q}{1+p/q}},\frac{2}{c_{1}c_{3}(\pi-\alpha_{0})^{p}}\frac{\|\Omega\|_{\infty}}{\kappa}\right\}$ (55)
and
$\kappa>\max\left\{\frac{2(2c_{2})^{p/q}}{c_{1}c_{3}}\cdot\frac{\|\Omega\|_{\infty}}{{\varepsilon}^{1+\frac{p}{q}}},\frac{2}{c_{1}c_{3}(\pi-\alpha_{0})^{p}}\cdot\frac{\|\Omega\|_{\infty}}{{\varepsilon}}\right\}.$
By Theorem 21, we have
$\mu^{\otimes N}\Big\{\{\theta_{i}^{0}\}_{i=1}^{N}\in[-\pi,\pi]^{N}:\sup_{t\geq 0}\theta_{i}(t)-\inf_{t\geq 0}\theta_{i}(t)<2\pi~\forall i=1,\cdots,N\Big\}$
$\geq\mu^{\otimes N}\left\{\{\theta_{i}^{0}\}_{i=1}^{N}\in[-\pi,\pi]^{N}:{R_{0}}=\frac{1}{N}\sum_{j=1}^{N}I(\theta_{j}^{0})>\varepsilon\right\}$
$\geq 1-\mu^{\otimes N}\left\{\{\theta_{i}^{0}\}_{i=1}^{N}\in[-\pi,\pi]^{N}:{R_{0}}=\frac{1}{N}\sum_{j=1}^{N}I(\theta_{j}^{0})\leq\varepsilon\right\}$
and combining (54) and (55) gives the stated result. ∎
The proof of Theorems 1, 7, 28, and 30 will be more involved since their proof require asking for “how much” initial data $\{\theta_{i}^{0}\}_{i=1}^{N}$ does the order parameter stay small up to time $T$. This is the subject of the next subsection.
5.2. Instability of states with small order parameter
In this subsection, we prove Theorems 1, 7, 28, and 30. We will control the volume of a singular set using the divergence of the flows (1) and (3); the argument of this section is inspired by Section 6 of [HR20].
5.2.1. The prototypical Winfree model (3)
In this subsubsection, we prove Theorems 1 and 7. The idea is that the set of states with sufficiently small order parameters cannot be positively invariant under the Winfree flow (3). By the identification $\theta_{i}\leftrightarrow e^{\mathrm{i}\theta_{i}}$, we may consider (3) to be a flow on the $N$-torus $\mathbb{T}^{N}$.
For any $\delta>0$ and $N$, we define the set
$U^{0}_{1-\delta}\coloneqq\left\{\Theta\in\mathbb{T}^{N}:~R(\Theta)<1-\delta\right\}.$
We will show that $U^{0}_{1-\delta}$ is almost surely not positively invariant under the Winfree flow (3) with $\kappa>0$, i.e., if the flow starts at a generic point in $U^{0}_{1-\delta}$, it cannot stay inside the region for all time, regardless of the magnitude of the natural frequencies. Then, as an application of this result, we will prove our main Theorem 1.
First, we begin with a heuristic argument for the fact that the set $U^{0}_{1-\delta}$ is not positively invariant. For a given natural frequency vector $\Omega$, consider the following two cases:
  • Case A ($\kappa\gg\|\Omega\|_{\infty}$): According to numerical simulations, the flow converges to the equilibrium of Theorem 23, and so the order parameter must eventually exceed $1+\cos\frac{\pi}{3}=\frac{3}{2}>1$.
  • Case B ($\kappa\ll\|\Omega\|_{\infty}$): In this case, the nonlinear coupling will contribute little to the dynamics of (3), so we may approximate (3) as
    $\dot{\theta}_{i}\approx\omega_{i}.$
    Thus, the phases will tend to be uniformly randomly distributed over the unit circle and thus the order parameter will attain the temporal average
    $\langle R\rangle=\frac{1}{N}\sum_{i}\langle 1+\cos\theta_{i}\rangle=1,$
    where the temporal average $\langle\cdot\rangle$ is defined by
    $\langle R\rangle\coloneqq\lim_{T\rightarrow\infty}\frac{1}{T}\int_{0}^{T}R(t)dt.$
    Thus, for any $\delta>0$, such a solution cannot stay in the region $U^{0}_{1-\delta}$ for all times $t\geq 0$.
Hence, it is reasonable to expect that given $\kappa>0$, $\Omega=(\omega_{1},\cdots,\omega_{N})$ and any $\delta>0$, a generic solution to (3) cannot stay in the region $U^{0}_{1-\delta}$ for all $t\geq 0$. Now, we will prove this heuristic argument rigorously using the divergence of the flow (3). First, note that the Winfree model gives the integral curves to the following vector field:
$F:\mathbb{T}^{N}\rightarrow T\mathbb{T}^{N},\quad F=(F_{1},\cdots,F_{N}),$ (56)
where
$F_{i}(\Theta)=\omega_{i}-\frac{\kappa}{N}\sum_{j=1}^{N}(1+\cos\theta_{j})\sin\theta_{i}.$ (57)
Let us denote the flow of this vector field by $\Phi_{t}$, i.e., $\Phi_{t}(\Theta^{0})=\Theta(t)$ is the solution to (32) with initial data $\Theta^{0}$. Also, for any subset $U\subset\mathbb{T}^{N}$ we denote
$\Phi_{t}(U)=\{\Phi_{t}(\Theta):\Theta\in U\}.$
Finally, we endow $\mathbb{T}^{N}$ with the standard Lebesgue measure $m$, normalized so that $m(\mathbb{T}^{N})=1$. Now the divergence of the vector field $F$ is
$\nabla\cdot F=\sum_{i=1}^{N}\frac{\partial F_{i}}{\partial\theta_{i}}=-\frac{\kappa}{N}\sum_{i,j=1,\cdots,N}(1+\cos\theta_{j})\cos\theta_{i}+\frac{\kappa}{N}\sum_{i=1}^{N}\sin^{2}\theta_{i}=\kappa\left(NR(1-R)+\frac{1}{N}\sum_{i=1}^{N}\sin^{2}\theta_{i}\right)$ (58)
Note that the divergence is independent of the choice of natural frequencies $\nu_{i}$, and depends only on $\kappa$, $N$ and $\Theta$.
With (58), we can immediately show the instability of equilibria in $U^{0}_{1-\delta}$.
Proposition 55.
The following two assertions hold.
  1. (a)
    If $\kappa>0$, then any equilibrium $\Theta$ of the Winfree model (3) with order parameter $R$ satisfying
    $NR(1-R)+\frac{\sum_{i}\omega_{i}^{2}}{N\kappa^{2}R^{2}}>0$
    is linearly unstable. In particular, any equilibrium with $0<R<1$ is linearly unstable.
  2. (b)
    If $\kappa<0$, then any equilibrium $\Theta$ of the Winfree model (3) with order parameter $R$ satisfying
    $NR(1-R)+\frac{\sum_{i}\omega_{i}^{2}}{N\kappa^{2}R^{2}}<0$
    is linearly unstable.
Proof.
If $\Theta$ is an equilibrium with $\nabla\cdot F(\Theta)>0$, then the Jacobian matrix of $F$ at $\Theta$ has a positive trace and thus has a complex eigenvalue with positive real part, so that $\Theta$ is an unstable equilibrium. Note that for an equilibrium $\Theta$, we have $\omega_{i}=\kappa R\sin\theta_{i}$, so that
$\nabla\cdot F(\Theta)=\kappa\left(NR(1-R)+\frac{1}{N\kappa^{2}R^{2}}\sum_{i=1}^{N}\omega_{i}^{2}\right).$
More generally, we can rigorously prove the heuristic argument in the beginning of this subsubsection.
Proposition 56.
Let $\kappa>0$, $N\geq 2$, [^5] If $N=1$, the dynamics of (3) is extremely well-understood; see Proposition 46. and $\Omega=(\omega_{1},\cdots,\omega_{N})$ be fixed, and consider the flow $\Phi_{t}$ generated by (56) and (57). Then, for any $0<\delta<1$, the Borel set
$U_{1-\delta}^{\infty}=\{\Theta^{0}\in\mathbb{T}^{N}:R(\Phi_{t}(\Theta^{0}))<1-\delta\mbox{ for all }t\geq 0\}$
has measure zero. Also, for each $T\geq 0$, the open set
$U_{1-\delta}^{T}\coloneqq\{\Theta^{0}\in\mathbb{T}^{N}:R(\Phi_{t}(\Theta^{0}))<1-\delta\mbox{ for all }0\leq t\leq T\}$
has measure
$m(U_{1-\delta}^{T})\leq\begin{cases}\frac{1}{N/2+1}\left(\frac{\sqrt{\pi e\delta}}{2}\right)^{N}\left(1-\exp\left(-\kappa N\delta(1-\delta)T\right)\right)+\exp\left(-N\delta^{2}-\kappa N\delta(1-\delta)T\right),&\mathrm{if~}0<\delta<\frac{1}{4}\mathrm{~and~}T\leq T_{0},\\ \left(\frac{4}{\pi e\delta}+\frac{4\kappa(1-\delta)}{\pi e}(T-T_{0})\right)^{-\frac{N}{2}},&\mathrm{if~}0<\delta<\frac{1}{4}\mathrm{~and~}T>T_{0},\\ \frac{1}{N/2+1}\left(\frac{\sqrt{\pi e}}{4}\right)^{N}\left(1-\exp\left(-\frac{3\kappa NT}{16}\right)\right)+\exp\left(-N\delta^{2}-\frac{3\kappa NT}{16}\right),&\mathrm{if~}\frac{1}{4}\leq\delta<\frac{1}{2}\mathrm{~and~}T\leq T_{0},\\ \left(\frac{16}{\pi e}+\frac{3\kappa}{\pi e}(T-T_{0})\right)^{-\frac{N}{2}},&\mathrm{if~}\frac{1}{4}\leq\delta<\frac{1}{2}\mathrm{~and~}T>T_{0},\\ \left(e^{2\delta^{2}}+\frac{4\kappa\delta T}{\pi e}\right)^{-\frac{N}{2}},&\mathrm{if~}\frac{1}{2}\leq\delta<\frac{3}{4},\\ \left(\frac{4}{\pi e(1-\delta)}+\frac{4\kappa\delta T}{\pi e}\right)^{-\frac{N}{2}},&\mathrm{if~}\frac{3}{4}\leq\delta<1,\end{cases}$ (59)
where
$T_{0}=\begin{cases}\frac{1}{\kappa N\delta(1-\delta)}\log\left(\left(1+\frac{2}{N}\right)\left(\frac{2}{\sqrt{\pi\delta}e^{\frac{1}{2}+\delta^{2}}}\right)^{N}-\frac{2}{N}\right)&\mathrm{if~}0<\delta<\frac{1}{4},\\ \frac{16}{3\kappa N}\log\left(\left(1+\frac{2}{N}\right)\left(\frac{4}{\sqrt{\pi}e^{\frac{1}{2}+\delta^{2}}}\right)^{N}-\frac{2}{N}\right)&\mathrm{if~}\frac{1}{4}\leq\delta<\frac{1}{2}.\end{cases}$
Proof.
Note that
$U_{\varepsilon}^{0}=\{\Theta^{0}\in\mathbb{T}^{N}:R(\Theta^{0})<\varepsilon\},\quad\varepsilon\in(0,1),$
which by Lemma 54 has measure
$m(U_{\varepsilon}^{0})\leq\left(\frac{\sqrt{\pi e\varepsilon}}{2}\right)^{N},\quad\varepsilon\in(0,1).$
By definition, we have
$\Phi_{t}(U_{1-\delta}^{T})\subset U_{1-\delta}^{0},\quad\forall~0\leq t\leq T.$
By the fact that the divergence of a flow determines the rate of change of volume, we have for $0\leq t\leq T$
[…]
We now prove (59) case by case.
  1. (1)
    Suppose $0<\delta<\frac{1}{4}$. We have
[…]
    $\geq\kappa N\delta(1-\delta)m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\setminus U^{0}_{\delta}\right)+\kappa N(1-\delta)\int_{\Phi_{t}\left(U_{1-\delta}^{T}\right)\cap U^{0}_{\delta}}Rdm(\Theta),$
    which we estimate term by term:
    $\kappa N\delta(1-\delta)m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\setminus U^{0}_{\delta}\right)$ $\geq\kappa N\delta(1-\delta)m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\right)-\kappa N\delta(1-\delta)m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\cap U^{0}_{\delta}\right)$
    and
    $\kappa N(1-\delta)\int_{\Phi_{t}\left(U_{1-\delta}^{T}\right)\cap U^{0}_{\delta}}Rdm(\Theta)$ $=\kappa N(1-\delta)\int_{\Phi_{t}\left(U_{1-\delta}^{T}\right)\cap U^{0}_{\delta}}\int_{0}^{\delta}\mathbbm{1}_{R\geq t}dtdm(\Theta)$
    $=\kappa N(1-\delta)\int_{0}^{\delta}m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\cap U^{0}_{\delta}\setminus U^{0}_{t}\right)dt$
    $\geq\kappa N(1-\delta)\int_{0}^{\delta}\max\left\{0,m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\cap U^{0}_{\delta}\right)-m\left(U^{0}_{t}\right)\right\}dt$
    $\geq\kappa N(1-\delta)\int_{0}^{\delta}\max\left\{0,m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\cap U^{0}_{\delta}\right)-\left(\frac{\sqrt{\pi et}}{2}\right)^{N}\right\}dt$
    $=\kappa N(1-\delta)\frac{4}{\pi e}\frac{1}{1+2/N}m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\cap U^{0}_{\delta}\right)^{1+\frac{2}{N}},$
    where in the last equality we used $m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\cap U^{0}_{\delta}\right)\leq m\left(U^{0}_{\delta}\right)\leq\left(\frac{\sqrt{\pi e\delta}}{2}\right)^{N}$. Harvesting estimates, we obtain
    $\begin{aligned} \frac{d}{dt}m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\right)&\geq\kappa N\delta(1-\delta)m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\right)-\kappa N\delta(1-\delta)m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\cap U^{0}_{\delta}\right)\\ &\quad+\kappa N(1-\delta)\frac{4}{\pi e}\frac{1}{1+2/N}m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\cap U^{0}_{\delta}\right)^{1+\frac{2}{N}}\\ &\geq\kappa N\delta(1-\delta)m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\right)+\kappa N(1-\delta)\inf_{0\leq y\leq m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\right)}\left(-\delta y+\frac{4}{\pi e}\frac{1}{1+2/N}y^{1+\frac{2}{N}}\right)\\ &=\begin{cases}\kappa N(1-\delta)\frac{4}{\pi e}\frac{1}{1+2/N}m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\right)^{1+2/N}&\mathrm{if~}m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\right)\leq\left(\frac{\sqrt{\pi e\delta}}{2}\right)^{N}\\ \kappa N\delta(1-\delta)\left(m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\right)-\frac{1}{N/2+1}\left(\frac{\sqrt{\pi e\delta}}{2}\right)^{N}\right)&\mathrm{if~}m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\right)>\left(\frac{\sqrt{\pi e\delta}}{2}\right)^{N}\end{cases}\\ &\geq\begin{cases}\frac{2\kappa N(1-\delta)}{\pi e}m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\right)^{1+2/N}&\mathrm{if~}m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\right)\leq\left(\frac{\sqrt{\pi e\delta}}{2}\right)^{N}\\ \kappa N\delta(1-\delta)\left(m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\right)-\frac{1}{N/2+1}\left(\frac{\sqrt{\pi e\delta}}{2}\right)^{N}\right)&\mathrm{if~}m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\right)>\left(\frac{\sqrt{\pi e\delta}}{2}\right)^{N}\end{cases}\end{aligned}$ (61)
    where in the second inequality we put $y=m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\cap U^{0}_{\delta}\right)$, and in the first equality we optimized to find $y=m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\right)$ if $m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\right)\leq\left(\frac{\sqrt{\pi e\delta}}{2}\right)^{N}$ and $y=\left(\frac{\sqrt{\pi e\delta}}{2}\right)^{N}$ if $m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\right)>\left(\frac{\sqrt{\pi e\delta}}{2}\right)^{N}$, and in the third inequality we used $N\geq 2$.
    Solving this differential inequality (61) along with the terminal condition
[…]
    gives the stated inequality (59) in the case $0<\delta<\frac{1}{4}$.
  2. (2)
    On the other hand, suppose $\frac{1}{4}\leq\delta<\frac{1}{2}$. We have
[…]
    $\geq\frac{3}{16}\kappa Nm\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\setminus U^{0}_{1/4}\right)+\frac{3}{4}\kappa N\int_{\Phi_{t}\left(U_{1-\delta}^{T}\right)\cap U^{0}_{1/4}}Rdm(\Theta),$
    which we again estimate term by term:
    $\frac{3}{16}\kappa Nm\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\setminus U^{0}_{1/4}\right)$ $\geq\frac{3}{16}\kappa Nm\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\right)-\frac{3}{16}\kappa Nm\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\cap U^{0}_{1/4}\right)$
    and
    $\frac{3}{4}\kappa N\int_{\Phi_{t}\left(U_{1-\delta}^{T}\right)\cap U^{0}_{1/4}}Rdm(\Theta)$ $=\frac{3}{4}\kappa N\int_{\Phi_{t}\left(U_{1-\delta}^{T}\right)\cap U^{0}_{1/4}}\int_{0}^{1/4}\mathbbm{1}_{R\geq t}dtdm(\Theta)$
    $=\frac{3}{4}\kappa N\int_{0}^{1/4}m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\cap U^{0}_{1/4}\setminus U^{0}_{t}\right)dt$
    $\geq\frac{3}{4}\kappa N\int_{0}^{1/4}\max\left\{0,m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\cap U^{0}_{1/4}\right)-m\left(U^{0}_{t}\right)\right\}dt$
    $\geq\frac{3}{4}\kappa N\int_{0}^{1/4}\max\left\{0,m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\cap U^{0}_{1/4}\right)-\left(\frac{\sqrt{\pi et}}{2}\right)^{N}\right\}dt$
    $=\frac{3\kappa N}{\pi e}\frac{1}{1+2/N}m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\cap U^{0}_{1/4}\right)^{1+\frac{2}{N}},$
    where in the last equality we used $m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\cap U^{0}_{1/4}\right)\leq m\left(U^{0}_{1/4}\right)\leq\left(\frac{\sqrt{\pi e}}{4}\right)^{N}$.
    Harvesting estimates, we obtain
    $\begin{aligned} \frac{d}{dt}m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\right)&\geq\frac{3}{16}\kappa Nm\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\right)-\frac{3}{16}\kappa Nm\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\cap U^{0}_{1/4}\right)\\ &\quad+\frac{3\kappa N}{\pi e}\frac{1}{1+2/N}m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\cap U^{0}_{1/4}\right)^{1+\frac{2}{N}}\\ &\geq\frac{3}{16}\kappa Nm\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\right)+\frac{3\kappa N}{16}\inf_{0\leq y\leq m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\right)}\left(-y+\frac{16}{\pi e}\frac{1}{1+2/N}y^{1+\frac{2}{N}}\right)\\ &=\begin{cases}\frac{3\kappa N}{\pi e}\frac{1}{1+2/N}m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\right)^{1+2/N}&\mathrm{if~}m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\right)\leq\left(\frac{\sqrt{\pi e}}{4}\right)^{N}\\ \frac{3\kappa N}{16}\left(m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\right)-\frac{1}{N/2+1}\left(\frac{\sqrt{\pi e}}{4}\right)^{N}\right)&\mathrm{if~}m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\right)>\left(\frac{\sqrt{\pi e}}{4}\right)^{N}\end{cases}\\ &\geq\begin{cases}\frac{3\kappa N}{2\pi e}m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\right)^{1+2/N}&\mathrm{if~}m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\right)\leq\left(\frac{\sqrt{\pi e}}{4}\right)^{N}\\ \frac{3\kappa N}{16}\left(m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\right)-\frac{1}{N/2+1}\left(\frac{\sqrt{\pi e}}{4}\right)^{N}\right)&\mathrm{if~}m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\right)>\left(\frac{\sqrt{\pi e}}{4}\right)^{N}\end{cases}\end{aligned}$ (62)
    where in the second inequality we put $y=m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\cap U^{0}_{1/4}\right)$, and in the first equality we optimized to find $y=m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\right)$ if $m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\right)\leq\left(\frac{\sqrt{\pi e}}{4}\right)^{N}$ and $y=\left(\frac{\sqrt{\pi e}}{4}\right)^{N}$ if $m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\right)>\left(\frac{\sqrt{\pi e}}{4}\right)^{N}$, and in the third inequality we used $N\geq 2$.
    Solving this differential inequality (62) along with the terminal condition
[…]
    gives the stated inequality (59) in the case $\frac{1}{4}\leq\delta<\frac{1}{2}$.
  3. (3)
    Now, suppose $\frac{1}{2}\leq\delta<1$. Evidently $\Phi_{t}\left(U_{1-\delta}^{T}\right)\subset U_{1-\delta}^{0}$, so
[…]
    where in the last equality we used $m\left(\Phi_{t}\left(U_{1-\delta}^{T}\right)\right)\leq m\left(U_{1-\delta}^{0}\right)\leq\left(\frac{\sqrt{\pi e(1-\delta)}}{2}\right)^{N}$ and in the last inequality we used $N\geq 2$.
[…]
It is now evident that
$U_{1-\delta}^{\infty}=\mathop{\bigcap}_{T\in\mathbb{N}}U_{1-\delta}^{T}=\mathop{\bigcap}_{T>0}U_{1-\delta}^{T}$
is a Borel set with measure $m\left(U_{1-\delta}^{\infty}\right)\leq\liminf_{T\to\infty}m\left(U_{1-\delta}^{T}\right)=0$. ∎
We are now ready to prove Theorem 1.
Proof of Theorem 1.
Choose a sequence $\{\mu_{n}\}_{n=1}^{\infty}$ of real numbers in $\left(1-\sqrt{\frac{1}{2}+\sqrt{\frac{1}{4}-\frac{\|\Omega\|_{\infty}^{2}}{\kappa^{2}}}},1-\frac{1}{\sqrt{2}}\right)$ such that $\lim_{n\to\infty}\mu_{n}=1-\sqrt{\frac{1}{2}+\sqrt{\frac{1}{4}-\frac{\|\Omega\|_{\infty}^{2}}{\kappa^{2}}}}$. As
$\kappa>\frac{\|\Omega\|_{\infty}}{(1-\mu_{n})\sqrt{\mu_{n}(2-\mu_{n})}},$
we may choose a sequence $\{\delta_{n}\}_{n=1}^{\infty}$ of real numbers in $(0,1)$ sufficiently close to 0 such that
$\kappa>\frac{\|\Omega\|_{\infty}}{(1-\delta_{n}-\mu_{n})\sqrt{\mu_{n}(2-\mu_{n})}}\quad\forall n\geq 1,\quad\mbox{and}\quad\lim_{n\rightarrow\infty}\delta_{n}=0.$
Since $\kappa>0$, Proposition 56 tells us that for every $n\geq 1$ the set
$U_{1-\delta_{n}}^{\infty}=\{\Theta^{0}\in\mathbb{T}^{N}:R(\Phi_{t}(\Theta^{0}))<1-\delta_{n}\mbox{ for all }t\geq 0\}$
has measure zero and thus the union
$U^{\infty}=\mathop{\bigcup}_{n=1}^{\infty}U_{1-\delta_{n}}^{\infty}$
[…]
$R(\Phi_{t_{n}}(\Theta^{0}))\geq 1-\delta_{n},$
and we apply Theorem 15 at this time $t=t_{n}$ to conclude the following:
  1. (a)
    By Theorem 15 (b), there exists $\lim_{t\to\infty}\theta_{i}(t)$ and $\lim_{t\to\infty}\dot{\theta}_{i}(t)=0$. This proves assertion (a) of Theorem 1.
  2. (b)
    By Theorem 15 (a), we have $\lim_{t\to\infty}R(t)\geq 1-\delta_{n}-\mu_{n}$. By taking the limit $n\rightarrow\infty$ we obtain assertion (b) of Theorem 1.
We are also ready to prove Theorem 7.
Proof of Theorem 7.
Similarly to the proof of Theorem 6, we have from Corollary 16 that
$m\Big\{\{\theta_{i}^{0}\}_{i=1}^{N}\in[-\pi,\pi]^{N}:~\forall t\geq T~R(t)\geq\frac{1}{\sqrt{2}}-\frac{3\varepsilon}{20},$
$\qquad\forall i=1,\cdots,N~\exists\lim_{t\to\infty}\theta_{i}(t),~\sup_{t\geq T}\theta_{i}(t)-\inf_{t\geq T}\theta_{i}(t)<2\pi,~\lim_{t\to\infty}\dot{\theta}_{i}(t)=0$
$\qquad\mathrm{and~}\forall i\in\{1,\cdots,N\}\forall t_{0}\geq T\mathrm{~if~}\cos\theta_{i}(t_{0})\geq-\frac{1}{\sqrt{2}}\mathrm{~then}$
$\qquad\cos\theta_{i}(t_{0})\geq-\frac{1}{\sqrt{2}}\mathrm{~for~}t\geq t_{0}\mathrm{~and~}\cos\theta_{i}(t)\geq\frac{1}{\sqrt{2}}\mathrm{~for~}t\geq t_{0}+\frac{(2+\varepsilon)\pi}{\kappa-(2+\varepsilon)\|\Omega\|_{\infty}}\Big\}$
$\geq m\Big\{\{\theta_{i}^{0}\}_{i=1}^{N}\in[-\pi,\pi]^{N}:~R(t)>1-\frac{\varepsilon}{5}\mathrm{~for~some~}t\in[0,T]\Big\}$
$=1-m\left(U_{1-\varepsilon/5}^{T}\right)$
at which point we may invoke Proposition 56 with $\delta=\frac{\varepsilon}{5}<\frac{1}{4}$ to see that
$m(U_{1-\varepsilon/5}^{T})\leq\begin{cases}\frac{1}{N/2+1}\left(\frac{\sqrt{\pi e\varepsilon}}{2\sqrt{5}}\right)^{N}\left(1-\exp\left(-\frac{\kappa N\varepsilon(5-\varepsilon)T}{25}\right)\right)+\exp\left(-\frac{N\varepsilon^{2}}{25}-\frac{\kappa N\varepsilon(5-\varepsilon)T}{25}\right),&\mathrm{if~}T\leq T_{0},\\ \left(\frac{20}{\pi e\varepsilon}+\frac{4\kappa(5-\varepsilon)}{5\pi e}(T-T_{0})\right)^{-\frac{N}{2}},&\mathrm{if~}T>T_{0},\end{cases}$
where $T_{0}=\frac{25}{\kappa N\varepsilon(5-\varepsilon)}\log\left(\left(1+\frac{2}{N}\right)\left(\frac{2\sqrt{5}}{\sqrt{\pi\varepsilon}e^{1/2+\varepsilon^{2}/25}}\right)^{N}-\frac{2}{N}\right)$.
Similarly, if $\kappa>8\|\Omega\|_{\infty}$ and $R_{0}>\left(\frac{2\sqrt{2}\|\Omega\|_{\infty}}{\kappa}\right)^{2/3}$, we have from Corollary 16 that
$m\Big\{\{\theta_{i}^{0}\}_{i=1}^{N}\in[-\pi,\pi]^{N}:\exists\lim_{t\to\infty}\theta_{i}(t),~\sup_{t\geq T}\theta_{i}(t)-\inf_{t\geq T}\theta_{i}(t)<2\pi,~\lim_{t\to\infty}\dot{\theta}_{i}(t)=0~\mbox{for all }i=1,\cdots,N\Big\}$
$\geq m\left\{\{\theta_{i}^{0}\}_{i=1}^{N}\in[-\pi,\pi]^{N}:~R(t)>\left(\frac{2\sqrt{2}\|\Omega\|_{\infty}}{\kappa}\right)^{2/3}\mathrm{~for~some~}t\in[0,T]\right\}=1-m\left(U^{T}_{\left(\frac{2\sqrt{2}\|\Omega\|_{\infty}}{\kappa}\right)^{2/3}}\right)$
and we invoke Proposition 56 to see that
$m\left(U^{T}_{\left(\frac{2\sqrt{2}\|\Omega\|_{\infty}}{\kappa}\right)^{2/3}}\right)\leq\begin{cases}\left(e^{1/2}+\frac{3\kappa T}{\pi e}\right)^{-\frac{N}{2}}&\mathrm{if~}8\|\Omega\|_{\infty}<\kappa\leq 16\sqrt{2}\|\Omega\|_{\infty}\\ \left(\frac{4}{\pi e}\left(\frac{\kappa}{2\sqrt{2}\|\Omega\|_{\infty}}\right)^{2/3}+\frac{3\kappa T}{\pi e}\right)^{-\frac{N}{2}}&\mathrm{if~}\kappa>16\sqrt{2}\|\Omega\|_{\infty}\end{cases}$
since $\left(\frac{2\sqrt{2}\|\Omega\|_{\infty}}{\kappa}\right)^{2/3}<\frac{1}{2}$ if $\kappa>8\|\Omega\|_{\infty}$ and $\left(\frac{2\sqrt{2}\|\Omega\|_{\infty}}{\kappa}\right)^{2/3}<\frac{1}{4}$ if $\kappa>16\sqrt{2}\|\Omega\|_{\infty}$. ∎
5.2.2. General interaction functions $I$ and $S$
In this subsection we prove Theorems 28 and 30. This time, note that the Winfree model (1) gives the integral curves to the following vector field:
$F^{I,S}:\mathbb{T}^{N}\rightarrow T\mathbb{T}^{N},\quad F^{I,S}=(F^{I,S}_{1},\cdots,F^{I,S}_{N}),$ (64)
where
$F^{I,S}_{i}(\Theta)=\omega_{i}+\frac{\kappa}{N}\sum_{j=1}^{N}I(\theta_{j})S(\theta_{i}).$ (65)
Denote the flow of this vector field by $\Phi^{I,S}_{t}$. The divergence of the vector field $F^{I,S}$ satisfies, under the assumptions (19), (20), and $I\geq 0$,
$\nabla\cdot F^{I,S}=\sum_{i=1}^{N}\frac{\partial F^{I,S}_{i}}{\partial\theta_{i}}=\frac{\kappa}{N}\sum_{i,j=1,\cdots,N}I(\theta_{j})S^{\prime}(\theta_{i})+\frac{\kappa}{N}\sum_{i=1}^{N}I^{\prime}(\theta_{i})S(\theta_{i})\geq c_{4}\kappa NR(I_{*}-R).$ (66)
Again, the right-hand side is independent of the choice of natural frequencies $\nu_{i}$, and depends only on $\kappa$, $N$ and $R$.
With (66), we can again show the instability of certain equilibria.
Proposition 57.
Consider the Winfree model (1) with interaction functions $I$ and $S$ satisfying (19), (20) and $I\geq 0$. If $\kappa>0$, then any equilibrium $\Theta$ of (1) with order parameter $R\in(0,I_{*})$ is linearly unstable.
The proof is the same as that of Proposition 55.
Again, we have the following generalization of Proposition 56, whose proof is the same.
Proposition 58.
Consider the Winfree model (1) with interaction functions $I$ and $S$ satisfying (19), (20), and (21). Let $\kappa>0$, $N$, and $\Omega=(\omega_{1},\cdots,\omega_{N})$ be fixed, and consider the flow $\Phi_{t}^{I,S}$ generated by (64)-(65). Then, for any $0<\delta<I_{*}$, the Borel set
$U_{I_{*}-\delta}^{\infty,I,S}=\{\Theta^{0}\in\mathbb{T}^{N}:R(\Phi_{t}^{I,S}(\Theta^{0}))<I_{*}-\delta\mbox{ for all }t\geq 0\}$
has measure zero.
Proof.
Defining for each $T\in[0,\infty)$ the open set
$U_{I_{*}-\delta}^{T,I,S}\coloneqq\{\Theta^{0}\in\mathbb{T}^{N}:R(\Phi_{t}^{I,S}(\Theta^{0}))<I_{*}-\delta\mbox{ for all }0\leq t\leq T\},$
we have that
$U_{I_{*}-\delta}^{\infty,I,S}=\bigcap_{T\in\mathbb{N}}U_{I_{*}-\delta}^{T,I,S}$
is a Borel set.
Note that
$U_{\varepsilon}^{0,I,S}=\{\Theta^{0}\in\mathbb{T}^{N}:R(\Theta^{0})<\varepsilon\},\quad\varepsilon>0,$
which by condition (21) satisfies
$\lim_{\varepsilon\to 0}m(U_{\varepsilon}^{0,I,S})=0.$
By definition, we have
$\Phi_{t}^{I,S}(U_{I_{*}-\delta}^{T,I,S})\subset U_{I_{*}-\delta}^{0,I,S},\quad\forall~0\leq t\leq T\leq\infty,~t<\infty.$
Also, since for $0<\delta_{1}<\delta_{2}<I_{*}$ and $0\leq T\leq\infty$ we have $U^{T,I,S}_{I_{*}-\delta_{2}}\subset U^{T,I,S}_{I_{*}-\delta_{1}}$, we may possibly shrink $\delta$ to assume that $\delta<I_{*}/2$.
By the fact that the divergence of a flow determines the rate of change of volume, we have for $0\leq t\leq T<\infty$
[…]
Let $0<\varepsilon<\delta$. We have
$\frac{d}{dt}m\left(\Phi_{t}^{I,S}\left(U_{I_{*}-\delta}^{T,I,S}\right)\right)$ $\geq c_{4}\kappa N\int_{\Phi_{t}^{I,S}\left(U_{I_{*}-\delta}^{T,I,S}\right)\setminus U^{0,I,S}_{\varepsilon}}R(I_{*}-R)dm(\Theta)\geq c_{4}\kappa N\varepsilon(I_{*}-\varepsilon)m\left(\Phi_{t}^{I,S}\left(U_{I_{*}-\delta}^{T,I,S}\right)\setminus U^{0,I,S}_{\varepsilon}\right)$
$\geq c_{4}\kappa N\varepsilon(I_{*}-\varepsilon)\max\left\{0,m\left(\Phi_{t}^{I,S}\left(U_{I_{*}-\delta}^{T,I,S}\right)\right)-m\left(U^{0,I,S}_{\varepsilon}\right)\right\}$
which we can solve as
$\max\left\{0,m\left(\Phi_{T}^{I,S}\left(U_{I_{*}-\delta}^{T,I,S}\right)\right)-m\left(U^{0,I,S}_{\varepsilon}\right)\right\}\geq\exp\left(c_{4}\kappa N\varepsilon(I_{*}-\varepsilon)T\right)\max\left\{0,m\left(U_{I_{*}-\delta}^{T,I,S}\right)-m\left(U^{0,I,S}_{\varepsilon}\right)\right\}.$
However, because of the bound $m\left(\Phi_{T}^{I,S}\left(U_{I_{*}-\delta}^{T,I,S}\right)\right)\leq 1$, we have
$m\left(U_{I_{*}-\delta}^{\infty,I,S}\right)\leq m\left(U_{I_{*}-\delta}^{T,I,S}\right)\leq m\left(U^{0,I,S}_{\varepsilon}\right)+\exp\left(-c_{4}\kappa N\varepsilon(I_{*}-\varepsilon)T\right),\quad\forall\varepsilon\in(0,\delta),~\forall T\in(0,\infty).$
First taking the limit $T\to\infty$ and then taking the limit $\varepsilon\to 0$ gives $m\left(U_{I_{*}-\delta}^{\infty,I,S}\right)=0$. ∎
The proof of Theorem 28 from Theorem 21 and Proposition 58 is essentially the same as the proof of Theorem 1 from Theorem 15 and Proposition 56.
Proof of Theorem 28.
Choose a sequence $\{\delta_{n}\}_{n=1}^{\infty}$ of positive real numbers such that
$\kappa>\max\left\{\frac{2(2c_{2})^{p/q}}{c_{1}c_{3}}\cdot\frac{\|\Omega\|_{\infty}}{{(I_{*}-\delta_{n})}^{1+\frac{p}{q}}},\frac{2}{c_{1}c_{3}(\pi-\alpha_{0})^{p}}\cdot\frac{\|\Omega\|_{\infty}}{{I_{*}-\delta_{n}}}\right\}\quad\forall n\geq 1,\quad\mbox{and}\quad\lim_{n\rightarrow\infty}\delta_{n}=0.$
Since $\kappa>0$, Proposition 56 tells us that for every $n\geq 1$ the set
$U_{I_{*}-\delta_{n}}^{\infty,I,S}=\{\Theta^{0}\in\mathbb{T}^{N}:R(\Phi_{t}^{I,S}(\Theta^{0}))<I_{*}-\delta_{n}\mbox{ for all }t\geq 0\}$
has measure zero and thus the union
$U^{\infty,I,S}=\mathop{\bigcup}_{n=1}^{\infty}U_{I_{*}-\delta_{n}}^{\infty,I,S}$
has measure zero as well. So, for almost all initial data $\Theta^{0}\in\mathbb{T}^{N}\backslash U^{\infty,I,S}$, for each $n\geq 1$ there exists a finite time $t_{n}\geq 0$ such that
$R(\Phi^{I,S}_{t_{n}}(\Theta^{0}))\geq I_{*}-\delta_{n},$
and we apply Theorem 21 at this time $t=t_{n}$ to conclude the following:
  1. (a)
    By Theorem 21 (a), we have $\liminf_{t\to\infty}R(t)\geq\frac{c_{3}R(t_{n})}{2}\geq\frac{c_{3}(I_{*}-\delta_{n})}{2}$. By taking the limit $n\to\infty$ we obtain assertion (a) of Theorem 28.
  2. (b)
    By Theorem 21 (b), $\limsup_{t\to\infty}\theta_{i}(t)-\liminf_{t\to\infty}\theta_{i}(t)\leq\sup_{t\geq t_{1}}\theta_{i}(t)-\inf_{t\geq t_{1}}\theta_{i}(t)<2\pi$, proving assertion (b) of Theorem 28.
Now we prepare for the proof of Theorem 30. We begin with a generalization of Lemma 54.
Lemma 59.
Assuming (22), we have
$m\Big\{\{\theta_{i}^{0}\}_{i=1}^{N}\in[-\pi,\pi]^{N}:~\frac{1}{N}\sum_{i=1}^{N}I(\theta_{i}^{0})\leq t\Big\}\leq\left(\frac{\Gamma(1/r)}{\pi r^{1-1/r}}\left(\frac{et}{c_{5}}\right)^{1/r}\right)^{N},\quad t>0.$
Assuming (19), we have
$m\Big\{\{\theta_{i}^{0}\}_{i=1}^{N}\in[-\pi,\pi]^{N}:~\frac{1}{N}\sum_{i=1}^{N}I(\theta_{i}^{0})\leq I_{*}-\delta\Big\}\leq\exp\left(-\frac{2N\delta^{2}}{\|I\|^{2}_{\sup}}\right),\quad\delta\in(0,I_{*}).$
Proof.
Assuming (22), we can compute
[…]
so that by Lemma 52 with $M=0$,
$\mathbb{P}\left[\frac{1}{N}\sum_{i=1}^{N}I(\theta_{i}^{0})\leq t\right]=\mathbb{P}\left[\frac{1}{N}\sum_{i=1}^{N}(-I(\theta_{i}^{0}))\geq-t\right]\leq\left(\frac{\Gamma(1/r)}{\pi r^{1-1/r}}\left(\frac{et}{c_{5}}\right)^{1/r}\right)^{N}.$
On the other hand, assume (19). Then
[…]
so that $I_{*}\leq R^{*}$, where $R^{*}$ is defined as in (15):
$R^{*}=\frac{1}{2\pi}\int_{-\pi}^{\pi}I(\theta)d\theta.$
The distribution $m$ means that we are independently and identically sampling $\theta_{i}^{0}$ according to the uniform Lebesgue distribution on $\pi,\pi]$, so $\frac{1}{N}\sum_{j=1}^{N}I(\theta_{j}^{0})$ is the sum of the independent random variables $\frac{1}{N}I(\theta_{j}^{0})$, $j=1,\cdots,N$. By the Hoeffding lemma, $\frac{1}{N}I(\theta_{j}^{0})$ is $\frac{\|I\|_{\sup}^{2}}{4N^{2}}$-subgaussian, so by independence, $\sum_{j=1}^{N}\frac{1}{N}I(\theta_{j}^{0})$ is $\frac{\|I\|_{\sup}^{2}}{4N}$-subgaussian. As $\mathbb{E}\sum_{j=1}^{N}\frac{1}{N}I(\theta_{j}^{0})=R^{*}$, we conclude
$m\left\{\{\theta_{i}^{0}\}_{i=1}^{N}\in[-\pi,\pi]^{N}:~\frac{1}{N}\sum_{i=1}^{N}I(\theta_{i}^{0})\leq I_{*}-\delta\right\}$ $\leq m\left\{\{\theta_{i}^{0}\}_{i=1}^{N}\in[-\pi,\pi]^{N}:~\frac{1}{N}\sum_{i=1}^{N}I(\theta_{i}^{0})\leq R^{*}-\delta\right\}$
$\leq\exp\left(-\frac{\delta^{2}}{2\cdot\|I\|^{2}_{\sup}/4N}\right)=\exp\left(-\frac{2N\delta^{2}}{\|I\|^{2}_{\sup}}\right).$
Next we obtain an analogue of Proposition 56.
Proposition 60.
Consider the Winfree model (1) with interaction functions $I$ and $S$ satisfying (19), (20), and (22). Let $\kappa>0$, $N$, and $\Omega=(\omega_{1},\cdots,\omega_{N})$ be fixed, and consider the flow $\Phi_{t}^{I,S}$ generated by (64)-(65). Then, for any $0<\delta\leq I_{*}/2$ and $T\geq 0$, the open set
$U_{\delta}^{T,I,S}\coloneqq\{\Theta^{0}\in\mathbb{T}^{N}:R(\Phi_{t}^{I,S}(\Theta^{0}))<\delta\mbox{ for all }0\leq t\leq T\}$
has measure
$m(U_{\delta}^{T,I,S})\leq\left(\max\left\{\frac{c_{5}\pi^{r}r^{r-1}}{\Gamma(1/r)^{r}e\delta},\exp\left(\frac{2r(I_{*}-\delta)^{2}}{\|I\|^{2}_{\sup}}\right)\right\}+\frac{c_{4}c_{5}\pi^{r}r^{r-1}}{\Gamma(1/r)^{r}(1+r/N)}\kappa N(I_{*}-\delta)T\right)^{-N/r}.$
Proof.
As in the proof of Proposition 56, we begin by noting that
$U_{\varepsilon}^{0,I,S}=\{\Theta^{0}\in\mathbb{T}^{N}:R(\Theta^{0})<\varepsilon\},\quad\varepsilon>0,$
and estimate its measure to be the following, via Lemma 59:
$m(U_{\varepsilon}^{0,I,S})\leq\left(\frac{\Gamma(1/r)}{\pi r^{1-1/r}}\left(\frac{e\varepsilon}{c_{5}}\right)^{1/r}\right)^{N}.$
By definition, we have
$\Phi_{t}^{I,S}(U_{\delta}^{T,I,S})\subset U_{\delta}^{0,I,S},\quad\forall~0\leq t\leq T.$
By the fact that the divergence of a flow determines the rate of change of volume, we have for $0\leq t\leq T$
[…]
Since $\Phi_{t}^{I,S}\left(U_{\delta}^{T,I,S}\right)\subset U^{0,I,S}_{\delta}$,
$\frac{d}{dt}m\left(\Phi_{t}^{I,S}\left(U_{\delta}^{T,I,S}\right)\right)$ $\geq c_{4}\kappa N(I_{*}-\delta)\int_{\Phi_{t}^{I,S}\left(U_{\delta}^{T,I,S}\right)}Rdm(\Theta),$
$=c_{4}\kappa N(I_{*}-\delta)\int_{\Phi_{t}^{I,S}\left(U_{\delta}^{T,I,S}\right)}\int_{0}^{\delta}\mathbbm{1}_{R\geq t}dtdm(\Theta)$
$=c_{4}\kappa N(I_{*}-\delta)\int_{0}^{\delta}m\left(\Phi_{t}^{I,S}\left(U_{\delta}^{T,I,S}\right)\setminus U^{0,I,S}_{t}\right)dt$
$\geq c_{4}\kappa N(I_{*}-\delta)\int_{0}^{\delta}\max\left\{0,m\left(\Phi_{t}^{I,S}\left(U_{\delta}^{T,I,S}\right)\right)-m\left(U^{0,I,S}_{t}\right)\right\}dt$
$\geq c_{4}\kappa N(I_{*}-\delta)\int_{0}^{\delta}\max\left\{0,m\left(\Phi_{t}^{I,S}\left(U_{\delta}^{T,I,S}\right)\right)-\left(\frac{\Gamma(1/r)}{\pi r^{1-1/r}}\left(\frac{et}{c_{5}}\right)^{1/r}\right)^{N}\right\}dt$
$=c_{4}\kappa N(I_{*}-\delta)\frac{\pi^{r}r^{r-1}c_{5}}{\Gamma(1/r)^{r}e}\frac{1}{1+r/N}m\left(\Phi_{t}^{I,S}\left(U_{\delta}^{T,I,S}\right)\right)^{1+\frac{r}{N}},$
where in the last equality we used $m\left(\Phi_{t}^{I,S}\left(U_{I_{*}-\delta}^{T,I,S}\right)\right)\leq m\left(U^{0,I,S}_{\delta}\right)\leq\left(\frac{\Gamma(1/r)}{\pi r^{1-1/r}}\left(\frac{e\delta}{c_{5}}\right)^{1/r}\right)^{N}$.
Solving this differential inequality along with the terminal condition
[…]
gives the stated inequality. ∎
We are now ready to prove Theorem 30.
Proof of Theorem 30.
Under the condition
$\kappa>\max\left\{\frac{4(4c_{2})^{p/q}}{c_{1}c_{3}}\cdot\frac{\|\Omega\|_{\infty}}{{I_{*}}^{1+\frac{p}{q}}},\frac{4}{c_{1}c_{3}(\pi-\alpha_{0})^{p}}\cdot\frac{\|\Omega\|_{\infty}}{{I_{*}}}\right\}.$ (23)
we have from Theorem 21 that
$m\Big\{\{\theta_{i}^{0}\}_{i=1}^{N}\in[-\pi,\pi]^{N}:R(t)>\frac{c_{3}I_{*}}{4}~\forall t\geq T\mbox{ and}$
$\qquad\qquad\qquad\qquad\qquad\qquad\sup_{t\geq T}\theta_{i}(t)-\inf_{t\geq T}\theta_{i}(t)<2\pi~\forall i=1,\cdots,N\Big\}$
$\geq m\Big\{\{\theta_{i}^{0}\}_{i=1}^{N}\in[-\pi,\pi]^{N}:~R(t)\geq I_{*}/2\mathrm{~for~some~}t\in[0,T]\Big\}$
$=1-m\left(U_{I_{*}/2}^{T,I,S}\right)$
$\geq 1-\left(\exp\left(\frac{rI_{*}^{2}}{2\|I\|^{2}_{\sup}}\right)+\frac{c_{4}c_{5}\pi^{r}r^{r-1}}{\Gamma(1/r)^{r}(1+r/N)}\kappa N(I_{*}-\delta)T\right)^{-N/r}$
where in the last inequality we used Proposition 60 with $\delta=\frac{I_{*}}{2}$.
Under condition (24), we have
$\kappa=\max\left\{\frac{2(2c_{2})^{p/q}}{c_{1}c_{3}}\frac{\|\Omega\|_{\infty}}{\varepsilon_{0}^{1+p/q}},\frac{2}{c_{1}c_{3}(\pi-\alpha_{0})^{p}}\frac{\|\Omega\|_{\infty}}{\varepsilon_{0}}\right\}$
for some $\varepsilon_{0}\in(0,I_{*}/2)$ such that $\frac{\Gamma(1/r)}{\pi r^{1-1/r}}\left(\frac{e\varepsilon_{0}}{c_{5}}\right)^{1/r}<\frac{1}{2^{1/r}}$, and we have
$m\left\{\{\theta_{i}^{0}\}_{i=1}^{N}\in[-\pi,\pi]^{N}:\sup_{t\geq T}\theta_{i}(t)-\inf_{t\geq T}\theta_{i}(t)<2\pi~\forall i=1,\cdots,N\right\}$
[…]
$=1-m\left(U_{\varepsilon_{0}}^{T,I,S}\right)$
[…]
6. On the equilibria of the Winfree model
The main purpose of this section is to prove Theorem 34. We will develop a framework to analyze the equilibria of (3).
This is similar to a framework developed by Ha, Kim, and the author in [HKR16] that is used to prove the finiteness of phase-locked states for the Kuramoto model; see Theorem 64.
6.1. Description of equilibria of (3)
In the case $\omega_{i}=0$ for all $i=1,\cdots,N$, the ODE (5) becomes
$\dot{\theta}_{i}=-\kappa R(t)\sin\theta_{i},$
and it is clear that the only equilibrium initial data are the “bipolar states”
$\theta_{i}^{0}=0\mbox{ or }\pi\mod 2\pi,\quad i=1,\cdots,N.$
In the general case where $\omega_{i}\neq 0$ for some $i$, we can expect the equilibria to be perturbations of the equilibria for the case $\omega_{i}=0$.
Indeed, fix $\kappa\neq 0$ and $\{\omega_{i}\}_{i=1}^{N}$, where $\omega_{i}\neq 0$ for some $i$. If $\{\theta_{i}^{0}\}_{i=1}^{N}$ is an equilibrium of (3), then, by (4) and (5), we must necessarily have
$R\coloneqq\frac{1}{N}\sum_{j=1}^{N}(1+\cos\theta_{j}^{0})=1+\frac{1}{N}\sum_{j=1}^{N}\cos\theta_{j}^{0}$
and
$0=\omega_{i}-\kappa R\sin\theta_{i}^{0},\quad i=1,\cdots,N.$
Clearly, we must have $R\geq\max_{i}\frac{|\omega_{i}|}{|\kappa|}$ (and a fortiori $R>0$). Now, for some choice $\sigma_{i}\in\{-1,1\}$ of signs,
$\sin\theta_{i}^{0}=\frac{\omega_{i}}{\kappa R},~\cos\theta_{i}^{0}=\sigma_{i}\sqrt{1-\frac{\omega_{i}^{2}}{\kappa^{2}R^{2}}},\quad i=1,\cdots,N,$
or equivalently
$\theta_{i}^{0}=\begin{cases}\sin^{-1}\frac{\omega_{i}}{\kappa R}\mod 2\pi&\mathrm{if}~\sigma_{i}=1,\\ \pi-\sin^{-1}\frac{\omega_{i}}{\kappa R}\mod 2\pi&\mathrm{if}~\sigma_{i}=-1,\\ \end{cases}\quad i=1,\cdots,N,$ (67)
and
$R=1+\frac{1}{N}\sum_{j=1}^{N}\sigma_{j}\sqrt{1-\frac{\omega_{j}^{2}}{\kappa^{2}R^{2}}}.$ (68)
Conversely, if (68) holds with some $\sigma_{i}\in\{-1,1\}$, $i=1,\cdots,N,$ and $R>0$, then it is easy to see that $\{\theta_{i}^{0}\}_{i=1}^{N}$ given as (67) is an equilibrium of (3).
We can summarize the discussion so far into the following proposition.
Proposition 61.
Fix $\kappa\neq 0$ and $\{\omega_{i}\}_{i=1}^{N}$, where $\omega_{i}\neq 0$ for some $i$. Given any $R\in[\max_{i}\frac{|\omega_{i}|}{|\kappa|},\infty)$ and $\{\sigma_{i}\}_{i=1}^{N}\subseteq\{-1,1\}$, the initial data (67) is an equilibrium for (3) if and only if (68) holds. There are no equilibria of (3) other than the ones just described.
Based on the characterization of equilibria provided by Proposition 61, one can easily construct equilibria with a given structure (namely, the signs $\sigma_{i}$’s) using the intermediate value theorem.
Proposition 62.
Fix $\kappa\neq 0$ and $\{\omega_{i}\}_{i=1}^{N}$. Let $\{\sigma_{i}\}_{i=1}^{N}\subseteq\{-1,1\}$, with not every $\sigma_{i}$ being 1-1, and denote
$\rho_{0}\coloneqq 1+\frac{1}{N}\sum_{j=1}^{N}\sigma_{j}>0.$
If
$\max_{i}\frac{|\omega_{i}|}{|\kappa|}<\frac{\rho_{0}^{1.5}}{4},$ (69)
then there exists $R\in[\frac{1}{2}\rho_{0},\frac{3}{2}\rho_{0}]$ such that the initial data (67) is an equilibrium to (3).
Proof.
By Proposition 61, it is enough to show that there exists a solution $R$ to (68) in the interval $\frac{1}{2}\rho_{0},\frac{3}{2}\rho_{0}]$. Define the function $f:[\max_{i}\frac{|\omega_{i}|}{|\kappa|},\infty)\to\mathbb{R}$ by
$f(r)\coloneqq r-1-\frac{1}{N}\sum_{j=1}^{N}\sigma_{j}\sqrt{1-\frac{\omega_{j}^{2}}{\kappa^{2}r^{2}}},\quad r\in[\max_{i}\frac{|\omega_{i}|}{|\kappa|},\infty).$
It is enough to see that $f$ has a zero in $\frac{1}{2}\rho_{0},\frac{3}{2}\rho_{0}]$. Using the fact that $|1-\sqrt{1-x}|\leq x$ for $0\leq x\leq 1$, we have
[…]
Observing that $\frac{1}{2}\rho_{0},\frac{3}{2}\rho_{0}]\subset[\max_{i}\frac{|\omega_{i}|}{|\kappa|},\infty)$,
[…]
and
[…]
we conclude by the intermediate value theorem that $f$ does indeed have a zero in $\frac{1}{2}\rho_{0},\frac{3}{2}\rho_{0}]$. ∎
We are now ready to prove our main result on the existence of certain equilibria, namely Theorem 34.
Proof of Theorem 34.
As $N\geq\frac{2}{\rho}$, we may find $1\leq m\leq N$ such that
$\frac{2m}{N}\leq\rho<\frac{2m+2}{N}.$
Choose $\sigma_{1}=\cdots=\sigma_{m}=1$ and $\sigma_{m+1}=\cdots=\sigma_{N}=-1$. Then, under the notation of Proposition 62, we have $\rho_{0}=\frac{2m}{N}$, so $\rho_{0}\leq\rho<2\rho_{0}$. Now
$\max_{i}\frac{|\omega_{i}|}{|\kappa|}<\frac{\rho^{1.5}}{16}<\frac{\rho_{0}^{1.5}}{4},$
so by Proposition 62 there exists $R\in[\frac{1}{2}\rho_{0},\frac{3}{2}\rho_{0}]\subset[\frac{1}{4}\rho,\frac{3}{2}\rho]$ such that the initial data (67) is an equilibrium to (3).
Moreover, it is easy to check that
$\frac{\rho}{4}<\frac{m}{N}\leq\frac{\rho}{2},$
while
[…]
and
[…]
6.2. The critical coupling strength
Based on the characterization (67)-(68) of equilibria by Proposition 61, we can compute the critical coupling strength for the Winfree model (3). Note that the analogous computation for the Kuramoto model (32) has been performed in [VM09].
Proposition 63.
Given $\Omega=\{\omega_{i}\}_{i=1}^{N}\neq 0$, we have
$\kappa_{\mathrm{c}}(\Omega)\coloneqq\frac{Nu_{*}}{N+\sum_{j=1}^{N}\sqrt{1-\frac{\omega_{j}^{2}}{u_{*}^{2}}}},$
where $u_{*}\in[\|\Omega\|_{\infty},\frac{2}{\sqrt{3}}\|\Omega\|\infty]$ is the solution to
$N+2\sum_{j=1}^{N}\sqrt{{1-\frac{\omega_{j}^{2}}{u_{*}^{2}}}}=\sum_{j=1}^{N}\frac{1}{\sqrt{1-\frac{\omega_{j}^{2}}{u_{*}^{2}}}}.$
More precisely, given $\Omega$ and $\kappa>0$, (3) admits an equilibrium initial data $\Theta^{0}$ if and only if $\kappa\geq\kappa_{\mathrm{c}}(\Omega)$. Also,
$\frac{2N\|\Omega\|_{\infty}}{4N-1}\leq\frac{16N\|\Omega\|_{\infty}}{(6N-3+\sqrt{4N^{2}-4N+9})\sqrt{5-2N+\sqrt{4N^{2}-4N+9}}\sqrt{3+2N-\sqrt{4N^{2}-4N+9}}}\leq\kappa_{\mathrm{c}}(\Omega)\leq\frac{4}{3\sqrt{3}}\|\Omega\|_{\infty},$
the equality case for the second inequality being when $\omega_{i}=0$ for all but at most one $i$, and the equality case for the third inequality being when $|\omega_{i}|=\|\Omega\|_{\infty}$ for all $i$.
Proof.
We first observe that (3) has an equilibrium if and only if it has an equilibrium of the form (67)-(68) with $\sigma_{i}=1$ for all $i$. The if direction is obvious; to see the only if direction, assume (3) has an equilibrium, which by Proposition 61 must be of the form (67)-(68) for some choice of $R=R_{1}\leq 2$ and $\sigma_{i}$’s. It follows that
$1+\frac{1}{N}\sum_{j=1}^{N}\sqrt{1-\frac{\omega_{j}^{2}}{\kappa^{2}R_{1}^{2}}}-R_{1}\geq 1+\frac{1}{N}\sum_{j=1}^{N}\sigma_{i}\sqrt{1-\frac{\omega_{j}^{2}}{\kappa^{2}R_{1}^{2}}}-R_{1}=0.$
As $1+\frac{1}{N}\sum_{j=1}^{N}\sqrt{1-\frac{\omega_{j}^{2}}{\kappa^{2}\cdot 2^{2}}}-2\leq 0$, it follows from the intermediate value theorem that there is $R_{2}\in[R_{1},2]$ such that (68) holds for $R=R_{2}$ and $\sigma_{i}=1$ for all $i$. This completes the proof of our observation.
By the transformation $u=\kappa R$, we see that (3) has an equilibrium if and only if
$\frac{1}{\kappa}=\frac{1+\frac{1}{N}\sum_{j}\sqrt{1-\omega_{j}^{2}/u^{2}}}{u}\quad\mathrm{for~some~}u\in[\|\Omega\|_{\infty},\infty).$
The right-hand side converges to 0 as $u\to\infty$ and is a unimodal function of $u$, increasing on $\|\Omega\|_{\infty},u_{*}]$ and decreasing on $u_{*},\infty)$, because it has derivative
$\frac{1}{u^{2}}\left(-1-\frac{2}{N}\sum_{j=1}^{N}\sqrt{{1-\frac{\omega_{j}^{2}}{u^{2}}}}+\frac{1}{N}\sum_{j=1}^{N}\frac{1}{\sqrt{1-\frac{\omega_{j}^{2}}{u^{2}}}}\right),$
the term in parentheses being a decreasing function of $u$ and taking the value $\infty$ at $u=\|\Omega\|_{\infty}$ and taking a value $\leq 0$ at $u=\frac{2}{\sqrt{3}}\|\Omega\|_{\infty}$. This completes the proof that (3) has an equilibrium if and only if $\kappa\geq\kappa_{\mathrm{c}}(\Omega)$.
We have
$\frac{1}{\kappa_{\mathrm{c}}(\Omega)}=\max_{u\in[\|\Omega\|_{\infty},\infty)}\frac{1+\frac{1}{N}\sum_{j}\sqrt{1-\omega_{j}^{2}/u^{2}}}{u}.$
Given $\|\Omega\|_{\infty}$, the function to be maximized is pointwise minimized when $|\omega_{j}|=\|\Omega\|_{\infty}$ for all $i$ and is maximized when $\omega_{i}=0$ for all but one $i$. This gives our equality cases. It is easy to compute that in these cases $u_{*}=\frac{2}{\sqrt{3}}\|\Omega\|_{\infty}$ and
$u_{*}=4\|\Omega\|_{\infty}/\sqrt{5-2N+\sqrt{4N^{2}-4N+9}}\sqrt{3+2N+\sqrt{4N^{2}-4N+9}},$
respectively, and that $\kappa_{\mathrm{c}}(\Omega)$ computes to the above values. The first inequality follows from
$\frac{1}{\kappa_{\mathrm{c}}(\Omega)}$ $=\max_{u\in[\|\Omega\|_{\infty},\infty)}\frac{1+\frac{1}{N}\sum_{j}\sqrt{1-\omega_{j}^{2}/u^{2}}}{u}\leq\max_{u\in[\|\Omega\|_{\infty},\infty)}\frac{1+\frac{N-1}{N}+\frac{1}{N}\sqrt{1-\|\Omega\|_{\infty}^{2}/u^{2}}}{u}$
$\leq\max_{u\in[\|\Omega\|_{\infty},\infty)}\frac{2N-1}{Nu}+\max_{u\in[\|\Omega\|_{\infty},\infty)}\frac{1}{N\|\Omega\|_{\infty}}\sqrt{\frac{\|\Omega_{\infty}^{2}}{u^{2}}\left(1-\frac{\|\Omega\|_{\infty}^{2}}{u^{2}}\right)}$
$=\frac{4N-1}{2N\|\Omega\|_{\infty}}.$
6.3. A polynomial description for the equilibria
In this subsection, we prove Theorem 36. We provide a method of understanding the behavior of equilibria of the sinusoidal Winfree model (3) in terms of a single univariate polynomial, which we will denote by $W(\cdot)$. We have seen in subsection 6.1 that the zero of an algebraic function serves as the order parameter of an equilibrium. If we multiply these functions over all the different signatures $\sigma_{i}$, this will happen to give us a rational function, with the property that its roots over $\frac{\|\Omega\|_{\infty}}{|\kappa|},\infty)$ are precisely the order parameters of the possible equilibria; we will let $W(\cdot)$ be the numerator of this rational function. With some careful analysis, it will turn out that the number of possible different configurations corresponding to a root of $W(\cdot)$ is bounded by the order of the zero with respect to $W(\cdot)$, and by bounding the degree of $W(\cdot)$ we have the bound stated in Theorem 36 on the total number of equilibria.
We remark that this polynomial description was first developed by the Ha, Kim, and the author in [HKR16], where we proved the following theorem for the Kuramoto model:
Theorem 64 ([HKR16, Theorem 6.1]).
For fixed data $(N,\{\Omega_{i}\}_{i=1}^{N})$ with $\kappa>0$, there are at most $2^{N}$ phase-locked states (i.e., relative equilibria) of the Kuramoto model (32) with positive order parameter, up to modulo $2\pi$ shifts and $\mathrm{U}(1)$ symmetry.
The proof method of Theorem 64 presented in [HKR16] is similar to the proof method of Theorem 36 that we present here. We also remark that this proof works only when the Winfree model (3) has the given specific sinusoidal interaction functions. It may also work (with worse results) when the interaction functions are trigonometric polynomials, or more generally when $I$ and $S$ are algebraically dependent analytic functions. However, it is likely that Theorem 36 will fail when we require less than analytic regularity on the interaction functions, say if we only require them to belong in the class $C^{2}$.
It is trivial that Theorem 36 holds when $\Omega=0$ (recall that in the beginning of subsection 6.1 we fully classified the equilibria into the bipolar states), so we will assume that $\omega_{i}\neq 0$ for some $i$.
To prove Theorem 36, let us begin by observing the following algebraic result, whose proof we omit.
Lemma 65.
Let $x_{1},\ldots,x_{N+1}$ be real variables. Then, the polynomial $S$ defined by
$S(x_{1},\ldots,x_{N+1})\coloneqq\prod_{(\sigma_{1},\ldots,\sigma_{N})\in\{-1,1\}^{N}}\left(\sum_{k=1}^{N}\sigma_{j}x_{j}+x_{N+1}\right)$
is a homogeneous polynomial of degree $2^{N-1}$ in $x_{1}^{2},\ldots,x_{N+1}^{2}$.
For $\Sigma=(\sigma_{1},\ldots,\sigma_{N})\in\{-1,1\}^{N}$, we introduce the auxiliary function
$P(r,\Sigma)\coloneqq 1+\frac{1}{N}\sum_{j=1}^{N}\sigma_{j}\sqrt{1-\left(\frac{\omega_{j}}{\kappa r}\right)^{2}}-r,\quad r\in\mathbb{R}.$
This is the function from (68) whose zeros were the order parameters of equilibria. We denote their normalized product
$W(r)\coloneqq r^{2^{N}}\prod_{\Sigma\in\{-1,1\}^{N}}P(r,\Sigma).$
Then, using Lemma 65 with
$x_{j}=\frac{1}{N}\sqrt{1-\left(\frac{\omega_{j}}{\kappa r}\right)^{2}},~j=1,\cdots,N,\quad x_{N+1}=1-r,$
we see that $\prod_{\Sigma\in\{-1,1\}^{N}}P(r,\Sigma)$ is a polynomial of homogeneous degree $2^{N-1}$ in $\frac{1}{N^{2}}\left(1-\left(\frac{\omega_{j}}{\kappa r}\right)^{2}\right)$, $j=1,\cdots,N$, and $(1-r)^{2}$, so $W(r)$ is a polynomial of degree at most $2^{N+1}$ in $r$. It is not hard to see that $W(\cdot)$ has degree exactly $2^{N+1}$.
Note that for any $r\in(0,2]$,
[…]
In fact, this correspondence can be made more precise, i.e., the number of distinct equilibria with order parameter $r$ is bounded above by the multiplicity of $r$ as a zero of $W$.
Proposition 66.
Suppose that $R\in\Big[\frac{\|\Omega\|_{\infty}}{|\kappa|},2\Big]$ is the order parameter of $m$ distinct (modulo $2\pi$) equilibria of the system (3). Then $R$ is a zero of $W$ of multiplicity at least $m$.
Proof.
$\bullet$ Case A: $R>\frac{\|\Omega\|_{\infty}}{|\kappa|}$. In this case
$\frac{\|\Omega\|_{\infty}}{|\kappa|R}<1,$
so for each of the $m$ equilibria there exists a unique $\Sigma\in\{-1,1\}^{N}$ such that the equilibrium is described as in (67). In other words, there are $m$ distinct $\Sigma_{1},\ldots,\Sigma_{m}\in\{-1,1\}^{N}$ satisfying $P(R,\Sigma_{i})=0$, and since $P(r,\Sigma_{i})$ is differentiable at $r=R$, we easily see that $R$ is a zero of $W(r)$ of multiplicity at least $m$.
$\bullet$ Case B: $R=\frac{\|\Omega\|_{\infty}}{|\kappa|}$.
By rearranging the indices we may assume, for some $1\leq a\leq N$,
$|\omega_{i}|=\|\Omega\|_{\infty}\mathrm{~for~}i=1,\cdots,a,\quad|\omega_{i}|<\|\Omega\|_{\infty}\mathrm{~for~}i=a+1,\cdots,N.$
For $\Sigma=(\sigma_{1},\ldots,\sigma_{N})$, define the class
$\Sigma]=\{(\tau_{1},\ldots,\tau_{a},\sigma_{a+1},\ldots,\sigma_{N}):\tau_{1},\ldots,\tau_{a}\in\{-1,1\}\}\subseteq\{-1,1\}^{N}.$
It is clear that these classes partition $\{-1,1\}^{N}$. As in Case A, for each equilibrium $\Theta^{*}$ there exists a unique class $\Sigma]$, all of whose elements along with $R$ give the equilibrium $\Theta^{*}$ via (67). Since there are $m$ distinct equilibria with order parameter $R$, there are $m$ distinct classes $\Sigma_{1}],\ldots,[\Sigma_{m}]$, with each element $\Sigma$ of each class $\Sigma_{i}]$ satisfying
$P(R,\Sigma)=0.$
For $\Sigma=(\sigma_{1},\ldots,\sigma_{N})\in\{-1,1\}^{N}$, define
$Q(r,[\Sigma])\coloneqq\prod_{\tau_{1},\ldots,\tau_{a}\in\{-1,1\}}P(r,(\tau_{1},\ldots,\tau_{a},\sigma_{a+1},\ldots,\sigma_{N}))=\prod_{T\in[\Sigma]}P(r,T).$
This is clearly well-defined, and
$Q(R,[\Sigma_{1}])=\cdots=Q(R,[\Sigma_{a}])=0.$
while again by Lemma 65 we see that $Q(r,[\Sigma_{i}])$ is a polynomial of
$\frac{1}{N}\left(1-\left(\frac{\omega_{1}}{\kappa r}\right)^{2}\right),\cdots,\frac{1}{N}\left(1-\left(\frac{\omega_{a}}{\kappa r}\right)^{2}\right),\mbox{ and }\left(1+\frac{1}{N}\sum_{j=a+1}^{N}\sigma_{j}\sqrt{1-\left(\frac{\omega_{j}}{\kappa r}\right)^{2}}-r\right)^{2}.$
Hence, each $Q(r,[\Sigma_{i}])$ is differentiable at $\|\Omega\|_{\infty}/|\kappa|$, and we again see that $W(r)$ has a zero of multiplicity at least $m$ at $R$. ∎
Now, based on the fact that $W$ is a polynomial of degree $2^{N+1}$ in $r$, Theorem 36 now follows from Proposition 66.
Given that the polynomial $W(r)$ encodes the equilibria of (3), it seems likely that $W$ carries some physical significance.
Question 67.
What is the physical significance of the polynomial $W(r)$ to the model (3), besides the fact that its zeros encode the order parameters of the equilibria?
Appendix A Proof of inequality (52)
We need to verify
$K_{c}\geq\frac{1}{\rho\sqrt{\mu(2-\mu)}},\quad 0<R_{0}\leq 2.$
For $1\leq R_{0}\leq 2$, this is just the inequality
$2(2-R_{0})+\frac{4}{3\sqrt{3}}(R_{0}-1)\geq\frac{16}{\left(3R_{0}-3+\sqrt{R_{0}^{2}-2R_{0}+9}\right)\sqrt{3+R_{0}-\sqrt{R_{0}^{2}-2R_{0}+9}}\sqrt{5-R_{0}+\sqrt{R_{0}^{2}-2R_{0}+9}}},$
and since the right-hand side is $2$ when $R_{0}=1$ and $\frac{4}{3\sqrt{3}}$ when $R_{0}=2$, it is enough to see that the right-hand side is convex. Indeed, its second derivative with respect to $R_{0}$ is
$\frac{1024\left((R_{0}-1)(-3R_{0}^{4}+12R_{0}^{3}-32R_{0}^{2}+40R_{0}+99)+(3R_{0}^{4}-4R_{0}^{3}+20R_{0}^{2}-16R_{0}+33)\sqrt{R_{0}^{2}-2R_{0}+9}\right)}{\sqrt{R_{0}^{2}-2R_{0}+9}\left(\left(-\sqrt{R_{0}^{2}-2R_{0}+9}+R_{0}+3\right)\left(\sqrt{R_{0}^{2}-2R_{0}+9}-R_{0}+5\right)\right)^{5/2}\left(\sqrt{R_{0}^{2}-2R_{0}+9}+3R_{0}-3\right)^{3}}$
which is easily seen to be positive when $1\leq R_{0}\leq 2$, as each term in brackets in the above expression is positive.
For $0<R_{0}\leq 1$, the inequality is equivalent to
$1\geq\frac{\left(3-3R_{0}+\sqrt{R_{0}^{2}-2R_{0}+9}\right)\sqrt{3+R_{0}+\sqrt{R_{0}^{2}-2R_{0}+9}}}{2\sqrt{2}(2-R_{0})\sqrt{5-R_{0}+\sqrt{R_{0}^{2}-2R_{0}+9}}},$
and since the right-hand side is 1 when $R_{0}=1$, it is enough to see that the right-hand side is increasing in $R_{0}$. Indeed, its derivative with respect to $R_{0}$ is
$\frac{\sqrt{2}\left((R_{0}-3)^{2}+3(1-R_{0})\sqrt{R_{0}^{2}-2R_{0}+9}\right)}{(2-R_{0})\sqrt{R_{0}^{2}-2R_{0}+9}\left(\sqrt{R_{0}^{2}-2R_{0}+9}-R_{0}+5\right)^{3/2}\sqrt{\sqrt{R_{0}^{2}-2R_{0}+9}+R_{0}+3}}>0,\quad 0<R_{0}<1.$